### v1

# Provable test-time adaptivity and distributional robustness of in-context learning

Abstract
We study in-context learning problems where a Transformer is pretrained on tasks drawn from a mixture distribution $\pi=\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$, called the pretraining prior, in which each mixture component $\pi_{\alpha}$ is a distribution on tasks of a specific difficulty level indexed by $\alpha$. Our goal is to understand the performance of the pretrained Transformer when evaluated on a different test distribution $\mu$, consisting of tasks of fixed difficulty $\beta\in\mathcal{A}$, and with potential distribution shift relative to $\pi_{\beta}$, subject to the chi-squared divergence $\chi^{2}(\mu,\pi_{\beta})$ being at most $\kappa$. In particular, we consider nonparametric regression problems with random smoothness, and multi-index models with random smoothness as well as random effective dimension. We prove that a large Transformer pretrained on sufficient data achieves the optimal rate of convergence corresponding to the difficulty level $\beta$, uniformly over test distributions $\mu$ in the chi-squared divergence ball. Thus, the pretrained Transformer is able to achieve faster rates of convergence on easier tasks and is robust to distribution shift at test time. Finally, we prove that even if an estimator had access to the test distribution $\mu$, the convergence rate of its expected risk over $\mu$ could not be faster than that of our pretrained Transformers, thereby providing a more appropriate optimality guarantee than minimax lower bounds.
1 Introduction
Transformers (Vaswani et al., 2017) have emerged as one of the dominant architectures in modern machine learning, as they have achieved state-of-the-art performance in many domains such as natural language processing (Devlin et al., 2019; Brown et al., 2020), computer vision (Dosovitskiy et al., 2021; Liu et al., 2021) and protein prediction (Jumper et al., 2021). A striking ability of pretrained Transformers, first observed by Brown et al. (2020) in Large Language Models (LLMs), is the phenomenon of in-context learning (ICL): given a prompt containing examples and a query, Transformers can learn the underlying pattern from the examples and produce accurate output for the query, without updating its parameters. For instance, given the values of the function $(x,y)\mapsto xy^{2}$ or $\text{subfield}\mapsto\text{field}$’ at a few distinct points, LLMs are able to recover the value of the function at a new point:
$\overbrace{\underbrace{f(1,1)=1,\;f(1,2)=4,\;f(2,2)=8,\;f(2,3)=18,}_{\text{examples}}\;\underbrace{f(3,3)=}_{\text{query}}}^{\text{prompt}}\underbrace{27\vphantom{f(2,2)}}_{\text{output}},$ (1)
$\overbrace{\underbrace{\text{Genetics}\to\text{Biology},\;\text{Relativity}\to\text{Physics},}_{\text{examples}}\;\underbrace{\text{Topology}\to}_{\text{query}}}^{\text{prompt}}\underbrace{\text{Mathematics}\vphantom{y}}_{\text{output}}.$
The ICL ability of Transformers was also exploited by Hollmann et al. (2025) to build tabular foundation models, which they show to outperform many classical machine learning methods, such as tree-based and boosting algorithms, on regression and classification problems for tabular data.
Extensive studies on ICL have been conducted from various perspectives (e.g. Garg et al., 2022; Xie et al., 2022; Von Oswald et al., 2023; Bai et al., 2023; Kim, Nakamaki and Suzuki, 2024); we refer the readers to Section 1.1 for a more detailed literature review. However, most theoretical analyses of ICL have focused on the setting where the distribution of pretraining data (which we call the pretraining prior) and the distribution of the test data (which we call the test distribution) are the same. In practice, Transformers are pretrained on a variety of tasks of different difficulties, and the test distribution will typically exhibit distribution shift from the pretraining prior. For example, we may train a Transformer on translation tasks between different languages, where Dutch–English translation is typically easier than Chinese–English, since Dutch and English are both West Germanic languages with shared structure. At test time, the Transformer is required to translate between two specific languages, while some linguistic patterns may appear more frequently in the test data than in the pretraining data. To formulate a theoretical framework, we focus on regression problems (in line with most prior work) and consider a Transformer pretrained on a mixture distribution containing tasks of different difficulties, but where the test distribution consists of tasks of a specific difficulty with potential distribution shift from the corresponding mixture component of the pretraining prior. We study the following natural but important questions:
Do pretrained Transformers adapt to the difficulty of the test distribution, and are they robust to distribution shift?
Our contributions
  • In Proposition 1, we provide a key test-time risk upper bound, which decomposes into a term that captures both the degree of distribution shift and the proximity of the pretrained Transformer to the posterior regression function, and another term that represents the rate of convergence of the posterior regression function to the true regression function.
  • In Section 3.2, we prove a universal approximation theorem for Transformers with softmax attention and non-polynomial activation in FFN layers. We further show that a large Transformer pretrained on sufficient data can learn the posterior regression function arbitrarily well when the activation in FFN layers is one of $\mathsf{ReLU}$, $\mathsf{GELU}$ or $\mathsf{SiLU}$.
  • In Sections 4.1 and 4.2, we consider general nonparametric regression problems and multi-index models respectively, with regression functions belonging to Besov smoothness classes. We prove that the posterior regression function is optimally adaptive to the difficulty of every task in the support of the pretraining prior (and hence robust to distribution shift). In combination with the results mentioned above, this establishes the adaptivity and distributional robustness of pretrained Transformers.
  • Finally, in Section 4.3, we show that the expected risk of any estimator over the test distribution $\mu$ cannot achieve a faster rate of convergence than our pretrained Transformers (even if the estimator is allowed to depend on $\mu$). This provides a more appropriate form of optimality guarantee for pretrained Transformers than that yielded by minimax lower bounds usually considered in ICL literature.
1.1 Related work
In-context learning
There is a large body of theoretical research that aims to explain the ICL ability of Transformers. One line of work shows that Transformers can perform ICL by approximating certain algorithms, such as gradient descent (Ahn et al., 2023; Akyürek et al., 2023; Bai et al., 2023; Von Oswald et al., 2023), functional gradient descent (Cheng, Chen and Sra, 2024), basis function regression (Kim, Nakamaki and Suzuki, 2024), reinforcement learning (Lin, Bai and Mei, 2024) and domain adaptation (Hataya, Matsui and Imaizumi, 2024). Other lines of work analyse ICL through the lens of Bayesian inference (Xie et al., 2022; Zhang et al., 2022; Nagler, 2023; Wang et al., 2023; Deora et al., 2025; Zhang et al., 2025), meta-learning (Dai et al., 2023; Jeon et al., 2024; Wu, Wang and Yao, 2025) and training dynamics (Oko et al., 2024; Zhang, Frei and Bartlett, 2024; Kumano, Kera and Yamasaki, 2025; Kuwataka and Suzuki, 2025). In particular, Bai et al. (2023) show that Transformers can perform gradient descent on penalised linear regression, generalised linear models and two-layer neural networks, as well as implement in-context algorithm selection, while Kim, Nakamaki and Suzuki (2024) prove that Transformers are minimax optimal for nonparametric regression. Both papers consider the case where the test distribution coincides with the pretraining prior. On the other hand, Xie et al. (2022) study the setting where the test prompts have different formats compared to the pretraining data, and provide consistency results under a Bayesian hidden Markov model. There are also numerous empirical studies on the behaviour and properties of ICL (e.g. Garg et al., 2022; Kirsch et al., 2022; Panwar, Ahuja and Goyal, 2024; Reuter et al., 2025). Finally, we note the concurrent and independent work of Wakayama and Suzuki (2025), who study the adaptivity of ICL from a Bayesian perspective. Their results are complementary to ours, and neither subsumes the other.
Posterior contraction theory
Posterior contraction theory is an active area of research in statistics that studies the concentration of the posterior distribution around a true data generating distribution (i.e. from a frequentist perspective). Pioneered by Ghosal, Ghosh and van der Vaart (2000) and Shen and Wasserman (2001), posterior contraction for Bayesian nonparametric models has been investigated in different problems (e.g. Ghosal, Lember and van der Vaart, 2008; van der Vaart and van Zanten, 2009; Finocchio and Schmidt-Hieber, 2023; Egels and Castillo, 2025). In this paper, we modify the tools developed by Ghosal, Lember and van der Vaart (2008) to derive rates of convergence for the posterior regression function. Interested readers are referred to the book by Ghosal and van der Vaart (2017) for an introduction to the field.
1.2 Notation
For $n\in\mathbb{N}$, we write $n]\coloneqq\{1,\ldots,n\}$. For $m,n\in\mathbb{N}$ with $m<n$, we write $m:n]\coloneqq\{m,m+1,\ldots,n\}$ and similarly $n:\infty)\coloneqq\{n,n+1,\ldots\}$. Given $a,b\in\mathbb{R}$, we write $a\wedge b\coloneqq\min\{a,b\}$ and $a\vee b\coloneqq\max\{a,b\}$. For $d\in\mathbb{N}$, we define $\bm{0}_{d}\coloneqq(0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ and $\bm{1}_{d}\coloneqq(1,\ldots,1)^{\top}\in\mathbb{R}^{d}$. For $R>0$, we define the clipping operator $\mathsf{clip}_{R}:\mathbb{R}\to[-R,R]$ by $\mathsf{clip}_{R}(x)\coloneqq-R\vee x\wedge R$ for $x\in\mathbb{R}$. Given a Borel measurable set $A\subseteq\mathbb{R}^{d}$, we write $\mathrm{Vol}_{d}(A)$ for the $d$-dimensional Lebesgue measure of $A$. For two $\sigma$-finite measures $P$ and $Q$ on a measurable space $(\mathcal{X},\Sigma)$, we say that $P$ is absolutely continuous with respect to $Q$, written as $P\ll Q$, if $P(A)=0$ whenever $A\in\Sigma$ is such that $Q(A)=0$, and in that case write $\frac{\mathrm{d}P}{\mathrm{d}Q}$ for the Radon–Nikodym derivative. When $P$ and $Q$ are probability measures, the $\chi^{2}$-divergence from $Q$ to $P$ is defined as
$\chi^{2}(P,Q)\coloneqq\begin{cases}\int_{\mathcal{X}}\bigl(\frac{\mathrm{d}P}{\mathrm{d}Q}\bigr)^{2}\,\mathrm{d}Q-1\quad&\text{if }P\ll Q\\ \infty&\text{otherwise}.\end{cases}$
Further, we write $P^{\otimes n}$ for the $n$-fold product measure of $P$ and $\mathrm{supp}(P)$ for the support of $P$, i.e. the intersection of all closed $A\in\Sigma$ with $P(A)=1$. For a set $\mathcal{F}$ of functions from $\mathcal{X}$ to $\mathcal{Y}$ and a set $\mathcal{G}$ of functions from $\mathcal{Y}$ to $\mathcal{Z}$, we write $\mathcal{F}\circ\mathcal{G}\coloneqq\{f\circ g:f\in\mathcal{F},\,g\in\mathcal{G}\}$, where $f\circ g$ denotes the composition of $f$ and $g$. For a function $f:\mathcal{X}\to\mathbb{R}$, we write $\|f\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|f(x)|$.
2 Problem set-up and ICL excess risk decomposition
We first present the generating mechanism of our pretraining data. Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}\subseteq\mathbb{R}$ be Borel measurable, and let $\mathcal{P}$ be a measurable space of probability distributions on $\mathcal{X}\times\mathcal{Y}$, dominated by a common $\sigma$-finite measure. The pretraining prior $\pi$ is a distribution on $\mathcal{P}$, and our pretraining data take the form $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$, generated as follows:
  1. 1.
    Draw random distributions $P_{1},\ldots,P_{T}\overset{\mathrm{iid}}{\sim}\pi$.
  2. 2.
    For each $t\in[T]$, draw $\mathcal{D}_{n}^{(t)}\coloneqq(X_{i}^{(t)},Y_{i}^{(t)})_{i=1}^{n}\sim P_{t}^{\otimes n}$ and $(X^{(t)},Y^{(t)})\sim P_{t}$ independently.
Thus, for each $t\in[T]$, $\mathcal{D}_{n}^{(t)}$ consists of $n$ examples, $X^{(t)}$ is our query and our goal is to predict its corresponding output $Y^{(t)}$; see (1). We emphasise that the test distribution $\mu$ on $\mathcal{P}$ of new prompts at test time will typically be different from the (possibly unknown) distribution $\pi$ that generated our pretraining data.
Next, for a class $\mathcal{F}$ of measurable functions from $\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}$ to $\mathcal{Y}$ (e.g. Transformers), define the empirical risk minimiser $\widehat{f}_{T}$ over $\mathcal{F}$ by
$\widehat{f}_{T}\in\operatorname*{argmin}_{f\in\mathcal{F}}\widehat{\mathcal{R}}_{T}(f),\quad\text{where }\widehat{\mathcal{R}}_{T}(f)\coloneqq\frac{1}{T}\sum_{t=1}^{T}\bigl\{Y^{(t)}-f(X^{(t)},\mathcal{D}_{n}^{(t)})\bigr\}^{2}$ (2)
and $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$ are generated according to $\pi$.
We remark that, given a new set of examples $D_{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$, the prediction at a query $x\in\mathcal{X}$ is given by $\widehat{f}_{T}(x,D_{n})$, and we do not need to update $\widehat{f}_{T}$ accordingly.
For a distribution $\nu$ on $\mathcal{P}$, we define the $\nu$-risk of a measurable $g:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ as [^1] We adopt the convention that for $P\in\mathcal{P}$, $\mathbb{E}_{P}$ means that the random variables inside the expectation are generated from $P$, i.e. $(\mathcal{D}_{n},X,Y)\sim P^{\otimes(n+1)}$, whereas for a distribution $\nu$ on $\mathcal{P}$, $\mathbb{E}_{\nu}\mathbb{E}_{P}$ means that the distribution $P$ is also random and drawn from $\nu$.
$\mathcal{R}_{\nu}(g)$ $\coloneqq\mathbb{E}_{P\sim\nu}\mathbb{E}_{(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]\equiv\mathbb{E}_{\nu}\mathbb{E}_{P}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr].$
If $\widetilde{f}_{T}$ is random and depends on the pretraining data $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$, then we interpret $\mathcal{R}_{\nu}(\widetilde{f}_{T})$ as the expectation conditional on the pretraining data, which is assumed to be independent of the test data $(\mathcal{D}_{n},X,Y)$, so $\mathcal{R}_{\nu}(\widetilde{f}_{T})$ is also random. For a test distribution $\mu$ on $\mathcal{P}$, we further define the ICL excess $\mu$-risk of $\widetilde{f}_{T}$ by
$\mathcal{R}^{\mathrm{ICL}}_{\mu}(\widetilde{f}_{T})$ $\coloneqq\mathbb{E}\mathcal{R}_{\mu}(\widetilde{f}_{T})-\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
$\hskip 2.56064pt=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$ (3)
where the term $\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ is the oracle risk assuming knowledge of the distribution $P$ that generates $(X,Y)$, and in particular of the true regression function $\mathbb{E}_{P}(Y\,|\,X)$ under $P$, and the equality is shown in the proof of Proposition 1. Finally, we define the posterior regression function $g_{\pi}:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ (with respect to $\pi$) by
$g_{\pi}(x,D_{n})$ $\coloneqq\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl(Y\,|\,X=x,\,\mathcal{D}_{n}=D_{n}\bigr)$ (4)
for $x\in\mathcal{X}$ and $D_{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$. Thus $g_{\pi}(x,D_{n})$ is the posterior mean of $Y$ given $X=x$ and $\mathcal{D}_{n}=D_{n}$, where $(\mathcal{D}_{n},X,Y)\,|\,P\sim P^{\otimes(n+1)}$ and $P\sim\pi$.
The following proposition provides a key upper bound on the ICL excess $\mu$-risk, enabling us to study the adaptivity and distributional robustness of ICL.
Proposition 1.
Let $R>0$, and suppose that for all $P\in\mathcal{P}$, we have $|\mathbb{E}_{P}(Y\,|\,X)|\leq R$ almost surely. Let $\widehat{f}_{T}$ be defined as in (2), let $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$ and let $\mathcal{E}(\widetilde{f}_{T})\coloneqq\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})-\mathcal{R}_{\pi}(g_{\pi})$. Then
$\mathcal{R}^{\mathrm{ICL}}_{\mu}(\widetilde{f}_{T})\leq 4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr].$ (5)
The upper bound in (5) consists of two terms. The first involves both the extent of the distribution shift measured by the $\chi^{2}$-divergence, as well as the expected difference between the $\pi$-risk of $\widetilde{f}_{T}$ and that of the posterior regression function $g_{\pi}$. It is well-known that $g_{\pi}$ minimises the $\pi$-risk $\mathcal{R}_{\pi}$ over all measurable functions, so intuitively, if $\widetilde{f}_{T}$ is a large Transformer pretrained on a sufficient corpus of data, then $\mathcal{E}(\widetilde{f}_{T})$ should be negligible. Indeed, this is proved in our Proposition 3. The second term reflects the proximity of the posterior regression function to the true regression function. In fact, in Sections 4.1 and 4.2, we will control $\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ uniformly for all $P$ in the support of $\mu$. This reveals the interesting feature that only the first term on the right-hand side of (5) depends on the degree of distribution shift, and this term can be made arbitrarily small by increasing the size of the Transformer and the pretraining dataset, so that the second term ultimately determines the rate of convergence.
3 Transformers, universal approximation and learnability
3.1 Transformers
In this section, we formally define the classes of (encoder-only) Transformers that we will consider. For a vector $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$, let $\operatorname{\mathsf{Softmax}}(v)\in\mathbb{R}^{d}$ denote the vector whose $j$th entry is given by $\exp(v_{j})/\sum_{\ell=1}^{d}\exp(v_{\ell})$. Often, we will apply the $\operatorname{\mathsf{Softmax}}$ function to matrices, by which we mean that the $\operatorname{\mathsf{Softmax}}$ function is applied row-wise. Given $d_{\mathrm{model}}\geq d+2$, called the model dimension, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a query $x\in\mathcal{X}$, the input matrix is defined as
$Z_{\mathrm{in}}\equiv Z_{\mathrm{in}}(D_{n},x)\coloneqq\begin{pmatrix}x_{1}^{\top}&y_{1}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ x_{n}^{\top}&y_{n}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ x^{\top}&0&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}},$ (6)
where the last column of $Z_{\mathrm{in}}$ is our positional encoding used to identify the query, and we pad each row with zeros to make them $d_{\mathrm{model}}$-dimensional. A Transformer iteratively applies attention layers and feed forward network (FFN) layers to the input matrix $Z_{\mathrm{in}}$. Letting $N\in\mathbb{N}$ denote an upper bound on the context length, our universal approximation theory (Theorem 2) holds uniformly for all $n\in[N]$, so it is convenient to consider Transformers as functions from [^2] The space $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ is equipped with the disjoint union topology and its corresponding Borel $\sigma$-algebra. $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ to $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$. We now define attention layers.
Definition 1 (Attention layer).
Let $H\in\mathbb{N}$ and for $h\in[H]$, let $Q_{h},K_{h},V_{h}\in\mathbb{R}^{d_{\mathrm{model}}\times d_{\mathrm{model}}}$. The attention layer with $H$ heads and parameters $\theta_{\mathrm{attn}}\coloneqq(Q_{h},K_{h},V_{h})_{h=1}^{H}$ is the function $\mathsf{Attn}_{\theta_{\mathrm{attn}}}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ such that for $n\in[N]$ and $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$,
$\operatorname{\mathsf{Attn}}_{\theta_{\mathrm{attn}}}(Z)$ $\coloneqq Z+\sum_{h=1}^{H}\operatorname{\mathsf{Softmax}}\biggl(\frac{ZQ_{h}K_{h}^{\top}Z^{\top}}{\sqrt{d_{\mathrm{model}}}}\biggr)ZV_{h}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}.$
Definition 2 (FFN layer).
Let $d_{\mathrm{ffn}}\in\mathbb{N}$, $W_{1}\in\mathbb{R}^{d_{\mathrm{model}}\times d_{\mathrm{ffn}}}$, $W_{2}\in\mathbb{R}^{d_{\mathrm{ffn}}\times d_{\mathrm{model}}}$, $v\in\mathbb{R}^{d_{\mathrm{ffn}}}$ and let $\rho:\mathbb{R}\rightarrow\mathbb{R}$. The FFN layer with parameters $\theta_{\mathrm{ffn}}\coloneqq(W_{1},W_{2},v)$ (and activation function $\rho$) is the function $\mathsf{FFN}_{\theta_{\mathrm{ffn}}}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ such that for $n\in[N]$ and $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$,
$\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(Z)\coloneqq Z+\rho(ZW_{1}+\bm{1}_{n+1}v^{\top})W_{2}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}},$
where $\rho$ is applied entrywise.
Now, a Transformer block is a composition of an attention layer and an FFN layer.
Definition 3 (Transformer block).
Let $\theta_{\mathrm{attn}}$ and $\theta_{\mathrm{ffn}}$ be as in Definitions 1 and 2 respectively. The Transformer block with parameters $(\theta_{\mathrm{attn}},\theta_{\mathrm{ffn}})$ is the function $\mathsf{TFBlock}_{(\theta_{\mathrm{attn}},\theta_{\mathrm{ffn}})}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ given by
$\mathsf{TFBlock}_{(\theta_{\mathrm{attn}},\theta_{\mathrm{ffn}})}\coloneqq\mathsf{FFN}_{\theta_{\mathrm{ffn}}}\circ\operatorname{\mathsf{Attn}}_{\theta_{\mathrm{attn}}}.$
We say that the Transformer block $\mathsf{TFBlock}_{(\theta_{\mathrm{attn}},\theta_{\mathrm{ffn}})}$ has $H$ heads, model dimension $d_{\mathrm{model}}$ and FFN width $d_{\mathrm{ffn}}$.
Finally, a Transformer is a composition of Transformer blocks.
Definition 4 (Transformers).
We define $\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$ to be the set of all Transformers $\mathsf{TF}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ of the form
$\mathsf{TF}=\mathsf{TFBlock}^{(L)}\circ\mathsf{TFBlock}^{(L-1)}\circ\cdots\circ\mathsf{TFBlock}^{(1)},$
where $\mathsf{TFBlock}^{(\ell)}\equiv\mathsf{TFBlock}_{(\theta_{\mathrm{attn}}^{(\ell)},\theta_{\mathrm{ffn}}^{(\ell)})}$ is a Transformer block with $H$ heads, model dimension $d_{\mathrm{model}}$ and FFN width $d_{\mathrm{ffn}}$ for all $\ell\in[L]$.
For a Transformer $\mathsf{TF}\in\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}$ and a query $x$, we will slightly abuse our notation by writing $\mathsf{TF}(x,D_{n})\coloneqq\mathsf{TF}\bigl(Z_{\mathrm{in}}(D_{n},x)\bigr)$ where the input matrix $Z_{\mathrm{in}}(D_{n},x)$ is defined by (6). Finally, we define $\mathsf{Read}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\mathbb{R}$ by $\mathsf{Read}(Z)\coloneqq Z_{n+1,d+1}$ for $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ and $n\in[N]$. Thus, $\mathsf{Read}\circ\mathsf{TF}(x,D_{n})$ can be interpreted as the predicted value of the response at the query $x$ based on examples $D_{n}$.
3.2 Universal approximation and learnability
We now consider pretraining priors induced by randomly drawn regression functions, and prove universal approximation and learnability results for Transformers. Let $\mathcal{G}$ be a measurable space of real-valued, measurable functions on $\mathcal{X}\subseteq\mathbb{R}^{d}$ that are uniformly bounded by $R>0$ and let $\widetilde{\pi}$ be a distribution on $\mathcal{G}$. Let $X$ be a random variable on $\mathcal{X}$ with distribution $P_{X}$, and let $\xi\perp\!\!\!\perp X$ be a zero-mean random variable on $\mathbb{R}$. For $g\in\mathcal{G}$, let $P_{g}$ be the distribution of $(X,Y_{g})$ where $Y_{g}=g(X)+\xi$. Further let $\pi$ be the distribution of the random measure $P_{\widetilde{g}}$, where the randomness comes from $\widetilde{g}\sim\widetilde{\pi}$. Suppose throughout this subsection that there exist $R^{\prime}>0$ and a $\sigma$-finite measure $\nu$ on $\mathbb{R}$ such that for all $g\in\mathcal{G}$ and almost all realisations of $g(X)$, the conditional distribution of $Y_{g}\,|\,g(X)$ is absolutely continuous with respect to $\nu$, with Radon–Nikodym derivative bounded by $R^{\prime}$. For example, if the noise $\xi$ has a $N(0,\sigma^{2})$ distribution, then the conditional distribution of $Y_{g}\,|\,\{g(X)=z\}$ has bounded Lebesgue density given by $y\mapsto\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-(y-z)^{2}/(2\sigma^{2})}$. The following theorem shows that Transformers are universal approximators for the posterior regression function $g_{\pi}$ given by (4).
Theorem 2.
Suppose that the activation function $\rho:\mathbb{R}\rightarrow\mathbb{R}$ in the FFN layers is Borel measurable and there is no polynomial $h$ such that $\rho=h$ Lebesgue almost everywhere. Then for any $\epsilon>0$ and $N\in\mathbb{N}$, there exist $d_{\mathrm{model}}\geq d+2$, $d_{\mathrm{ffn}}\in\mathbb{N}$ and a Transformer $\mathsf{TF}\in\mathcal{F}_{\mathrm{TF}}(3,1,d_{\mathrm{model}},d_{\mathrm{ffn}})$ such that
$\max_{n\in[N]}\mathbb{E}_{P\sim\pi,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\Bigl\{\bigl(\mathsf{Read}\circ\mathsf{TF}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}\leq\epsilon.$
From now on, we will assume that the activation $\rho$ belongs to the set $\{\mathsf{ReLU},\mathsf{GELU},\mathsf{SiLU}\}$, which are commonly used activation functions in modern Transformer models; see (14) for their definitions. The following proposition shows that, when the noise $\xi$ is sub-Gaussian, a large Transformer pretrained with sufficient data can achieve a $\pi$-risk that is arbitrarily close to the $\pi$-risk of $g_{\pi}$. Note that we do not require the input matrix or the parameters of the Transformer to be bounded. We achieve this by showing that the pseudo-dimension of the class is finite using results from a branch of mathematical logic called model theory.
Proposition 3.
Suppose that the noise $\xi$ is sub-Gaussian. For any $n\in\mathbb{N}$ and $\epsilon>0$, there exist $d_{\mathrm{model}}^{\circ}\geq d+2$ and $d_{\mathrm{ffn}}^{\circ}\in\mathbb{N}$ such that the following holds. Suppose $d_{\mathrm{model}}\geq d_{\mathrm{model}}^{\circ}$, $d_{\mathrm{ffn}}\geq d_{\mathrm{ffn}}^{\circ}$, and $\widehat{f}_{T}$ is defined as in (2) with $\mathcal{F}\coloneqq\mathsf{Read}\circ\mathcal{F}_{\mathrm{TF}}(3,1,d_{\mathrm{model}},d_{\mathrm{ffn}})$ and pretraining data generated according to $\pi$. Let $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$. Then, for all $T$ sufficiently large,
$\mathcal{E}(\widetilde{f}_{T})\coloneqq\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})-\mathcal{R}_{\pi}(g_{\pi})\leq\epsilon.$
This proposition confirms formally the intuition provided after Proposition 1 that the first term in the decomposition (5) can be made negligible.
4 Adaptivity and robustness of ICL
In this section, we consider mixture distribution pretraining priors $\pi=\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$ for some finite index set $\mathcal{A}$, and test distribution $\mu$ on $\mathcal{P}$ with $\chi^{2}(\mu,\pi_{\beta})\leq\kappa$ for some $\beta\in\mathcal{A}$, so that draws from $\mu$ contain tasks of difficulty $\beta$ with potential distribution shift relative to $\pi_{\beta}$. We have already seen that Proposition 1 provides a decomposition of the ICL excess $\mu$-risk, and Proposition 3 shows that the first term on the right-hand side of (5) can be made negligible. Our goal here, then, is to control the key quantity in the other term, namely $\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$, uniformly over all $P$ in the support of $\pi_{\beta}$. To this end, we specialise the general posterior contraction theory for nonparametric regression that we provide in Appendix E.2 to Besov regression functions with random smoothness (in Section 4.1) and multi-index models (in Section 4.2).
4.1 In-context learning for Besov functions
For the remainder of this section, we take [^3] Our results in this section extend to any compact domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ by a scaling argument. $\mathcal{X}\coloneqq[0,1]^{d}$, $\mathcal{Y}\coloneqq\mathbb{R}$. Since Besov functions can be characterised by their wavelet decomposition, we will consider regression functions with random wavelet coefficients. Let $L^{2}([0,1]^{d})$ denote the set of square integrable functions on $0,1]^{d}$, let $S\in\mathbb{N}$ and let
$\bigl\{\Phi_{k}:k\in K\bigr\}\cup\bigl\{\Psi_{\ell,\gamma}:\ell\in[{\ell_{0}}:\infty),\,\gamma\in\Gamma_{\ell}\bigr\}$ (7)
denote the tensor product Cohen–Daubechies–Vial (CDV) wavelet basis for $L^{2}([0,1]^{d})$, constructed from $S$-regular and $S$-times continuously differentiable wavelets. The precise definitions of these wavelet functions are not important for us, and we will only use some basic properties of this wavelet basis, summarised in Appendix D.2. Let $c_{0}\in(0,1]$ and $C_{0}\geq 1$. For $\alpha\in(0,S)$, let $\widetilde{\pi}_{\alpha}$ be the distribution of the random function
$\widetilde{g}_{\alpha}\coloneqq\sum_{k\in K}C_{0}2^{-\ell_{0}d/2}a_{k}^{(\alpha)}\Phi_{k}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}C_{0}2^{-\ell(\alpha+d/2)}b_{\ell,\gamma}^{(\alpha)}\Psi_{\ell,\gamma},$ (8)
where $\bigl(a_{k}^{(\alpha)},b_{\ell,\gamma}^{(\alpha)}:k\in K,\ell\in[\ell_{0}:\infty),\gamma\in\Gamma_{\ell}\bigr)$ are independent random variables supported on [1,1][-1,1] with Lebesgue density bounded between $c_{0}/2$ and $c_{0}^{-1}/2$. The scaling of the wavelet coefficients in (8) is chosen to ensure that $\widetilde{g}_{\alpha}$ belongs to the Besov space $B_{\infty,\infty}^{\alpha}([0,1]^{d})$ with smoothness $\alpha$. Indeed, a function $g\coloneqq\sum_{k\in K}a_{k}^{\prime}\Phi_{k}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}b_{\ell,\gamma}^{\prime}\Psi_{\ell,\gamma}$, with deterministic wavelet coefficients, belongs to the Besov ball $B_{\infty,\infty}^{\alpha}([0,1]^{d},C)$ of radius $C>0$ if and only if $2^{\ell_{0}d/2}\max_{k}|a_{k}^{\prime}|+\sup_{\ell,\gamma}2^{\ell(\alpha+d/2)}|b_{\ell,\gamma}^{\prime}|\leq C$; see Appendix D. Thus, writing $\mathcal{G}_{\alpha}\coloneqq\mathrm{supp}(\widetilde{\pi}_{\alpha})$, we have
$B_{\infty,\infty}^{\alpha}([0,1]^{d},C_{0})\subseteq\mathcal{G}_{\alpha}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{d},2C_{0}).$ (9)
We remark that the Besov space $B_{\infty,\infty}^{\alpha}([0,1]^{d})$ coincides with the Hölder space $H^{\alpha}([0,1]^{d})$ for $\alpha\notin\mathbb{N}$, and it contains $H^{\alpha}([0,1]^{d})$ for $\alpha\in\mathbb{N}$; see Appendix D.1 for more details. Moreover, there exists $C_{\alpha,S,d}>0$ such that $\sup_{g\in\mathcal{G}_{\alpha}}\|g\|_{\infty}\leq C_{\alpha,S,d}$; see (54).
Now let $P_{X}$ be a Borel probability distribution on $0,1]^{d}$ with the property that there exists a hypercube $A\subseteq[0,1]^{d}$ such that $P_{X}(A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0})$ for all measurable $A_{0}\subseteq A$. Let $X\sim P_{X}$ and $\xi\sim N(0,\sigma^{2})$ be independent. For a Borel measurable function $g:[0,1]^{d}\to\mathbb{R}$, let $P_{g}$ denote the distribution of $(X,Y_{g})$ where $Y_{g}=g(X)+\xi$. We then define $\pi_{\alpha}$ to be the distribution of the random measure $P_{\widetilde{g}_{\alpha}}$, where the randomness comes from the random regression function $\widetilde{g}_{\alpha}\sim\widetilde{\pi}_{\alpha}$, and let $\mathcal{P}_{\alpha}\coloneqq\mathrm{supp}(\pi_{\alpha})$. Letting $\mathcal{A}\subseteq(0,S)$ be a finite set and letting $(\lambda_{\alpha})_{\alpha\in\mathcal{A}}$ be a sequence of positive weights such that $\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}=1$, our pretraining prior $\pi$ is defined to be the mixture distribution
$\pi\coloneqq\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}.$ (10)
Thus $\pi$ consists of random regression tasks where the smoothness of the regression functions is randomly drawn from $\mathcal{A}$ with probabilities given by $(\lambda_{\alpha})_{\alpha\in\mathcal{A}}$. Finally, we define $\mathcal{P}\coloneqq\bigcup_{\alpha\in\mathcal{A}}\mathcal{P}_{\alpha}$ and $\mathcal{G}\coloneqq\bigcup_{\alpha\in\mathcal{A}}\mathcal{G}_{\alpha}$. By (9), the set $\mathcal{G}$ contains a union of Besov balls with smoothness parameters in $\mathcal{A}$.
The following proposition provides the rate of convergence of the posterior regression, when the true regression belongs to $\mathcal{G}_{\beta}$ for some $\beta\in\mathcal{A}$.
Proposition 4.
For any $\beta\in\mathcal{A}$ and $g^{\circ}\in\mathcal{G}_{\beta}$, write $P_{0}\equiv P_{g^{\circ}}$, let $\pi$ be the prior distribution defined in (10) and let $g_{\pi}$ denote the posterior regression function with respect to $\pi$. Then there exists $C>0$, not depending on $n$, such that
$\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+d)}$
for all $n\in\mathbb{N}$.
Combining Propositions 13 and 4, we are now ready to state the main theorem of this section.
Theorem 5.
For any $n\in\mathbb{N}$ and $\kappa>0$, there exist $d_{\mathrm{model}}^{\circ}\geq d+2$ and $d_{\mathrm{ffn}}^{\circ}\in\mathbb{N}$ such that the following holds. Suppose $d_{\mathrm{model}}\geq d_{\mathrm{model}}^{\circ}$, $d_{\mathrm{ffn}}\geq d_{\mathrm{ffn}}^{\circ}$, $\widehat{f}_{T}$ is defined as in (2) with $\mathcal{F}\coloneqq\mathsf{Read}\circ\mathcal{F}_{\mathrm{TF}}(3,1,d_{\mathrm{model}},d_{\mathrm{ffn}})$ and pretraining prior $\pi$ defined in (10). Let $R\coloneqq\sup_{g\in\mathcal{G}}\|g\|_{\infty}$ and $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$. Then, for all $T$ sufficiently large and $\beta\in\mathcal{A}$,
$\sup_{\mu\,:\,\chi^{2}(\mu,\pi_{\beta})\leq\kappa}\mathcal{R}^{\mathrm{ICL}}_{\mu}(\widetilde{f}_{T})\leq Cn^{-2\beta/(2\beta+d)},$
where $C>0$ does not depend on $n$ and $\kappa$.
The Transformer $\widetilde{f}_{T}$ in Theorem 5 is pretrained on the mixture distribution $\pi=\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$ with different smoothness levels, whereas the test distribution $\mu$ consists of random regression functions of a fixed smoothness level $\beta$, with a distribution shift such that $\chi^{2}(\mu,\pi_{\beta})\leq\kappa$. When we evaluate the performance of $\widetilde{f}_{T}$ under the test distribution $\mu$, Theorem 5 shows that the ICL excess $\mu$-risk of $\widetilde{f}_{T}$ is adaptive to the unknown smoothness $\beta$, in the sense that its rate of convergence $n^{-2\beta/(2\beta+d)}$ is optimal even if $\beta$ were known; see Theorem 8. Moreover, this risk bound holds uniformly over test distributions in a $\chi^{2}$-divergence ball around $\pi_{\beta}$. Thus, although the Transformer is pretrained on tasks of different levels of difficulties, it is able to achieve faster (and optimal) rates of convergence on easier tasks (without knowing the difficulty of the tasks) and is robust to distribution shift. Finally, we remark that $d_{\mathrm{model}}^{\circ},d_{\mathrm{ffn}}^{\circ}$ and $T$ should increase with $\kappa$ in order to keep $C$ independent of $\kappa$. Thus, for a Transformer with fixed architecture, fine tuning for downstream tasks may still be beneficial in cases where (i) the Transformer is not large enough, (ii) the amount of pretraining data is not sufficient, or (iii) there is significant distribution shift ($\kappa$ is large).
4.2 In-context learning for multi-index models
To avoid the curse of dimensionality in the previous section, we will now consider multi-index models with [^4] Again, our results in this section extend to any compact domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ by a scaling argument. $\mathcal{X}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}\eqqcolon\mathbb{B}^{d}$ and $\mathcal{Y}=\mathbb{R}$. For $p\in[d]$, let $V_{p}(\mathbb{R}^{d})\coloneqq\bigl\{U\in\mathbb{R}^{d\times p}:U^{\top}U=I_{p}\bigr\}$ denote the set of all projection matrices from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ (also called the Stiefel manifold). We will consider regression functions on $\mathbb{B}^{d}$ of the form $x\mapsto g\bigl((U^{\top}x+\bm{1}_{p})/2\bigr)$, where $U\in V_{p}(\mathbb{R}^{d})$ for some effective dimension $p\in[d]$ and Borel measurable $g:[0,1]^{p}\to\mathbb{R}$. Note that we translate and scale $U^{\top}x$ to ensure that it belongs to $0,1]^{p}$.
For $\alpha\in(0,S)$ and $p\in[d]$, let $\widetilde{g}_{\alpha}^{(p)}:[0,1]^{p}\to\mathbb{R}$ be the random function (with smoothness $\alpha$) defined as in (8), but with $d$ replaced by $p$ in the wavelet coefficients, and using the tensor product CDV wavelet basis for $L^{2}([0,1]^{p})$ instead of $L^{2}([0,1]^{d})$. Let $\widetilde{\pi}_{\alpha,p}$ be the distribution of the random function $\widetilde{g}_{\alpha,p}:\mathbb{B}^{d}\to\mathbb{R}$ defined by
$\widetilde{g}_{\alpha,p}(x)\coloneqq\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr),$ (11)
where $U^{(p)}$ is uniformly distributed [^5] One way to generate such a random matrix is to set $U^{(p)}\coloneqq Z(Z^{\top}Z)^{-1/2}$, where $Z$ is a $d\times p$ random matrix with independent $N(0,1)$ entries (e.g. Chikuse, 2003, Theorem 2.2.1(iii)). on the Stiefel manifold $V_{p}(\mathbb{R}^{d})$, independently of $\widetilde{g}_{\alpha}^{(p)}$. We write $\mathcal{G}_{\alpha,p}$ for the support of $\widetilde{\pi}_{\alpha,p}$.
Next, let $P_{X}$ be a distribution on $\mathbb{B}^{d}$ with the property that there exist $c_{0}>0$ and a non-empty Euclidean ball $A\subseteq\mathbb{B}^{d}$ such that $P_{X}(A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0})$ for all measurable $A_{0}\subseteq A$. We then define $\pi_{\alpha,p}$ to be the distribution of the random measure $P_{\widetilde{g}_{\alpha,p}}$, where the randomness comes from the random regression function $\widetilde{g}_{\alpha,p}\sim\widetilde{\pi}_{\alpha,p}$, and let $\mathcal{P}_{\alpha,p}\coloneqq\mathrm{supp}(\pi_{\alpha,p})$. For a finite set $\mathcal{A}^{\prime}\subseteq(0,S)\times[d]$ and a sequence of positive weights $(\lambda_{\alpha,p})_{(\alpha,p)\in\mathcal{A}^{\prime}}$ such that $\sum_{(\alpha,p)\in\mathcal{A}^{\prime}}\lambda_{\alpha,p}=1$, our pretraining prior $\pi$ for multi-index models is defined to be the mixture distribution
$\pi\coloneqq\sum_{(\alpha,p)\in\mathcal{A}^{\prime}}\lambda_{\alpha,p}\pi_{\alpha,p}.$ (12)
Thus $\pi$ generates random multi-index regression tasks, where the smoothness and effective dimensions of the regression functions are randomly drawn from $\mathcal{A}^{\prime}$ with probabilities given by $(\lambda_{\alpha,p})_{(\alpha,p)\in\mathcal{A}^{\prime}}$. Finally, we define $\mathcal{P}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{P}_{\alpha,p}$ and $\mathcal{G}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{G}_{\alpha,p}$. Proposition 6 is the analogue for multi-index models of Proposition 4.
Proposition 6.
For any $(\beta,r)\in\mathcal{A}^{\prime}$ and $g^{\circ}\in\mathcal{G}_{\beta,r}$, write $P_{0}\equiv P_{g^{\circ}}$, let $\pi$ be the prior distribution defined in (12) and let $g_{\pi}$ denote the posterior regression function with respect to $\pi$. Then there exists $C>0$, not depending on $n$, such that
$\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+r)}$
for all $n\in\mathbb{N}$.
Combining Propositions 13 and 6, we can now conclude that a pretrained Transformer is adaptive to the effective dimension of the multi-index models, as well as adapting to the smoothness of the regression function and being robust to distribution shift.
Theorem 7.
For any $n\in\mathbb{N}$ and $\kappa>0$, there exist $d_{\mathrm{model}}^{\circ}\geq d+2$ and $d_{\mathrm{ffn}}^{\circ}\in\mathbb{N}$ such that the following holds. Suppose $d_{\mathrm{model}}\geq d_{\mathrm{model}}^{\circ}$, $d_{\mathrm{ffn}}\geq d_{\mathrm{ffn}}^{\circ}$, $\widehat{f}_{T}$ is defined as in (2) with $\mathcal{F}\coloneqq\mathsf{Read}\circ\mathcal{F}_{\mathrm{TF}}(3,1,d_{\mathrm{model}},d_{\mathrm{ffn}})$ and pretraining prior $\pi$ defined in (12). Let $R\coloneqq\sup_{g\in\mathcal{G}}\|g\|_{\infty}$ and $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$. Then, for all $T$ sufficiently large and $(\beta,r)\in\mathcal{A}^{\prime}$,
$\sup_{\mu\,:\,\chi^{2}(\mu,\pi_{\beta,r})\leq\kappa}\mathcal{R}^{\mathrm{ICL}}_{\mu}(\widetilde{f}_{T})\leq Cn^{-2\beta/(2\beta+r)},$
where $C>0$ does not depend on $n$ and $\kappa$.
4.3 Lower bounds on the ICL excess risk
Our final theoretical contribution is to provide lower bounds on the ICL excess risk. Prior work has sought to establish the optimality of ICL algorithms via lower bounds on the minimax risk
$\inf_{\widehat{g}_{n}\in\widehat{\mathcal{G}}_{n}}\sup_{P\in\mathrm{supp}(\mu)}\mathbb{E}_{P}\bigl[\bigl\{\widehat{g}_{n}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$ (13)
where $\widehat{\mathcal{G}}_{n}$ denotes the set of all measurable functions from $\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}$ to $\mathcal{Y}$. Although the minimax risk is often regarded as the gold standard in traditional statistical learning problems, the ICL excess $\mu$-risk $\mathcal{R}_{\mu}^{\mathrm{ICL}}$ is an expectation over $P\sim\mu$, not a worst-case risk over $P\in\mathrm{supp}(\mu)$. Thus lower bounds on (13) do not provide lower bounds for the ICL excess $\mu$-risk. Theorem 8 below, however, does indeed provide a lower bound on the ICL excess $\mu$-risk, and in a strong sense: even an estimator that has knowledge of the test distribution $\mu$ cannot achieve a better rate of convergence than the pretrained Transformers in Theorems 5 and 7.
Theorem 8.
(a) Under the setting of Section 4.1, let $\beta\in\mathcal{A}$ and let $\mu\ll\pi_{\beta}$ be such that $\sup_{P\in\mathcal{P}_{\beta}}\frac{\mathrm{d}\mu}{\mathrm{d}\pi_{\beta}}(P)<\infty$. Then there exists $c>0$, not depending on $n$, such that
$\inf_{\widehat{g}_{n}\in\widehat{\mathcal{G}}_{n}}\mathcal{R}_{\mu}^{\mathrm{ICL}}(\widehat{g}_{n})\geq cn^{-2\beta/(2\beta+d)}.$
(b) In the setting of Section 4.2, let $(\beta,r)\in\mathcal{A}^{\prime}$ and let $\mu\ll\pi_{\beta,r}$ be such that $\sup_{P\in\mathcal{P}_{\beta,r}}\frac{\mathrm{d}\mu}{\mathrm{d}\pi_{\beta,r}}(P)<\infty$. Then there exists $c^{\prime}>0$, not depending on $n$, such that
$\inf_{\widehat{g}_{n}\in\widehat{\mathcal{G}}_{n}}\mathcal{R}_{\mu}^{\mathrm{ICL}}(\widehat{g}_{n})\geq c^{\prime}n^{-2\beta/(2\beta+r)}.$
We remark that, if $\widetilde{f}_{T}$ is defined as in Theorems 5 or 7, then conditional on the pretraining data, we have $\widetilde{f}_{T}\in\widehat{\mathcal{G}}_{n}$.
Acknowledgement: The research of the first and third author was supported by RJS’s European Research Council Advanced Grant 101019498.
Appendix A Notation for the appendix
Let $(\mathcal{P},d)$ be a metric space, $\epsilon>0$ and $\mathcal{P}^{\prime}\subseteq\mathcal{P}$, we write $\mathcal{N}(\epsilon,\mathcal{P}^{\prime},d)$ for the minimum number of closed $\epsilon$-balls in $d$ needed to cover $\mathcal{P}^{\prime}$. For $m,n\in\mathbb{N}$, let $I_{n}\in\mathbb{R}^{n\times n}$ be the identity matrix; for a matrix $A\in\mathbb{R}^{n\times m}$, define the operator norm of $A$ by $\|A\|_{\mathrm{op}}\coloneqq\sup_{x\in\mathbb{R}^{m}:\|x\|_{2}=1}\|Ax\|_{2}$ and define its entrywise $\ell_{\infty}$-norm by $\|A\|_{\infty}\coloneqq\max_{i\in[n],j\in[m]}|A_{ij}|$.
Let $P$ and $Q$ be probability measures on $\mathcal{Z}$, and let $\nu$ be a measure on $\mathcal{Z}$ such that $P\ll\nu$ and $Q\ll\nu$ (such a measure always exists, e.g. $\nu=P+Q$). Write $p=\frac{\mathrm{d}P}{\mathrm{d}\nu}$ and $q=\frac{\mathrm{d}Q}{\mathrm{d}\nu}$, then the Hellinger distance between $P$ and $Q$ is defined as
$d_{\mathrm{H}}(P,Q)\coloneqq\Bigl(\int_{\mathcal{Z}}\bigl(\sqrt{p(z)}-\sqrt{q(z)}\bigr)^{2}\,\mathrm{d}\nu(z)\Bigr)^{1/2}.$
The definition is independent of the choice of the dominating measure $\nu$, and $d_{\mathrm{H}}$ is a metric on the space of distributions on $\mathcal{Z}$. The Kullback–Leibler divergence from $Q$ to $P$ is defined as
$\mathrm{KL}(P,Q)\coloneqq\mathbb{E}_{P}\biggl(\log\frac{\mathrm{d}P}{\mathrm{d}Q}(Z)\biggr)=\int_{\mathcal{Z}}\log\frac{\mathrm{d}P}{\mathrm{d}Q}(z)\,\mathrm{d}P(z)$
if $P\ll Q$, and infinity otherwise; this is the $P$-expectation of the log-likelihood ratio between $P$ and $Q$. We further define
$\mathrm{V}_{2}(P,Q)\coloneqq\mathrm{Var}_{P}\biggl(\log\frac{\mathrm{d}P}{\mathrm{d}Q}(Z)\biggr)=\int_{\mathcal{Z}}\biggl(\log\frac{\mathrm{d}P}{\mathrm{d}Q}(z)-\mathrm{KL}(P,Q)\biggr)^{2}\,\mathrm{d}P(z)$
if $P\ll Q$, and infinity otherwise; this is the $P$-variance of the log-likelihood ratio between $P$ and $Q$.
For $d\in\mathbb{N}$, $p\in[1,\infty)$, a Borel measurable function $f:\mathbb{R}^{d}\to\mathbb{R}$ and a Borel measure $\mu$ on $\mathbb{R}^{d}$, we define
$\|f\|_{L^{p}(\mu)}\coloneqq\biggl(\int_{\mathbb{R}^{d}}|f(x)|^{p}\,\mathrm{d}\mu(x)\biggr)^{1/p}.$
and write $L^{p}(\mu)$ for the set of functions $f:\mathbb{R}^{d}\to\mathbb{R}$ with $\|f\|_{L^{p}(\mu)}<\infty$. If the domain of $f$ is restricted to $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mu$ is the Lebesgue measure restricted to a Borel set $\mathcal{X}$, then we write $\|f\|_{L^{p}(\mathcal{X})}\equiv\|f\|_{L^{p}(\mu)}$ and $L^{p}(\mathcal{X})\equiv L^{p}(\mu)$; when $\mathcal{X}$ is clear from context, we further write $\|f\|_{L^{p}}\equiv\|f\|_{L^{p}(\mathcal{X})}$.
For a zero-mean random variable $X$ on $\mathbb{R}$, we say that $X$ is sub-Gamma in the right tail with variance parameter $\sigma^{2}>0$ and scale parameter $c>0$ if for all $\lambda\in[0,1/c)$, we have
$\mathbb{E}e^{\lambda X}\leq\exp\biggl(\frac{\sigma^{2}\lambda^{2}}{2(1-c\lambda)}\biggr).$
Similarly, $X$ is sub-Gamma in the left tail if X-X is sub-Gamma in the right tail. And we say $X$ is sub-Gamma in both tails is it is both sub-Gamma in the right tail and sub-Gamma in the left tail.
Finally, we define three commonly used activation functions in Transformers. For $x\in\mathbb{R}$, define
$\mathsf{ReLU}(x)\coloneqq\max\{0,x\},\quad\mathsf{GELU}(x)\coloneqq x\Phi(x)\quad\text{and}\quad\mathsf{SiLU}(x)\coloneqq\frac{x}{1+e^{-x}},$ (14)
where $\Phi(x)\coloneqq\int_{-\infty}^{x}\frac{1}{\sqrt{2\pi}}e^{-t^{2}/2}\,\mathrm{d}t$ is the distribution function of standard Gaussian distribution.
Appendix B Proof of Proposition 1
Proof of Proposition 1.
We may assume without loss of generality that $\mu\ll\pi$ since otherwise, the right-hand side of (5) is infinite. Note that
$\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})$ $-\mathcal{R}_{\pi}(g_{\pi})=\mathbb{E}\,\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl[\bigl\{Y-\widetilde{f}_{T}(X,\mathcal{D}_{n})\bigr\}^{2}-\bigl\{Y-g_{\pi}(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]$
$=\mathbb{E}\,\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl\{-2Y\widetilde{f}_{T}(X,\mathcal{D}_{n})+\widetilde{f}_{T}^{2}(X,\mathcal{D}_{n})+2Yg_{\pi}(X,\mathcal{D}_{n})-g_{\pi}^{2}(X,\mathcal{D}_{n})\bigr\},$
where the outer expectation is taken over the randomness of $\widetilde{f}_{T}$. Moreover,
$\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl\{Y\widetilde{f}_{T}(X,\mathcal{D}_{n})\bigr\}$ $=\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl[\mathbb{E}_{P\sim\pi,\,(\mathcal{D}_{n},X,Y)\sim P^{\otimes(n+1)}}\bigl\{Y\widetilde{f}_{T}(X,\mathcal{D}_{n})\,|\,X,\mathcal{D}_{n}\bigr\}\bigr]$
$=\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P\sim\pi,\,(\mathcal{D}_{n},X,Y)\sim P^{\otimes(n+1)}}\bigl(Y\,|\,X,\mathcal{D}_{n}\bigr)\bigr\}$
$=\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})g_{\pi}(X,\mathcal{D}_{n})\bigr\},$
and similarly
$\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl\{Yg_{\pi}(X,\mathcal{D}_{n})\bigr\}$ $=\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl[\mathbb{E}_{P\sim\pi,\,(\mathcal{D}_{n},X,Y)\sim P^{\otimes(n+1)}}\bigl\{Yg_{\pi}(X,\mathcal{D}_{n})\,|\,X,\mathcal{D}_{n}\bigr\}\bigr]$
$=\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl\{g_{\pi}^{2}(X,\mathcal{D}_{n})\bigr\}.$
Therefore,
$\mathcal{E}(\widetilde{f}_{T})$ $=\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})-\mathcal{R}_{\pi}(g_{\pi})=\mathbb{E}\,\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl\{-2g_{\pi}(X,\mathcal{D}_{n})\widetilde{f}_{T}(X,\mathcal{D}_{n})+\widetilde{f}_{T}^{2}(X,\mathcal{D}_{n})+g_{\pi}^{2}(X,\mathcal{D}_{n})\bigr\}$
$=\mathbb{E}\,\mathbb{E}_{\pi}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr\}^{2}\bigr].$
By (4), we have $|g_{\pi}(X,\mathcal{D}_{n})|\leq R$ almost surely. Now, for every fixed $P\in\mathcal{P}$, define $Z_{P}\coloneqq\mathbb{E}\,\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]$, so that by Fubini’s theorem and the Cauchy–Schwarz inequality,
$\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]=\mathbb{E}_{\mu}(Z_{P})=\mathbb{E}_{\pi}\biggl(Z_{P}\cdot\frac{\mathrm{d}\mu}{\mathrm{d}\pi}(P)\biggr)$
$\leq\sqrt{\mathbb{E}_{\pi}(Z_{P}^{2})\cdot\mathbb{E}_{\pi}\biggl\{\biggl(\frac{\mathrm{d}\mu}{\mathrm{d}\pi}(P)\biggr)^{2}\biggr\}}\leq\sqrt{4R^{2}\mathbb{E}_{\pi}(Z_{P})\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}}=2R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}.$ (15)
Hence,
$\mathbb{E}\mathcal{R}_{\mu}(\widetilde{f}_{T})$ $-\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
$=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{-2Y\widetilde{f}_{T}(X,\mathcal{D}_{n})+\widetilde{f}_{T}^{2}(X,\mathcal{D}_{n})+2Y\mathbb{E}_{P}(Y\,|\,X)-\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}$
$\hskip-7.00032pt\overset{(i)}{=}\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
$\leq 2\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
$\hskip-8.72289pt\overset{(ii)}{\leq}4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr].$
Here, $(ii)$ follows from (15), and $(i)$ follows since
$\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{Y\widetilde{f}_{T}(X,\mathcal{D}_{n})\bigr\}$ $=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\mathbb{E}_{P}\bigl\{Y\widetilde{f}_{T}(X,\mathcal{D}_{n})\,|\,X,\mathcal{D}_{n}\bigr\}\bigr]$
$=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P}(Y\,|\,X,\mathcal{D}_{n})\bigr\}$
$=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P}(Y\,|\,X)\bigr\},$
and similarly,
$\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\bigr\}=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\,|\,X\bigr\}\bigr]=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}.$
This proves the claim. ∎
Appendix C Proofs for Section 3.2
C.1 Proof of Theorem 2
We work under the assumptions of Section 3.2 and we assume that $\rho$ is not a polynomial. For $g\in\mathcal{G}$, $(X,Y_{g})\sim P_{g}$ and $z\in\mathbb{R}$, let $y\mapsto f_{\xi}(y,z)$ denote the conditional density of $Y_{g}\,|\,\{g(X)=z\}$ with respect to the dominating measure $\nu$. Further let $\bar{f}_{\xi}\coloneqq f_{\xi}/R^{\prime}$. By Bayes’ theorem and (41), the posterior regression function $g_{\pi}$ can be written as
$g_{\pi}(x,D_{n})=\frac{\int_{\mathcal{G}}g(x)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i}-g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)}{\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i}-g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)}.$ (16)
Throughout this section, we assume that $\bigl((X_{1},Y_{1}),\ldots,(X_{N},Y_{N}),(X,Y)\bigr)\,|\,P\sim P^{\otimes(N+1)}$ with $P\sim\pi$. For $n\in[N]$ and $d_{\mathrm{model}}\geq d+2$, define
$Z_{\mathrm{in}}^{(n,d_{\mathrm{model}})}\coloneqq\begin{pmatrix}X_{1}^{\top}&Y_{1}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ X_{n}^{\top}&Y_{n}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ X^{\top}&0&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}.$ (17)
We first write down the rows of attention layers and FFN layers before we start. Consider the notation in Definition 1, if $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ has rows $Z_{1}^{\top},\ldots,Z_{n+1}^{\top}\in\mathbb{R}^{1\times d_{\mathrm{model}}}$, then the $i$th row $\operatorname{\mathsf{Attn}}_{\theta_{\mathrm{attn}}}(Z)]_{i,:}$ of $\operatorname{\mathsf{Attn}}_{\theta_{\mathrm{attn}}}(Z)$ can be written as
$[\operatorname{\mathsf{Attn}}_{\theta_{\mathrm{attn}}}(Z)]_{i,:}=Z_{i}^{\top}+\sum_{h=1}^{H}\sum_{j=1}^{n+1}Z_{j}^{\top}V_{h}\cdot\frac{\exp(Z_{i}^{\top}Q_{h}K_{h}^{\top}Z_{j}/\sqrt{d_{\mathrm{model}}})}{\sum_{\ell=1}^{n+1}\exp(Z_{i}^{\top}Q_{h}K_{h}^{\top}Z_{\ell}/\sqrt{d_{\mathrm{model}}})}\in\mathbb{R}^{1\times d_{\mathrm{model}}}.$
Next, consider the notation in Definition 2, if $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ has rows $Z_{1}^{\top},\ldots,Z_{n+1}^{\top}\in\mathbb{R}^{1\times d_{\mathrm{model}}}$, then the $i$th row $\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(Z)]_{i,:}$ of $\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(Z)$ can be written as
$[\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(Z)]_{i,:}=\bigl(Z_{i}+W_{2}^{\top}\rho(W_{1}^{\top}Z_{i}+v)\bigr)^{\top}.$
Thus the FFN layer applies a single-hidden layer neural network to each row of $Z$.
Lemma 9.
For any $\epsilon,\delta>0$, there exist $M\in\mathbb{N}$ and $g_{1},\ldots,g_{M}\in\mathcal{G}$ such that with probability at least $1-\delta$, we have
$\max_{n\in[N]}\biggl|\frac{1}{M}\sum_{m=1}^{M}g_{m}(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)-\int_{\mathcal{G}}g(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|\leq\epsilon$
and
$\max_{n\in[N]}\biggl|\frac{1}{M}\sum_{m=1}^{M}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)-\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|\leq\epsilon.$
Proof.
Let $M\coloneqq 2\epsilon^{-2}(R^{2}\vee 1)\log(2N/\delta)$ and let $\widetilde{g}_{1},\ldots,\widetilde{g}_{M}\overset{\mathrm{iid}}{\sim}\widetilde{\pi}$ be independent of $(X_{i},Y_{i})_{i=1}^{N}$ and $(X,Y)$. For $n\in[N]$, define events
$A_{n}\coloneqq\Biggl\{\biggl|\frac{1}{M}\sum_{m=1}^{M}\widetilde{g}_{m}(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},\widetilde{g}_{m}(X_{i})\bigr)-\int_{\mathcal{G}}g(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|\leq\epsilon\Biggr\}$
and
$B_{n}\coloneqq\Biggl\{\biggl|\frac{1}{M}\sum_{m=1}^{M}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},\widetilde{g}_{m}(X_{i})\bigr)-\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|\leq\epsilon\Biggr\}.$
We have by Hoeffding’s inequality that
$\mathbb{P}(A_{n}^{\mathrm{c}}\,|\,X_{1},Y_{1},\ldots,X_{n},Y_{n},X)\leq\frac{\delta}{2N}\quad\text{and}\quad\mathbb{P}(B_{n}^{\mathrm{c}}\,|\,X_{1},Y_{1},\ldots,X_{n},Y_{n},X)\leq\frac{\delta}{2N}.$
Then by a union bound,
$\mathbb{E}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\bigg|\,\widetilde{g}_{1},\ldots,\widetilde{g}_{M}\biggr)\biggr\}=\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\biggr)$
$\hskip 113.81102pt=\mathbb{E}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\bigg|\,X_{1},Y_{1},\ldots,X_{N},Y_{N},X,Y\biggr)\biggr\}\leq\delta.$
Therefore,
$\mathbb{P}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\biggm|\,\widetilde{g}_{1},\ldots,\widetilde{g}_{M}\biggr)\leq\delta\biggr\}>0.$
This proves the claim. ∎
Lemma 10.
For all $n\in[N]$, we have
$\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)>0$
almost surely.
Proof.
The density (with respect to $\nu$) of $(Y_{1},\ldots,Y_{n})$ conditional on $(X_{1},\ldots,X_{n})$ is given by
$(y_{1},\ldots,y_{n})\mapsto\int_{\mathcal{G}}\prod_{i=1}^{n}f_{\xi}\bigl(y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g).$
Hence, by definition,
$\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\biggr)$
$\hskip 113.81102pt=\mathbb{E}\biggl\{\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}f_{\xi}\bigl(y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\,\biggm|\,X_{1},\ldots,X_{n}\biggr)\biggr\}=0.$
This holds for all $n\in[N]$, which proves the claim. ∎
Lemma 11.
Let $d_{1},d_{2}\in\mathbb{N}$, let $f:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$ be Borel measurable and let $\mu$ be a Borel probability measure on $\mathbb{R}^{d_{1}}$. Then for every $\epsilon>0$ and $\delta\in(0,1)$, there exist $d^{\prime}\in\mathbb{N}$, $W_{1}\in\mathbb{R}^{d_{1}\times d^{\prime}}$, $W_{2}\in\mathbb{R}^{d^{\prime}\times d_{2}}$ and $v\in\mathbb{R}^{d^{\prime}}$ such that
$\mu\Bigl(\bigl\{x\in\mathbb{R}^{d_{1}}:\bigl\|f(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr\|_{\infty}\leq\epsilon\bigr\}\Bigr)\geq 1-\delta.$
Proof.
We first prove the claim for $d_{2}=1$. Let $K_{0}\subseteq\mathbb{R}^{d_{1}}$ be a compact set such that $\mu(K_{0})\geq 1-\delta/3$. By Lusin’s theorem (Folland, 1999, Theorems 7.8 and 7.10), there exists a (compactly supported) continuous function $h:\mathbb{R}^{d_{1}}\to\mathbb{R}$ such that $\mu\bigl(\{x\in K_{0}:f(x)\neq h(x)\}\bigr)\leq\delta/3$. Since one-hidden layer neural networks with non-polynomial activation are dense in the set of continuous functions on compact domains equipped with the uniform norm (Leshno et al., 1993, Theorem 1), there exist $d^{\prime}\in\mathbb{N}$, $W_{1}\in\mathbb{R}^{d_{1}\times d^{\prime}}$, $W_{2}\in\mathbb{R}^{d^{\prime}\times d_{2}}$ and $v\in\mathbb{R}^{d^{\prime}}$ such that for all $x\in K_{0}$,
$\bigl|h(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\leq\frac{\delta\epsilon}{3}.$
It follows by Markov’s inequality that
$\mu\Bigl(\bigl\{x\in\mathbb{R}^{d_{1}}:\bigl|f(x)$ $-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)$
$\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|f(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}$
$\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|h(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}+\frac{\delta}{3}\leq\delta.$
This proves the claim when $d_{2}=1$. For $d_{2}>1$, we may approximate the component functions in parallel and apply a union bound to deduce the final result. ∎
Lemma 12.
Let $d_{\mathrm{model}}\geq d+2$ and let $f_{1}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d_{\mathrm{model}}}$ and $f_{2}:\mathbb{R}^{d}\to\mathbb{R}^{d_{\mathrm{model}}}$ be Borel measurable. For $n\in[N]$, let $Z_{\mathrm{in}}^{(n,d_{\mathrm{model}})}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ be defined as in (17) and let
$U^{(n)}\coloneqq\begin{pmatrix}f_{1}^{\top}(X_{1},Y_{1})\\ \vdots\\ f_{1}^{\top}(X_{n},Y_{n})\\ f_{2}^{\top}(X)\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}.$
Then for $\delta\in(0,1)$ and $\epsilon>0$, there exists an FFN layer $\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ such that with probability at least $1-\delta$,
$\max_{n\in[N]}\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(Z_{\mathrm{in}}^{(n,d_{\mathrm{model}})})-U^{(n)}\bigr\|_{\infty}\leq\epsilon.$
Proof.
Define a Borel measurable function $f:\mathbb{R}^{d_{\mathrm{model}}}\to\mathbb{R}^{d_{\mathrm{model}}}$ by $f(x,y,\bm{0}_{d_{\mathrm{model}}-d-2},0)\coloneqq f_{1}(x,y)-(x,y,\bm{0}_{d_{\mathrm{model}}-d-2},0)$ and $f(x,0,\bm{0}_{d_{\mathrm{model}}-d-2},1)\coloneqq f_{2}(x)-(x,0,\bm{0}_{d_{\mathrm{model}}-d-2},1)$ for $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$ and set $f$ to be zero otherwise. Let $\mu_{1}$ denote the distribution of $(X,Y)$ and let $\mu_{2}$ denote the marginal distribution of $X$. Define a distribution $\mu$ on $\mathbb{R}^{d_{\mathrm{model}}}$ by
$\mu\biggl(\prod_{j=1}^{d_{\mathrm{model}}}A_{j}\biggr)\coloneqq\frac{\mu_{1}(\prod_{j=1}^{d+1}A_{j})\mathbbm{1}_{\{0\in A_{d_{\mathrm{model}}}\}}+\mu_{2}(\prod_{j=1}^{d}A_{j})\mathbbm{1}_{\{1\in A_{d_{\mathrm{model}}}\}}}{2}\cdot\prod_{j=d+2}^{{d_{\mathrm{model}}}-1}\mathbbm{1}_{\{0\in A_{j}\}},$
for all measurable sets $A_{1},\ldots,A_{d_{\mathrm{model}}}\subseteq\mathbb{R}$. By Lemma 11 and the definition of $\mu$ above, there exists an FFN layer $\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ such that
$\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ $(X,Y,\bm{0}_{{d_{\mathrm{model}}}-d-2},0)-f_{1}(X,Y)\bigr\|_{\infty}>\epsilon\Bigr)$
$+\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(X,0,\bm{0}_{{d_{\mathrm{model}}}-d-2},1)-f_{2}(X)\bigr\|_{\infty}>\epsilon\Bigr)\leq\frac{\delta}{N+1}.$
We conclude the claim by a union bound. ∎
Proof of Theorem 2.
Let $\delta\coloneqq\frac{7\epsilon}{80R^{2}}\wedge 1$. By Lemma 10 and a union bound, there exists $b\in(0,1]$ such that the event
$\mathcal{E}_{0}\coloneqq\Bigl\{\min_{n\in[N]}\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\geq b\Bigr\}$
has probability at least $1-\delta$. Let $\epsilon_{1}\coloneqq\frac{b}{8}\wedge\frac{b\sqrt{\epsilon}}{64(R\vee 1)}$. By Lemma 9, there exist $M\geq d$ and $g_{1},\ldots,g_{M}\in\mathcal{G}$ such that the event
$\mathcal{E}_{1}\coloneqq\biggl\{\max_{n\in[N]}\biggl|\frac{1}{M}\sum_{m=1}^{M}g_{m}(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)-\int_{\mathcal{G}}g(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|\leq\epsilon_{1}\biggr\}$
$\bigcap\biggl\{\max_{n\in[N]}\biggl|\frac{1}{M}\sum_{m=1}^{M}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)-\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|\leq\epsilon_{1}\biggr\}$
has probability at least $1-\delta$. Now let ${d_{\mathrm{model}}}\coloneqq 4M+1$ and define $f_{1}=(f_{1,1},\ldots,f_{1,{d_{\mathrm{model}}}}):\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d_{\mathrm{model}}}$ by
$f_{1,j}(x,y)\coloneqq\begin{cases}\log\bar{f}_{\xi}\bigl(y,g_{j}(x)\bigr)\quad&\text{if }j\in[M]\\ \log\bar{f}_{\xi}\bigl(y,g_{j-M}(x)\bigr)\quad&\text{if }j\in[M+1:2M]\\ 0&\text{if }j\in[2M+1:4M+1].\end{cases}$
Also define $f_{2}=(f_{2,1},\ldots,f_{2,{d_{\mathrm{model}}}}):\mathbb{R}^{d}\to\mathbb{R}^{d_{\mathrm{model}}}$ by
$f_{2,j}(x)\coloneqq\begin{cases}\log g_{j}(x)\quad&\text{if }j\in[M]\\ 0&\text{if }j\in[M+1:4M]\\ 1&\text{if }j=4M+1.\end{cases}$
Finally, for $n\in[N]$, define $U^{(n)}\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$ by
$U^{(n)}\coloneqq\begin{pmatrix}f_{1}(X_{1},Y_{1})\\ \vdots\\ f_{1}(X_{n},Y_{n})\\ f_{2}(X)\end{pmatrix}$.
By Lemma 12, there exists an FFN layer $\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}$ with $\theta_{\mathrm{ffn}}^{(1)}=(W_{1},W_{2},v)$, such that writing $\epsilon_{2}\coloneqq\frac{b}{32(N+1)}\wedge\frac{b\sqrt{\epsilon}}{256R(N+1)}$, the event
$\mathcal{E}_{2}\coloneqq\Bigl\{\max_{n\in[N]}\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})-U^{(n)}\bigr\|_{\infty}\leq\epsilon_{2}\Bigr\}$
has probability at least $1-\delta$. Since the final column of $W_{2}$ only affects the final column of $\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})$, we can choose the final column of $W_{2}$ to be zero, so that
$\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\bigr]_{i,{d_{\mathrm{model}}}}=\bigl[Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})}\bigr]_{i,{d_{\mathrm{model}}}}=\mathbbm{1}_{\{i=n+1\}}=\bigl[U^{(n)}\bigr]_{i,{d_{\mathrm{model}}}}\quad\text{for }i\in[n+1].$ (18)
Next, we define $Q^{(2)}=K^{(2)}\coloneqq\bm{0}_{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ and define $V^{(2)}\in\mathbb{R}^{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ by $V^{(2)}_{j,2M+j}\coloneqq 1$ for $j\in[2M]$, $V^{(2)}_{4M+1,4M+1}\coloneqq 1$ and zero otherwise. Let $\theta^{(2)}_{\mathrm{attn}}\coloneqq(Q^{(2)},K^{(2)},V^{(2)})$, $\widetilde{Z}^{(n)}\coloneqq\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$ and $\bar{Z}^{(n)}\coloneqq\mathsf{Attn}_{\theta^{(2)}_{\mathrm{attn}}}(\widetilde{Z}^{(n)})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$. Then by construction and (18), the $(n+1)$th row of $\bar{Z}^{(n)}$ is given by
$\bar{Z}^{(n)}_{n+1,j}=\begin{cases}\widetilde{Z}^{(n)}_{n+1,j}\quad&\text{if }j\in[2M]\\ \widetilde{Z}^{(n)}_{n+1,j}+\frac{1}{n+1}\sum_{i=1}^{n+1}\widetilde{Z}^{(n)}_{i,j-2M}\quad&\text{if }j\in[2M+1:4M]\\ 1+\frac{1}{n+1}\quad&\text{if }j=4M+1.\end{cases}$ (19)
Thus, for $n\in[N]$ and $m\in[M]$, we have that on $\mathcal{E}_{2}$,
$\biggl|\bar{Z}^{(n)}_{n+1,2M+m}-\frac{1}{n+1}\sum_{i=1}^{n}$ $\log\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)-\frac{1}{n+1}\log g_{m}(X)\biggr|$
[…]
$\leq\bigl|\widetilde{Z}^{(n)}_{n+1,2M+m}-0\bigr|+\frac{1}{n+1}\sum_{i=1}^{n+1}\bigl|\widetilde{Z}^{(n)}_{i,m}-U^{(n)}_{i,m}\bigr|\leq 2\epsilon_{2}.$ (20)
Similarly, for $n\in[N]$ and $m\in[M]$, we have that on $\mathcal{E}_{2}$,
$\biggl|\bar{Z}^{(n)}_{n+1,3M+m}-\frac{1}{n+1}\sum_{i=1}^{n}\log\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)\biggr|\leq 2\epsilon_{2}.$ (21)
Now define a continuous function $f_{3}:\mathbb{R}^{4M+1}\to\mathbb{R}$ by
$f_{3}(z_{1},\ldots,z_{4M+1})\coloneqq\frac{\sum_{m=1}^{M}\exp\{z_{2M+m}/(z_{4M+1}-1)\}}{\sum_{m=1}^{M}\exp\{z_{3M+m}/(z_{4M+1}-1)\}}.$
For $n\in[N]$, let $\bar{P}^{(n)}$ be the distribution of the last row of $\bar{Z}^{(n)}$. By applying Lemma 11 with $\mu=\frac{1}{N}\sum_{n=1}^{N}\bar{P}^{(n)}$ therein, there exists an FFN layer $\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(2)}}$ such that the event
$\mathcal{E}_{3}\coloneqq\Bigl\{\max_{n\in[N]}\Bigl|\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(2)}}(\bar{Z}^{(n)})\bigr]_{n+1,d+1}-f_{3}\bigl(\bar{Z}^{(n)}_{n+1,1},\ldots,\bar{Z}^{(n)}_{n+1,{d_{\mathrm{model}}}}\bigr)\Bigr|\leq\sqrt{\epsilon}/4\Bigr\}$
has probability at least $1-\delta$. Since $\|g_{m}\|_{\infty}\leq R$ and $\|\bar{f}_{\xi}\|_{\infty}\leq 1$, we have by (19) and (20) that on $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, for all $n\in[N]$,
$\biggl|\frac{1}{M}\sum_{m=1}^{M}\exp\bigl\{\bar{Z}^{(n)}_{n+1,2M+m}/(\bar{Z}^{(n)}_{n+1,4M+1}-1)\bigr\}-\int_{\mathcal{G}}g(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|$
$\leq\biggl|\frac{1}{M}\sum_{m=1}^{M}\exp\bigl\{\bar{Z}^{(n)}_{n+1,2M+m}/(\bar{Z}^{(n)}_{n+1,4M+1}-1)\bigr\}-\frac{1}{M}\sum_{m=1}^{M}g_{m}(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)\biggr|+\epsilon_{1}$
$\leq\frac{1}{M}\sum_{m\in[M]}\biggl|\exp\bigl\{(n+1)\bar{Z}^{(n)}_{n+1,2M+m}\bigr\}-g_{m}(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)\biggr|+\epsilon_{1}$
$\leq R(e^{2(N+1)\epsilon_{2}}-1)+\epsilon_{1}\leq\frac{b\sqrt{\epsilon}}{32}\eqqcolon\epsilon_{3}.$ (22)
Similarly, on $\mathcal{E}_{1}\cap\mathcal{E}_{2}$ we have by (19) and (21) that for all $n\in[N]$,
$\biggl|\frac{1}{M}\sum_{m=1}^{M}\exp\bigl\{\bar{Z}^{(n)}_{n+1,3M+m}/(\bar{Z}^{(n)}_{n+1,4M+1}$ $-1)\bigr\}-\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)\biggr|$
$\leq(e^{2(N+1)\epsilon_{2}}-1)+\epsilon_{1}\leq\frac{b\sqrt{\epsilon}}{32R}\wedge\frac{b}{4}\eqqcolon\epsilon_{4}.$ (23)
Now, on the event $\mathcal{E}_{0}\cap\mathcal{E}_{1}$, since $\epsilon_{1}\leq b/2$, we have that
$\min_{n\in[N]}\frac{1}{M}\sum_{m=1}^{M}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(Y_{i},g_{m}(X_{i})\bigr)\geq\frac{b}{2}.$ (24)
Therefore, on $\mathcal{E}_{0}\cap\mathcal{E}_{1}\cap\mathcal{E}_{2}$, we have by (22), (23), (24) and (16) that for all $n\in[N]$
$\bigl|f_{3}$ $(\bar{Z}^{(n)}_{n+1,1},\ldots,\bar{Z}^{(n)}_{n+1,{d_{\mathrm{model}}}})-g_{\pi}(X,\mathcal{D}_{n})\bigr|$
$=\Biggl|\frac{\sum_{m=1}^{M}\exp\{\bar{Z}^{(n)}_{n+1,2M+m}/(\bar{Z}^{(n)}_{n+1,4M+1}-1)\}}{\sum_{m=1}^{M}\exp\{\bar{Z}^{(n)}_{n+1,3M+m}/(\bar{Z}^{(n)}_{n+1,4M+1}-1)\}}-\frac{\int_{\mathcal{G}}g(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)}{\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)}\Biggr|$
$\leq\biggl(\frac{1}{1-2\epsilon_{4}/b}-1\biggr)\Biggl|\frac{\int_{\mathcal{G}}g(X)\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)}{\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)}\Biggr|+\frac{\epsilon_{3}}{b/2-\epsilon_{4}}$
$\leq\biggl(\frac{1}{1-2\epsilon_{4}/b}-1\biggr)R+\frac{2\epsilon_{3}}{b-2\epsilon_{4}}\leq\frac{4R\epsilon_{4}}{b}+\frac{4\epsilon_{3}}{b}\leq\frac{\sqrt{\epsilon}}{4},$ (25)
where in the penultimate inequality, we used the fact that $\frac{1}{1-x}-1\leq 2x$ for $x\in[0,1/2]$ and that $\epsilon_{4}\leq b/4$, and the final inequality uses the definitions of $\epsilon_{3}$ and $\epsilon_{4}$. If we set the first attention layer $\mathsf{Attn}_{\theta^{(1)}_{\mathrm{attn}}}$ to be the identity map (i.e. set the parameters of the attention layer to zero), and let $\check{Z}^{(n)}\coloneqq\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(2)}}\circ\mathsf{Attn}_{\theta^{(2)}_{\mathrm{attn}}}\circ\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}\circ\mathsf{Attn}_{\theta^{(1)}_{\mathrm{attn}}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})$, then by (25) and the triangle inequality, we have on $\mathcal{E}_{0}\cap\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}$ that for all $n\in[N]$
$\Bigl|\check{Z}^{(n)}_{n+1,d+1}-g_{\pi}(X,\mathcal{D}_{n})\Bigr|=\Bigl|\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(2)}}(\bar{Z}^{(n)})\bigr]_{n+1,d+1}-g_{\pi}(X,\mathcal{D}_{n})\Bigr|\leq\frac{\sqrt{\epsilon}}{2}.$ (26)
For $n\in[N]$, let $\check{P}^{(n)}$ be the distribution of the last row of $\check{Z}^{(n)}$. By applying Lemma 11 with $\mu=\frac{1}{N}\sum_{n=1}^{N}\check{P}^{(n)}$ therein, there exists an FFN layer $\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(3)}}$ such that the event
$\mathcal{E}_{4}\coloneqq\Bigl\{\max_{n\in[N]}\Bigl|\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(3)}}(\check{Z}^{(n)})\bigr]_{n+1,d+1}-\mathsf{clip}_{R}(\check{Z}^{(n)}_{n+1,d+1})\Bigr|\leq\sqrt{\epsilon}/4\Bigr\}$
has probability at least $1-\delta$. Finally, let $\mathsf{Attn}_{\theta^{(3)}_{\mathrm{attn}}}$ be the identity map and let
$\mathsf{TF}\coloneqq\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(3)}}\circ\mathsf{Attn}_{\theta^{(3)}_{\mathrm{attn}}}\circ\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(2)}}\circ\mathsf{Attn}_{\theta^{(2)}_{\mathrm{attn}}}\circ\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}\circ\mathsf{Attn}_{\theta^{(1)}_{\mathrm{attn}}}$.
Since $\|g_{\pi}\|_{\infty}\leq R$, we have by (26) that on $\mathcal{E}\coloneqq\mathcal{E}_{0}\cap\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}\cap\mathcal{E}_{4}$, for all $n\in[N]$,
$\bigl|\mathsf{Read}\circ\mathsf{TF}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})-g_{\pi}(X,\mathcal{D}_{n})\bigr|$
$\leq\Bigl|\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(3)}}(\check{Z}^{(n)})\bigr]_{n+1,d+1}-\mathsf{clip}_{R}(\check{Z}^{(n)}_{n+1,d+1})\Bigr|+\Bigl|\mathsf{clip}_{R}(\check{Z}^{(n)}_{n+1,d+1})-g_{\pi}(X,\mathcal{D}_{n})\Bigr|\leq\frac{3\sqrt{\epsilon}}{4}.$
By a union bound, the event $\mathcal{E}$ has probability at least $1-5\delta$. We conclude that
$\max_{n\in[N]}\mathbb{E}_{P\sim\pi,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\Bigl\{\bigl(\mathsf{Read}\circ\mathsf{TF}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}\leq\frac{9\epsilon}{16}+5R^{2}\delta\leq\epsilon,$
as required. ∎
C.2 Proof of Proposition 3
The main idea behind the proof of Proposition 3 is to use empirical process theory to upper bound $\mathcal{E}(\widetilde{f}_{T})$ by the approximation error and a complexity measure of $\mathcal{F}$. In the literature, the complexity measure is often taken to be the covering number with respect to the uniform norm, and upper bounds on this covering number for Transformers are available, provided the input matrix and parameters are bounded. There, the proof strategy is to argue that for any input matrix from a bounded set, the output of the Transformer with bounded parameters is uniformly Lipschitz in its parameters. However, this argument does not work in our setting, since our input matrix and parameters are unbounded. Instead, we use the pseudo-dimension of $\mathcal{F}$ (see the definition below) as a complexity measure and show that the pseudo-dimension of Transformers is finite. This result (Lemma 15 below) may be of independent interest.
Definition 5 (VC-dimension and pseudo-dimension).
Let $X$ be a set and let $\mathcal{C}$ be a collection of subsets of $X$. For $S\coloneqq\{x_{1},\ldots,x_{n}\}\subseteq X$, we say that $\mathcal{C}$ shatters $S$ if for every $T\subseteq S$, there exists $C\in\mathcal{C}$ such that $C\cap S=T$. The VC-dimension of $\mathcal{C}$ is the cardinality of the largest subset of $X$ that can be shattered by $\mathcal{C}$.
Now let $\mathcal{G}$ be a set of functions from $\mathcal{X}$ to $\mathbb{R}$. The pseudo-dimension $\mathrm{Pdim}(\mathcal{G})$ of $\mathcal{G}$ is the VC-dimension of its subgraph class $\bigl\{\{(x,y)\in\mathcal{X}\times\mathbb{R}:y\leq g(x)\}:g\in\mathcal{G}\bigr\}$.
Lemma 13 below controls $\mathcal{E}(\widetilde{f}_{T})$ in terms of the approximation error and pseudo-dimension of $\mathcal{F}$.
Lemma 13.
Under the setting of Proposition 3, there exists $C>0$, depending only on $\|\xi\|_{\psi_{2}}$ and $R$, such that for $T\geq 2$,
$\mathcal{E}(\widetilde{f}_{T})\leq 2\inf_{f\in\mathcal{F}}\mathbb{E}_{P\sim\pi,\,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\Bigl\{\bigl(f(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}+\frac{C\log^{3}(T)\cdot\mathrm{Pdim}(\mathcal{F})}{T}.$
Proof.
Let $P\sim\pi$, $(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}$ and let $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t\in[T]}$ be independent copies of $(\mathcal{D}_{n},X,Y)$. For $t\in[T]$, let $Z^{(t)}\coloneqq(X^{(t)},\mathcal{D}_{n}^{(t)})$ and let $Z\coloneqq(X,\mathcal{D}_{n})$. Now note that $g_{\pi}(Z)=\mathbb{E}(Y\,|\,Z)$,
$\widehat{f}_{T}\in\operatorname*{argmin}_{f\in\mathcal{F}}\frac{1}{T}\sum_{t=1}^{T}\bigl(Y^{(t)}-f(Z^{(t)})\bigr)^{2}$
and
$\mathcal{E}(\widetilde{f}_{T})=\mathbb{E}\bigl\{\bigl(Y-\mathsf{clip}_{R}\widehat{f}_{T}(Z)\bigr)^{2}\bigr\}-\mathbb{E}\bigl\{\bigl(Y-g_{\pi}(Z)\bigr)^{2}\bigr\}.$
We have therefore reduced our problem to bounding the excess risk of a nonparametric regression estimator in a problem with bounded regression function and sub-Gaussian noise. The conclusion now follows by applying, for example, Ma, Wang and Samworth (2025, Theorem 1) with the following minor modifications: (i) we ignore their $\Omega$ vectors since our data do not have missing values, (ii) by inspecting their proof, we may truncate $\widehat{f}_{T}$ at level $R$ since this is a known bound on the regression function, (iii) the optimisation error term (the first term in their upper bound) is zero since we consider the empirical risk minimiser, and (iv) by inspecting their proof and using Györfi et al. (2002, Theorem 9.4), we can replace the third term in their upper bound by $\frac{C\log^{3}(T)\cdot\mathrm{Pdim}(\mathcal{F})}{T}$, where $C>0$ depends only on $\|\xi\|_{\psi_{2}}$ and $R$. ∎
Next we will show that the class of Transformers with fixed architecture has finite pseudo-dimension. The proof uses terminology and results from model theory, a branch of mathematical logic. Since this field will probably be unfamiliar to most readers, we now provide a minimal glossary (sufficient for our proof) of the relevant terms; we refer the readers to Marker (2002) for a formal introduction to model theory. We will be concerned with an object called a structure, which is a set equipped with a collection of distinguished functions, relations and elements that together give meaning to the symbols of a formal language. Our structure of interest is $\mathbb{R}_{\exp,\Phi}\coloneqq(\mathbb{R};0,1,+,-,\cdot,<,\exp,\Phi)$, where $\cdot$ refers to multiplication, $\exp:\mathbb{R}\to(0,\infty)$ refers to the exponential function and $\Phi:\mathbb{R}\to[0,1]$ refers to the distribution function of the standard Gaussian distribution. Our goal is to show that Transformers can be defined within $\mathbb{R}_{\exp,\Phi}$, in a sense that we will make precise below.
Given a countable set $\mathcal{V}$ of variable symbols, which we intrepret as syntactic placeholders, the set of terms in the structure $\mathbb{R}_{\exp,\Phi}\coloneqq(\mathbb{R};0,1,+,-,\cdot,<,\exp,\Phi)$ is the smallest set $\mathcal{T}$ such that (i) $0,1\in\mathcal{T}$, (ii) $\mathcal{V}\subseteq\mathcal{T}$, (iii) if $t_{1},t_{2}\in\mathcal{T}$, then $t_{1}+t_{2},\,t_{1}-t_{2},\,t_{1}\cdot t_{2}\in\mathcal{T}$, and (iv) if $t\in\mathcal{T}$, then $\exp(t),\Phi(t)\in\mathcal{T}$. We say that $\varphi$ is an atomic formula in $\mathbb{R}_{\exp,\Phi}$ if $\varphi$ is either $t_{1}=t_{2}$ or $t_{1}<t_{2}$, for some $t_{1},t_{2}\in\mathcal{T}$. The set of (first-order) formulas $\mathcal{W}$ in (the language of) $\mathbb{R}_{\exp,\Phi}$ is the smallest set such that (i) every atomic formula is in $\mathcal{W}$, (ii) if $\varphi,\psi\in\mathcal{W}$, then $(\neg\varphi),(\varphi\vee\psi),(\varphi\wedge\psi)\in\mathcal{W}$, where the logic symbolds $\neg,\vee,\wedge$ stand for negation, disjunction and conjunction, and (iii) if $\varphi\in\mathcal{W}$ and $v\in\mathcal{V}$, then $\exists v\;\varphi$ and $\forall v\;\varphi$ are in $\mathcal{W}$. For a formula $\varphi\in\mathcal{W}$, we say that a variable $v$ is free in $\varphi$ if it is not inside a $\exists v$ or $\forall v$ quantifier; e.g. $v_{1}$ is free in the formula $(v_{1}=0)\vee(v_{1}>0)$, but it is not free in the formula $\exists v_{1}\;v_{2}\cdot v_{2}=v_{1}$. When a formula $\varphi$ has free variables $v_{1},\ldots,v_{n}$, we will often write $\varphi(v_{1},\ldots,v_{n})$ to make the free variables explicit. Now, letting $\varphi(v_{1},\ldots,v_{n})\in\mathcal{W}$ be a formula and $a=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$, we write $\mathbb{R}_{\exp,\Phi}\models\varphi(a)$ (we read $\mathbb{R}_{\exp,\Phi}$ entails $\varphi(a)$’ or $\mathbb{R}_{\exp,\Phi}$ models $\varphi(a)$’) if $\varphi(a)$ is true in $\mathbb{R}_{\exp,\Phi}$, and otherwise we write $\mathbb{R}_{\exp,\Phi}\not\models\varphi(a)$. For example, if $\varphi(v_{1},v_{2})$ is the formula $(v_{1}=v_{2})\vee(v_{1}+v_{2}=0)$, then $\mathbb{R}_{\exp,\Phi}\models\varphi(1,1)$ and $\mathbb{R}_{\exp,\Phi}\models\varphi(-1,1)$, but $\mathbb{R}_{\exp,\Phi}\not\models\varphi(1,2)$.
For $d\in\mathbb{N}$ and a set $A\subseteq\mathbb{R}^{d}$, we say that $A$ is definable (in $\mathbb{R}_{\exp,\Phi}$) if there exist $m\in\mathbb{N}$, a formula $\varphi(v_{1},\ldots,v_{d},w_{1},\ldots,w_{m})$ in $\mathbb{R}_{\exp,\Phi}$ and $b\in\mathbb{R}^{m}$ such that $A=\{a\in\mathbb{R}^{d}:\mathbb{R}_{\exp,\Phi}\models\varphi(a,b)\}$; we will sometimes abbreviate the notation and write $A=\{a\in\mathbb{R}^{d}:\varphi(a,b)\}$ instead. For $d_{1},d_{2}\in\mathbb{N}$ and a function $f:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$, we say that $f$ is definable (in $\mathbb{R}_{\exp,\Phi}$) if its graph $\{(x,y)\in\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}:y=f(x)\}$ is a definable set in $\mathbb{R}_{\exp,\Phi}$. Moreover, for a collection $\mathcal{C}$ of subsets of $\mathbb{R}^{d}$, we say that $\mathcal{C}$ is uniformly definable (in $\mathbb{R}_{\exp,\Phi}$) if there exists a formula $\varphi(v,w)$, where $v=(v_{1},\ldots,v_{d})$ and $w=(w_{1},\ldots,w_{m})$ for some $m\in\mathbb{N}$, such that for every $C\in\mathcal{C}$, there exists $b_{C}\in\mathbb{R}^{m}$ with $C=\{a\in\mathbb{R}^{d}:\varphi(a,b_{C})\}$. Similarly, for a collection $\mathcal{F}$ of functions from $\mathbb{R}^{d_{1}}$ to $\mathbb{R}^{d_{2}}$, we say that $\mathcal{F}$ is uniformly definable (in $\mathbb{R}_{\exp,\Phi}$) if the collection of graphs of functions in $\mathcal{F}$ is uniformly definable.
Lemma 14.
  • (a)
    Let $\mathcal{F}$ be a collection of uniformly definable functions from $\mathbb{R}^{d_{1}}$ to $\mathbb{R}$, and let $\mathcal{G}$ be a collection of uniformly definable functions from $\mathbb{R}^{d_{2}}$ to $\mathbb{R}$. Then $\bigl\{\{(x,y)\in\mathbb{R}^{d_{1}}\times\mathbb{R}:y\leq f(x)\}:f\in\mathcal{F}\bigr\}$, $\{(x,z)\mapsto f(x)+g(z):f\in\mathcal{F},g\in\mathcal{G}\}$ and $\{(x,z)\mapsto f(x)\cdot g(z):f\in\mathcal{F},g\in\mathcal{G}\}$ are all uniformly definable.
  • (b)
    Let $\mathcal{F}$ be a collection of uniformly definable functions from $\mathbb{R}^{d}$ to $\mathbb{R}\setminus\{0\}$. Then $\{x\mapsto 1/f(x):f\in\mathcal{F}\}$ is uniformly definable.
  • (c)
    Let $p\in\mathbb{N}$ and for $j\in[p]$, let $\mathcal{G}_{j}$ be a collection of uniformly definable functions from $\mathbb{R}^{d_{j}}$ to $\mathbb{R}$ for some $d_{j}\in\mathbb{N}$. Further let $\mathcal{F}$ be a collection of uniformly definable functions from $\mathbb{R}^{p}$ to $\mathbb{R}$. Then $\bigl\{(x_{1},\ldots,x_{p})\mapsto f\bigl(g_{1}(x_{1}),\ldots,g_{p}(x_{p})\bigr):f\in\mathcal{F},g_{j}\in\mathcal{G}_{j}\,\forall j\in[p]\bigr\}$ is uniformly definable.
Proof.
(a) Suppose that for each $f\in\mathcal{F}$, the graph of $f$ is $\{(x,y)\in\mathbb{R}^{d_{1}}\times\mathbb{R}:\varphi(x,y,b_{f})\}$ for some formula $\varphi$, some $b_{f}\in\mathbb{R}^{m}$ and some $m\in\mathbb{N}$; also suppose that for each $g\in\mathcal{G}$, the graph of $g$ is $\{(z,y)\in\mathbb{R}^{d_{2}}\times\mathbb{R}:\psi(z,y,c_{g})\}$ for some formula $\psi$, some $c_{g}\in\mathbb{R}^{n}$ and some $n\in\mathbb{N}$. The subgraph $\{(x,y)\in\mathbb{R}^{d_{1}}\times\mathbb{R}:y\leq f(x)\}$ of $f\in\mathcal{F}$ can be written as $\bigl\{(x,y)\in\mathbb{R}^{d_{1}}\times\mathbb{R}:\exists t\;\bigl(\varphi(x,t,b_{f})\wedge(y<t\vee y=t)\bigr)\bigr\}$. Hence the collection of subgraphs of $f\in\mathcal{F}$ is uniformly definable. For $f\in\mathcal{F}$ and $g\in\mathcal{G}$, the graph of the function $(x,z)\mapsto f(x)+g(z)$ can be written as $\bigl\{(x,z,y)\in\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\times\mathbb{R}:\exists u,v\;\bigl(\varphi(x,u,b_{f})\wedge\psi(z,v,c_{g})\wedge(y=u+v)\bigr)\bigr\}$. Finally, the graph of the function $(x,z)\mapsto f(x)\cdot g(z)$ can be written as $\bigl\{(x,z,y)\in\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\times\mathbb{R}:\exists u,v\;\bigl(\varphi(x,u,b_{f})\wedge\psi(z,v,c_{g})\wedge(y=u\cdot v)\bigr)\bigr\}$.
(b) Again, suppose that for each $f\in\mathcal{F}$, the graph of $f$ is $\{(x,y)\in\mathbb{R}^{d}\times\mathbb{R}:\varphi(x,y,b_{f})\}$ for some formula $\varphi$, some $b_{f}\in\mathbb{R}^{m}$ and some $m\in\mathbb{N}$. Then, for $f\in\mathcal{F}$, the graph of the function $x\mapsto 1/f(x)$ can be written as $\bigl\{(x,y)\in\mathbb{R}^{d}\times\mathbb{R}:\exists u\;\bigl(\varphi(x,u,b_{f})\wedge(y\cdot u=1)\bigr)\bigr\}$. This proves the claim.
(c) Suppose that for each $f\in\mathcal{F}$, the graph of $f$ is $\{(x,y)\in\mathbb{R}^{p}\times\mathbb{R}:\varphi(x,y,b_{f})\}$ for some formula $\varphi$, some $b_{f}\in\mathbb{R}^{m}$ and some $m\in\mathbb{N}$; also suppose that for $j\in[p]$ and $g_{j}\in\mathcal{G}_{j}$, the graph of $g_{j}$ is $\{(z,y)\in\mathbb{R}^{d_{j}}\times\mathbb{R}:\psi_{j}(z,y,c_{g_{j}})\}$ for some formula $\psi_{j}$, some $c_{g_{j}}\in\mathbb{R}^{n_{j}}$ and some $n_{j}\in\mathbb{N}$. Then the graph of the function $(x_{1},\ldots,x_{p})\mapsto f\bigl(g_{1}(x_{1}),\ldots,g_{p}(x_{p})\bigr)$ can be written as
$\biggl\{(x_{1},\ldots,x_{p},y)\in\mathbb{R}^{p+1}:\exists v_{1},\ldots,v_{p}\biggl(\bigwedge_{j=1}^{p}\psi_{j}(x_{j},v_{j},c_{g_{j}})\wedge\varphi(v_{1},\ldots,v_{p},y,b_{f})\biggr)\biggr\}.$
This proves the claim. ∎
Lemma 15.
For any $L,H,{d_{\mathrm{model}}},d_{\mathrm{ffn}}\in\mathbb{N}$, the class $\mathcal{F}^{\prime}\coloneqq\mathsf{Read}\circ\mathcal{F}_{\mathrm{TF}}(L,H,{d_{\mathrm{model}}},d_{\mathrm{ffn}})$ has finite pseudo-dimension.
Remark.
It is possible to obtain upper bounds on the pseudo-dimension of $\mathcal{F}^{\prime}$ with polynomial dependence on $L$, $H$, $d_{\mathrm{model}}$} and $d_{\mathrm{ffn}}$ via a more detailed analysis, but a finite pseudo-dimension is already enough for our purpose.
Proof.
Note that $\mathcal{F}^{\prime}$ is defined on the space $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$, equipped with the disjoint union topology. Thus, it suffices to show that, for each $n\in[N]$, the restriction $\mathcal{F}^{\prime}_{n}$ of the functions in $\mathcal{F}^{\prime}$ to $\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$ has finite pseudo-dimension. Now let
$\mathcal{C}_{n}\coloneqq\Bigl\{\bigl\{(Z,y)\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}\times\mathbb{R}:y\leq f(Z)\bigr\}:f\in\mathcal{F}^{\prime}_{n}\Bigr\}$
denote the collection of subgraphs of $f\in\mathcal{F}^{\prime}_{n}$. Then by definition, the pseudo-dimension of $\mathcal{F}^{\prime}_{n}$ is equal to the VC-dimension of $\mathcal{C}_{n}$. We will now show that $\mathcal{C}_{n}$ is uniformly definable in $\mathbb{R}_{\exp,\Phi}$. The graph $\{(x,y)\in\mathbb{R}^{2}:y=\operatorname{\mathsf{ReLU}}(x)\}$ of $\operatorname{\mathsf{ReLU}}$ can be written as
$\bigl\{(x,y)\in\mathbb{R}^{2}:\bigl(\neg(x<0)\wedge y=x\bigr)\vee(x<0\wedge y=0)\bigr\},$
so $\operatorname{\mathsf{ReLU}}:\mathbb{R}\to\mathbb{R}$ is definable on $\mathbb{R}_{\exp,\Phi}$. By Lemma 14, sums, products, reciprocals and compositions of definable functions are definable. Thus, for any $d,p\in\mathbb{N}$, the set of polynomials on $\mathbb{R}^{d}$ with degree at most $p$ is uniformly definable in $\mathbb{R}_{\exp,\Phi}$. Moreover, since $\exp$ and $\Phi$ are definable in $\mathbb{R}_{\exp,\Phi}$ by definition, we deduce that $\mathsf{GELU}$, $\mathsf{SiLU}$, and $\operatorname{\mathsf{Softmax}}$ are also definable in $\mathbb{R}_{\exp,\Phi}$. Each function $f\in\mathcal{F}^{\prime}_{n}$ comprises compositions, sums and products of $\rho\in\{\mathsf{ReLU},\mathsf{GELU},\mathsf{SiLU}\}$, $\operatorname{\mathsf{Softmax}}$ and polynomials of degree at most two, we deduce by Lemma 14 that $\mathcal{F}^{\prime}_{n}$ is uniformly definable in $\mathbb{R}_{\exp,\Phi}$. Further, by Lemma 14(a), we have that the subgraph class of $\mathcal{F}^{\prime}_{n}$ is also uniformly definable on $\mathbb{R}_{\exp,\Phi}$ and
$\mathcal{C}_{n}\subseteq\Bigl\{\bigl\{a\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}\times\mathbb{R}:\mathbb{R}_{\exp,\Phi}\models\varphi(a,b)\bigr\}:b\in\mathbb{R}^{m}\Bigr\}.$
By Marker (2002, Therorem 3.4.37), the structure $\mathbb{R}_{\exp}\coloneqq(\mathbb{R};0,1,+,-,\cdot,<,\exp,\Phi)$ is o-minimal, i.e. every definable set is a finite union of open intervals and singleton points. Further, the derivative $\Phi^{\prime}$ given by $\Phi^{\prime}(x)=\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}$ is definable in $\mathbb{R}_{\exp}$, so by Speissegger (1999), there exists an o-minimal extension $\widetilde{\mathbb{R}}$ of $\mathbb{R}_{\exp}$ in which $\Phi$ is definable. Now, every definable set in $\mathbb{R}_{\exp,\Phi}$ is definable in $\widetilde{\mathbb{R}}$, so $\mathbb{R}_{\exp,\Phi}$ is o-minimal. We may therefore apply Laskowski (1992, Theorem 3.2) (see also Karpinski and Macintyre, 1995, Theorem 2) to deduce that the VC-dimension of $\mathcal{C}_{n}$ is finite. ∎
Proof of Proposition 3.
By Theorem 2, there exist $d_{\mathrm{model}}^{\circ}\geq d+2$ and $d_{\mathrm{ffn}}^{\circ}\in\mathbb{N}$ such that if ${d_{\mathrm{model}}}\geq d_{\mathrm{model}}^{\circ}$, $d_{\mathrm{ffn}}\geq d_{\mathrm{ffn}}^{\circ}$ and $\mathcal{F}\coloneqq\mathsf{Read}\circ\mathcal{F}_{\mathrm{TF}}(3,1,{d_{\mathrm{model}}},d_{\mathrm{ffn}})$, then
$\inf_{f\in\mathcal{F}}\mathbb{E}_{P\sim\pi,\,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\Bigl\{\bigl(f(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}\leq\frac{\epsilon}{4},$
Now by Lemma 15, we have that $\mathrm{Pdim}(\mathcal{F})$ is finite. Thus, for all $T$ sufficiently large, we have $\frac{C\log^{3}(T)\cdot\mathrm{Pdim}(\mathcal{F})}{T}\leq\frac{\epsilon}{2}$, where $C>0$ is taken from Lemma 13. The claim then follows from Lemma 13. ∎
Appendix D Function spaces and wavelets
D.1 Function spaces
In this section, we restrict the domain of the functions to $0,1]^{d}$. By Jensen’s inequality, $\|f\|_{L^{p_{1}}}\leq\|f\|_{L^{p_{2}}}$ for all $1\leq p_{1}\leq p_{2}\leq\infty$. For $f,g\in L^{2}([0,1]^{d})$, define the inner product
$\langle f,g\rangle\coloneqq\int_{[0,1]^{d}}f(x)g(x)\,\mathrm{d}x.$
Let $C([0,1]^{d})$ denote the set of continuous functions from $0,1]^{d}$ to $\mathbb{R}$. For $m=(m_{1},\ldots,m_{d})^{\top}\in\mathbb{N}_{0}^{d}$ with $k\coloneqq\|m\|_{1}$, and a $k$-times differentiable $f:[0,1]^{d}\rightarrow\mathbb{R}$, we use the multi-index notation $\partial^{m}f\coloneqq\partial^{m_{1}}\cdots\partial^{m_{d}}f$ for partial derivatives.
Definition 6 (Hölder space).
For $\alpha>0$, let $\alpha_{0}\coloneqq\lceil\alpha\rceil-1$ be the largest integer strictly smaller than $\alpha$ and define the Hölder space $H^{\alpha}([0,1]^{d})$ to be the set of $\alpha_{0}$-times differentiable $f:[0,1]^{d}\to\mathbb{R}$ with
$\sup_{x\neq y}\sum_{m\in\mathbb{N}_{0}^{d}:\|m\|_{1}=\alpha_{0}}\frac{\partial^{m}f(x)-\partial^{m}f(y)}{\|x-y\|_{2}^{\alpha-\alpha_{0}}}<\infty.$
For $r\in\mathbb{N}$, define the $r$th order difference operator $\Delta_{h}^{r}:L^{1}([0,1]^{d})\rightarrow L^{1}([0,1]^{d})$ by
$\Delta_{h}^{r}(f)(x)=\mathbbm{1}_{\{x+rh\in[0,1]^{d}\}}\sum_{k=0}^{r}\binom{r}{k}(-1)^{r-k}f(x+kh).$
Note that $\Delta_{h}^{1}(f)(x)=\{f(x+h)-f(x)\}\mathbbm{1}_{\{x+h\in[0,1]^{d}\}}$ and $\Delta_{h}^{r}(f)(x)=\Delta_{h}^{1}\bigl(\Delta_{h}^{r-1}(f)(x)\bigr)$ if $x+rh\in[0,1]^{d}$. Now, for $p\in[1,\infty]$, $f\in L^{p}([0,1]^{d})$ and $t\geq 0$, define the $r$th modulus of smoothness
$\omega_{r,p}(f,t)\coloneqq\sup_{\|h\|_{2}\leq t}\|\Delta_{h}^{r}(f)\|_{L^{p}}.$
Definition 7 (Besov space).
For $p,q\in[1,\infty]$ and $\alpha>0$, let $r\coloneqq\lfloor\alpha\rfloor+1$ and define the Besov space $B_{p,q}^{\alpha}([0,1]^{d})$ by
$B_{p,q}^{\alpha}([0,1]^{d})\coloneqq\begin{cases}{\displaystyle\biggl\{f\in L^{p}([0,1]^{d}):\biggl[\int_{0}^{\infty}\bigl(t^{-\alpha}\omega_{r,p}(f,t)\bigr)^{q}\frac{\mathrm{d}t}{t}\biggr]^{1/q}<\infty\biggr\}}\quad&\text{for }q\in[1,\infty)\\[10.00002pt] {\displaystyle\Bigl\{f\in L^{p}([0,1]^{d}):\sup_{t>0}t^{-\alpha}\omega_{r,p}(f,t)<\infty\Bigr\}}\quad&\text{for }q=\infty.\end{cases}$
For non-integer smoothness levels, Hölder spaces are special cases of Besov spaces in the sense that $H^{\alpha}([0,1]^{d})=B^{\alpha}_{\infty,\infty}([0,1]^{d})$ for all $\alpha\in(0,\infty)\setminus\mathbb{N}$, while for $\alpha\in\mathbb{N}$, we have $H^{\alpha}([0,1]^{d})\subseteq B^{\alpha}_{\infty,\infty}([0,1]^{d})$, and in fact, $B_{\infty,\infty}^{\alpha}([0,1]^{d})$ coincides with Zygmund spaces when $\alpha\in\mathbb{N}$; see e.g. Giné and Nickl (2021, p. 351) or Triebel (1983, Chapter 2.5.7).
D.2 Wavelets
We start with the univariate setting where the notation is simpler. For $S\in\mathbb{N}$, denote the Cohen–Daubechies–Vial (CDV) wavelet basis on $[0,1]$ (Cohen, Daubechies and Vial, 1993), constructed from $S$-regular and $S$-times continuously differentiable father and mother Daubechies wavelets (Daubechies, 1988), by
$\bigl\{\phi_{k}:k\in[0:2^{\ell_{0}}-1]\bigr\}\bigcup\bigl\{\psi_{\ell,k}:\ell\in[{\ell_{0}}:\infty),\,k\in[0:2^{\ell}-1]\bigr\},$ (27)
where $\ell_{0}\in\mathbb{N}$ depends only on $S$. The precise definition and construction of this wavelet basis is not important for us, but the functions in (27) are $S$-times continuously differentiable and form an orthonormal basis for $L^{2}([0,1])$ (Cohen, Daubechies and Vial, 1993, Theorem 4.4). Moreover, $\phi_{k}$ is supported on $\bigl[\frac{k}{2^{\ell_{0}}},\frac{k+1}{2^{\ell_{0}}}\bigr]$ for $k\in[0:2^{\ell_{0}}-1]$, and $\psi_{\ell,k}$ is supported on $\bigl[\frac{k}{2^{\ell}},\frac{k+1}{2^{\ell}}\bigr]$ for $\ell\in[{\ell_{0}}:\infty),\,k\in[0:2^{\ell}-1]$.
The one-dimensional CDV wavelets in (27) can be extended to an orthonormal basis for $L^{2}([0,1]^{d})$ via tensor products. For $k=(k_{1},\ldots,k_{d})^{\top}\in[0:2^{\ell_{0}}-1]^{d}$ and $x=(x_{1},\ldots,x_{d})^{\top}\in[0,1]^{d}$, define
$\Phi_{k}(x)\coloneqq\prod_{j=1}^{d}\phi_{k_{j}}(x_{j}).$ (28)
Further, for $\tau=(\tau_{1},\ldots,\tau_{d})^{\top}\in\{0,1\}^{d}\setminus\{\bm{0}_{d}\}$, $\ell\in[{\ell_{0}}:\infty)$, $k=(k_{1},\ldots,k_{d})^{\top}\in[0:2^{\ell}-1]^{d}$ and $x=(x_{1},\ldots,x_{d})^{\top}\in[0,1]^{d}$, define
$\Psi_{\ell,k}^{(\tau)}(x)\coloneqq\prod_{j=1}^{d}\psi^{(\tau_{j})}_{\ell,k_{j}}(x_{j}),$ (29)
where $\psi^{(1)}_{\ell,k^{\prime}}=\psi_{\ell,k^{\prime}}$ for all $k^{\prime}\in[0:2^{\ell}-1]$, and $(\psi^{(0)}_{\ell,k^{\prime}})_{k^{\prime}=0}^{2^{\ell}-1}$ are the boundary-corrected ($S$-regular and $S$-times continuously differentiable) father Daubechies wavelets at resolution $\ell$. Again, the precise definition and construction are not important for us, but we will exploit the following properties. To simplify notation, define
$K\coloneqq[0:2^{\ell_{0}}-1]^{d}\quad\text{and}\quad\Gamma_{\ell}\coloneqq[0:2^{\ell}-1]^{d}\times\bigl(\{0,1\}^{d}\setminus\{\bm{0}_{d}\}\bigr)\text{ for }\ell\in[\ell_{0}:\infty),$ (30)
and for $\gamma=(k,\tau)\in\Gamma_{\ell}$, we write $\Psi_{\ell,\gamma}\coloneqq\Psi_{\ell,k}^{(\tau)}$. Note that
$|K|=2^{\ell_{0}d}\quad\text{and}\quad|\Gamma_{\ell}|=2^{\ell d}(2^{d}-1)<2^{(\ell+1)d}\text{ for }\ell\in[\ell_{0}:\infty).$ (31)
Each term in the products in (28) and (29) is $S$-times continuously differentiable, and the functions
$\bigl\{\Phi_{k}:k\in K\bigr\}\cup\bigl\{\Psi_{\ell,\gamma}:\ell\in[{\ell_{0}}:\infty),\,\gamma\in\Gamma_{\ell}\bigr\}$
form an orthonormal basis for $L^{2}([0,1]^{d})$; see e.g. Giné and Nickl (2021, Chapter 4.3.6) and Dahmen (1997, Section 4). By construction, there exist $C,C^{\prime}>0$ (depending only on $S$ and $d$) such that
$\|\Phi_{k}\|_{\infty}\leq C2^{\ell_{0}d/2}\quad\text{and}\quad\|\Psi_{\ell,\gamma}\|_{\infty}\leq C2^{\ell d/2}$ (32)
for $k\in K$, $\ell\in[\ell_{0}:\infty)$ and $\gamma\in\Gamma_{\ell}$, and that
$\sum_{k\in K}\mathbbm{1}_{\{\Phi_{k}(x)\neq 0\}}\vee\sup_{\ell\in[\ell_{0}:\infty)}\sum_{\gamma\in\Gamma_{\ell}}\mathbbm{1}_{\{\Psi_{\ell,\gamma}(x)\neq 0\}}\leq C^{\prime}\quad\text{for all }x\in[0,1]^{d}.$ (33)
For $f\in L^{2}([0,1]^{d})$, we have the wavelet decomposition
$f=\sum_{k\in K}a_{k}\Phi_{k}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}b_{\ell,\gamma}\Psi_{\ell,\gamma},$ (34)
where $a_{k}\coloneqq\langle f,\Phi_{k}\rangle$ and $b_{\ell,\gamma}\coloneqq\langle f,\Psi_{\ell,\gamma}\rangle$. Define
$a\coloneqq(a_{k})_{k\in K}\in\mathbb{R}^{K}\quad\text{and}\quad b_{\ell}\coloneqq(b_{\ell,\gamma})_{\gamma\in\Gamma_{\ell}}\in\mathbb{R}^{\Gamma_{\ell}}\text{ for }\ell\in[\ell_{0}:\infty).$ (35)
For $f\in L^{2}([0,1]^{d})$, $p\in[2,\infty]$, $q\in[1,\infty]$ and $\alpha\in(0,S)$, define its Besov norm by
$\|f\|_{B_{p,q}^{\alpha}}\coloneqq\begin{cases}\displaystyle 2^{\ell_{0}(d/2-d/p)}\|a\|_{p}+\biggl\{\sum_{\ell={\ell_{0}}}^{\infty}\bigl(2^{\ell(\alpha+d/2-d/p)}\|b_{\ell}\|_{p}\bigr)^{q}\biggr\}^{1/q}\quad&\text{for }q\in[1,\infty)\\ \displaystyle 2^{\ell_{0}(d/2-d/p)}\|a\|_{p}+\sup_{\ell\in[\ell_{0}:\infty)}2^{\ell(\alpha+d/2-d/p)}\|b_{\ell}\|_{p}\quad&\text{for }q=\infty.\end{cases}$
For $p\in[2,\infty]$, $q\in[1,\infty]$ and $\alpha\in(0,S)$, the Besov space $B_{p,q}^{\alpha}([0,1]^{d})$ in Definition 7 can be equivalently characterised by wavelet coefficients in the sense that $f\in B_{p,q}^{\alpha}([0,1]^{d})$ if and only if $f\in L^{2}([0,1]^{d})$ and $\|f\|_{B_{p,q}^{\alpha}}<\infty$ (e.g. Giné and Nickl, 2021, Chapter 4.3.6). Therefore, we can define a closed ball $B_{p,q}^{\alpha}([0,1]^{d},R)$ of radius $R>0$ in the space $B_{p,q}^{\alpha}([0,1]^{d})$ by
$B_{p,q}^{\alpha}([0,1]^{d},R)\coloneqq\{f\in L^{2}([0,1]^{d}):\|f\|_{B_{p,q}^{\alpha}}\leq R\}.$ (36)
As a special case, we have $\|f\|_{B_{\infty,\infty}^{\alpha}}=2^{\ell_{0}d/2}\|a\|_{\infty}+\sup_{\ell\in[\ell_{0}:\infty)}2^{\ell(\alpha+d/2)}\|b_{\ell}\|_{\infty}$.
Appendix E Posterior contraction and adaptation
E.1 General theory
This section is primarily based on material from Ghosal, Lember and van der Vaart (2008); see also Ghosal and van der Vaart (2017, Section 10.2). Here, we make some modifications in order to obtain finite-sample posterior contraction at an exponential rate.
Let $\mathcal{Z}$ be a measurable space and let $\mathcal{A}$ be a countable index set. For $\alpha\in\mathcal{A}$, let $\mathcal{P}_{\alpha}$ be a space of distributions on $\mathcal{Z}$ dominated by a $\sigma$-finite measure $\nu$, and let $\mathcal{P}\coloneqq\bigcup_{\alpha\in\mathcal{A}}\mathcal{P}_{\alpha}$, which is a metric space when equipped with the Hellinger metric $d_{\mathrm{H}}$. For example, $\alpha>0$ might index smoothness, in which case $\mathcal{P}_{\alpha}$ might denote the set of distributions of $(X,Y)$ for which the regression function $x\mapsto\mathbb{E}(Y\,|\,X=x)$ is $\alpha$-smooth. For $\alpha\in\mathcal{A}$, let $\pi_{\alpha}$ be a distribution on $\mathcal{P}_{\alpha}$ and let $\lambda=(\lambda_{\alpha})_{\alpha\in\mathcal{A}}$ be such that $\lambda_{\alpha}>0$ for $\alpha\in\mathcal{A}$ and $\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}=1$. We consider a prior distribution $\pi$ on $\mathcal{P}$ of the form
$\pi\coloneqq\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}.$
For $z_{1},\ldots,z_{n}\in\mathcal{Z}$, we may define a distribution on $\mathcal{P}$ by
$\pi(B\,|\,z_{1},\ldots,z_{n})\coloneqq\frac{\int_{B}\prod_{i=1}^{n}\frac{\mathrm{d}P}{\mathrm{d}\nu}(z_{i})\,\mathrm{d}\pi(P)}{\int_{\mathcal{P}}\prod_{i=1}^{n}\frac{\mathrm{d}P}{\mathrm{d}\nu}(z_{i})\,\mathrm{d}\pi(P)},$
for any measurable $B\subseteq\mathcal{P}$. Thus, $\pi(\cdot\,|\,Z_{1},\ldots,Z_{n})$ would be the posterior distribution when $P$ has prior distribution $\pi$ and $Z_{1},\ldots,Z_{n}\,|\,P\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P$. In the sequel, however, we will assume instead that $Z_{1},\ldots,Z_{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P_{0}$ for some fixed distribution $P_{0}$ on $\mathcal{Z}$, and seek bounds on the expected mass assigned by $\pi(\cdot\,|\,Z_{1},\ldots,Z_{n})$ to (Hellinger) balls around $P_{0}$.
For $\alpha\in\mathcal{A}$, a distribution $P_{0}$ on $\mathcal{Z}$ dominated by $\nu$, and for $\epsilon>0$, define
$B_{\alpha}(P_{0},\epsilon)\coloneqq\Bigl\{P\in\mathcal{P}_{\alpha}:\mathrm{KL}(P_{0},P)\leq\epsilon^{2},\,\mathrm{V}_{2}(P_{0},P)\leq\epsilon^{2}\Bigr\}$ (37)
and
$B^{\prime}_{\alpha}(P_{0},\epsilon)\coloneqq\Bigl\{P\in\mathcal{P}_{\alpha}:d_{\mathrm{H}}(P_{0},P)\leq\epsilon\Bigr\}.$ (38)
Let $(\epsilon_{n,\alpha})_{\alpha\in\mathcal{A}}$ be a positive sequence, thought of as the rate of convergence in Hellinger distance on each $\mathcal{P}_{\alpha}$. Fix $\beta\in\mathcal{A}$ and $H\geq 1$, and define
$\mathcal{A}_{\geq\beta}$ $\coloneqq\bigl\{\alpha\in\mathcal{A}:\epsilon_{n,\alpha}^{2}\leq H\epsilon_{n,\beta}^{2}\bigr\}$
$\mathcal{A}_{<\beta}$ $\coloneqq\bigl\{\alpha\in\mathcal{A}:\epsilon_{n,\alpha}^{2}>H\epsilon_{n,\beta}^{2}\bigr\},$
to be the sets of indices that are ‘at least as regular as $\beta$’ and ‘less regular than $\beta$’, respectively. Here $\beta$ should be thought of as the index of the ‘best model’ for $P_{0}$.
We now state five conditions for our posterior contraction theory:
  1. (C1)
    Covering number bound: There exists a positive sequence $(E_{\alpha})_{\alpha\in\mathcal{A}}$ such that
    $\sup_{\epsilon\geq\epsilon_{n,\alpha}}\log\mathcal{N}\bigl(\epsilon/3,\,B^{\prime}_{\alpha}(P_{0},2\epsilon),\,d_{\mathrm{H}}\bigr)\leq E_{\alpha}n\epsilon_{n,\alpha}^{2}\quad\text{for all }\alpha\in\mathcal{A}.$
  2. (C2)
    Prior mass lower bound for $\beta$: There exists $F>0$ such that
    $\pi_{\beta}\bigl(B_{\beta}(P_{0},\epsilon_{n,\beta})\bigr)\geq e^{-Fn\epsilon_{n,\beta}^{2}}.$
  3. (C3)
    Mixture weights: There exists a positive sequence $(\mu_{n,\alpha})_{\alpha\in\mathcal{A}}$ such that
    $\frac{\lambda_{\alpha}}{\lambda_{\beta}}\leq\mu_{n,\alpha}\,e^{n(H^{-1}\epsilon_{n,\alpha}^{2}\vee\epsilon_{n,\beta}^{2})}\quad\text{for all }\alpha\in\mathcal{A}.$
  4. (C4)
    Prior mass upper bound for $\mathcal{A}_{<\beta}$: There exist $M,G>0$ such that
    $\sum_{\alpha\in\mathcal{A}_{<\beta}}\frac{\lambda_{\alpha}}{\lambda_{\beta}}\cdot\pi_{\alpha}\bigl(B^{\prime}_{\alpha}(P_{0},M\epsilon_{n,\alpha})\bigr)\leq Ge^{-(F+3)n\epsilon_{n,\beta}^{2}}.$
  5. (C5)
    Uniformity of covering number constants: There exists $E\in(0,\infty)$ such that
    $E\geq\sup_{\alpha\in\mathcal{A}_{<\beta}}E_{\alpha}\vee\sup_{\alpha\in\mathcal{A}_{\geq\beta}}\frac{E_{\alpha}\epsilon_{n,\alpha}^{2}}{\epsilon_{n,\beta}^{2}}.$
The set $\mathcal{A}_{<\beta}$ may be empty; in that case, condition (C4) is automatically satisfied, and the supremum over the empty set in (C5) is defined as -$\infty$.
The following theorem is a modification of Ghosal, Lember and van der Vaart (2008, Theorem 2.1), where we use a slightly larger constant in the exponent in (C4) and impose an additional sub-Gamma assumption on the log-likelihood ratios in order to obtain exponential posterior contraction (as opposed to the weaker conclusion that the left-hand side of (39) converges to zero as $n\to\infty$). In our in-context learning problem, this is crucial to derive rates of convergence for the posterior regression function. The sub-Gamma assumption is verified in Lemma 24(c) under a nonparametric regression setting with independent Gaussian noise.
Proposition 16.
Let $P_{0}$ be a distribution on $\mathcal{Z}$, let $\beta\in\mathcal{A}$ and let $Z_{1},\ldots,Z_{n}\overset{\mathrm{iid}}{\sim}P_{0}$. Assume that the conditions (C1)(C5) hold with $M^{2}>243(F+1)(HE+1)$ and $\sum_{\alpha\in\mathcal{A}}\sqrt{\mu_{n,\alpha}}\leq e^{n\epsilon_{n,\beta}^{2}}$. For $i\in[n]$ and $P\in\mathcal{P}_{\beta}$ with with $P\ll P_{0}$, let
$W(P,Z_{i})\coloneqq\log\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i})-\mathbb{E}\log\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i}).$
Assume that there exist $a,c>0$ such that for all $P\in B_{\beta}(P_{0},\epsilon_{n,\beta})$, we have that $W(P,Z_{1})$ is sub-Gamma in the left tail with variance parameter $a\epsilon_{n,\beta}^{2}$ and scale parameter $c$. Then there exist $C_{1},c_{2},c_{3}>0$, not depending on $n$, such that if $n\epsilon_{n,\beta}^{2}\geq c_{3}$, then
$\mathbb{E}_{(Z_{1},\ldots,Z_{n})\sim P_{0}^{\otimes n}}\bigl[\pi\bigl(\{P\in\mathcal{P}:d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}\}\,\big|\,Z_{1},\ldots,Z_{n}\bigr)\bigr]\leq C_{1}e^{-c_{2}n\epsilon_{n,\beta}^{2}}.$ (39)
Proof.
Our assumptions ensure that the conditions of Ghosal, Lember and van der Vaart (2008, Theorem 2.1) hold [^6] As we seek a finite-sample bound, we do not require their condition that $n\epsilon_{n,\beta}^{2}\to\infty$ as $n\to\infty$. with $K=1/9$, $L=K/3=1/27$, $I=\sqrt{27(F+1)}$ and $B=M/I=M/\sqrt{27(F+1)}$ therein. Moreover, our conditions (C2) and (C4) ensure that
$\sum_{\alpha\in\mathcal{A}_{<\beta}}\frac{\lambda_{\alpha}}{\lambda_{\beta}}\cdot\frac{\pi_{\alpha}\bigl(B^{\prime}_{\alpha}(P_{0},M\epsilon_{n,\alpha})\bigr)}{\pi_{\beta}\bigl(B_{\beta}(P_{0},\epsilon_{n,\beta})\bigr)}\leq Ge^{-3n\epsilon_{n,\beta}^{2}}.$ (40)
Now, by following the proof of Ghosal, Lember and van der Vaart (2008, Theorem 2.1), replacing the use of their Eq. (2.4) by our (40), and the use of their Lemma 6.3 by our Lemma 23 (which applies because of our sub-Gamma assumption on $W(P,Z_{1})$), we deduce the desired conclusion. ∎
E.2 Bayesian nonparametric regression
Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$. In this section, we specialise to the setting of regression and take $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$. Let $P_{X}$ be a Borel distribution on $\mathcal{X}$, let $X\sim P_{X}$ and let $\xi\sim N(0,\sigma^{2})$ be independent of $X$. Let $R>0$, and for $\alpha\in\mathcal{A}$, let $\mathcal{G}_{\alpha}$ be a space of measurable functions from $\mathcal{X}$ to [R,R][-R,R]. For any Borel measurable function $g:\mathcal{X}\to\mathbb{R}$, we write $P_{g}$ for the distribution of $(X,Y_{g})$ where $Y_{g}=g(X)+\xi$. For $\alpha\in\mathcal{A}$, let $\widetilde{\pi}_{\alpha}$ be a distribution on $\mathcal{G}_{\alpha}$, let $\widetilde{g}_{\alpha}\sim\widetilde{\pi}_{\alpha}$ and let $\pi_{\alpha}$ denote the distribution of the random measure $P_{\widetilde{g}_{\alpha}}$, where the randomness comes from $\widetilde{g}_{\alpha}\sim\widetilde{\pi}_{\alpha}$. Finally, let $\widetilde{\pi}\coloneqq\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\widetilde{\pi}_{\alpha}$ and $\pi\coloneqq\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$, where $\lambda_{\alpha}\geq 0$ for all $\alpha\in\mathcal{A}$ and $\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}=1$, let $\mathcal{P}$ be the support of $\pi$ and $\mathcal{G}\coloneqq\bigcup_{\alpha\in\mathcal{A}}\mathcal{G}_{\alpha}$.
We first provide an alternative expression for the $\pi$-posterior regression function $g_{\pi}$, defined in (4). For $P\in\mathcal{P}$, let $g_{P}:\mathcal{X}\to\mathcal{Y}$ be the regression function under $P$, so that $g_{P}(x)\coloneqq\mathbb{E}_{(X,Y)\sim P}(Y\,|\,X=x)$ for $x\in\mathcal{X}$. Then the $\pi$-posterior regression function $g_{\pi}$ satisfies
$g_{\pi}(x,D_{n})$ $=\mathbb{E}_{P\sim\pi,\,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\bigl\{\mathbb{E}(Y\,|\,X=x,\,P)\,|\,X=x,\,\mathcal{D}_{n}=D_{n}\bigr\}$
$=\mathbb{E}_{P\sim\pi,\,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\bigl(g_{P}(x)\,|\,\mathcal{D}_{n}=D_{n}\bigr)$
$=\int_{\mathcal{P}}g_{P}(x)\,\mathrm{d}\pi(P\,|\,D_{n})=\int_{\mathcal{G}}g(x)\,\mathrm{d}\widetilde{\pi}(g\,|\,D_{n}),$ (41)
so $g_{\pi}(\cdot,D_{n})$ can be thought of as the posterior mean of the regression function.
Corollary 17.
Let $\beta\in\mathcal{A}$, let $g^{0}:\mathcal{X}\to[-R,R]$ be measurable and write $P_{0}\equiv P_{g^{0}}$. Assume that the conditions (C1)(C5) hold with $M^{2}>243(F+1)(HE+1)$ and $\sum_{\alpha\in\mathcal{A}}\sqrt{\mu_{n,\alpha}}\leq e^{n\epsilon_{n,\beta}^{2}}$. There exist $C_{1},C_{2},c_{3},c_{4}>0$, not depending on $n$, such that if $n\epsilon_{n,\beta}^{2}\geq c_{4}$, then
$\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{0}(X)\bigr\}^{2}\bigr]\leq C_{1}\epsilon_{n,\beta}^{2}+C_{2}e^{-c_{3}n\epsilon_{n,\beta}^{2}}.$
Proof.
We write $\mathbb{E}_{P_{0}}$ for $\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}$ and $C_{1}^{\prime}\coloneqq 2R^{2}/\bigl(1-e^{-R^{2}/(2\sigma^{2})}\bigr)$. We have by (41) that
$\mathbb{E}_{P_{0}}\bigl[\bigl\{g_{\pi}(X$ $,\mathcal{D}_{n})-g^{0}(X)\bigr\}^{2}\bigr]=\mathbb{E}_{P_{0}}\biggl[\biggl\{\int_{\mathcal{P}}g_{P}(X)\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})-g^{0}(X)\biggr\}^{2}\biggr]$
$\hskip-7.00032pt\overset{(i)}{\leq}\mathbb{E}_{P_{0}}\biggl[\int_{\mathcal{P}}\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
$\hskip-8.72289pt\overset{(ii)}{=}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl(\int_{\mathcal{P}}\mathbb{E}_{(X,Y)\sim P_{0}}\bigl[\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\bigr]\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr)$
$\hskip-10.44545pt\overset{(iii)}{\leq}C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl[\int_{\mathcal{P}}d_{\mathrm{H}}^{2}(P,P_{0})\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
$\hskip-9.60333pt\overset{(iv)}{\leq}C_{1}^{\prime}M^{2}\epsilon_{n,\beta}^{2}+2C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\bigl[\pi\bigl(\{P\in\mathcal{P}:d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}\}\,\big|\,\mathcal{D}_{n}\bigr)\bigr]$
$\hskip-7.88077pt\overset{(v)}{\leq}C_{1}\epsilon_{n,\beta}^{2}+C_{2}e^{-c_{3}n\epsilon_{n,\beta}^{2}}.$
Here, $(i)$ follows from Jensen’s inequality, $(ii)$ follows by Fubini’s theorem, $(iii)$ follows from Lemma 24(d), $(iv)$ follows by splitting the integral into regions where $d_{\mathrm{H}}(P,P_{0})<M\epsilon_{n,\beta}$ and $d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}$ and $(v)$ follows from Proposition 16 where the sub-Gamma assumption is verified in Lemma 24(c). ∎
Appendix F Proofs for Section 4
F.1 Proof of Proposition 4
We begin with two preliminary lemmas in the setting of Section 4.1. Suppose that $g^{\circ}\in\mathcal{G}_{\beta}$ has wavelet decomposition
$g^{\circ}=\sum_{k\in K}C_{0}2^{-\ell_{0}d/2}a_{k}^{\circ}\Phi_{k}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}C_{0}2^{-\ell(\beta+d/2)}b_{\ell,\gamma}^{\circ}\Psi_{\ell,\gamma},$ (42)
where $(a_{k}^{\circ})_{k\in K}$ and $(b_{\ell,\gamma}^{\circ})_{\ell\in[\ell_{0}:\infty),\gamma\in\Gamma_{\ell}}$ are deterministic and satisfy $|a^{\circ}_{k}|\leq 1$ and $|b^{\circ}_{\ell,\gamma}|\leq 1$ for $k\in K$, $\ell\in[\ell_{0}:\infty)$ and $\gamma\in\Gamma_{\ell}$. Lemma 18 below provides a lower bound on the prior mass; similar results for the one-dimensional case exist in the literature, e.g. Giné and Nickl (2011, Proposition 1) and Reiß and Schmidt-Hieber (2020, Lemma 4.1).
Lemma 18.
Let $n\in\mathbb{N}$ and assume that (42) holds. Then there exists $C>0$, depending only on $\beta,d,c_{0},C_{0}$ and $S$, such that
$\widetilde{\pi}_{\beta}\bigl(\bigl\{g\in\mathcal{G}_{\beta}:\|g-g^{\circ}\|_{\infty}^{2}\leq Cn^{-2\beta/(2\beta+d)}\bigr\}\bigr)\geq\exp\bigl(-Cn^{d/(2\beta+d)}\bigr).$ (43)
Moreover, there exists $C^{\prime}>0$, depending only on $\beta,d,c_{0},C_{0},S$ and $\sigma$, such that
$\pi_{\beta}\bigl(B_{\beta}(P_{g^{\circ}},\epsilon_{n,\beta})\bigr)\geq\exp\bigl(-C^{\prime}n\epsilon_{n,\beta}^{2}\bigr),$ (44)
where $\epsilon_{n,\beta}\coloneqq C^{\prime}n^{-\beta/(2\beta+d)}$ and $B_{\beta}(P_{g^{\circ}},\epsilon_{n,\beta})$ is defined by (37).
Proof.
Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that may depend only on $\beta,d,c_{0},C_{0}$ and $S$, while $C_{1}^{\prime},C_{2}^{\prime},\ldots$ may in addition depend on $\sigma$. Let $\widetilde{g}_{\beta}$ be given by (8) so that $\widetilde{g}_{\beta}\sim\widetilde{\pi}_{\beta}$, and write $a_{k}\equiv a_{k}^{(\beta)}$ and $b_{\ell,\gamma}\equiv b_{\ell,\gamma}^{(\beta)}$ where the random variables $(a_{k}^{(\beta)})$ and $(b_{\ell,\gamma}^{(\beta)})$ satisfy the conditions below (8). By (32) and (33), writing $L\coloneqq\bigl\lceil\frac{\log_{2}n}{2\beta+d}\bigr\rceil$ (and assuming for now that $L\geq\ell_{0}$), we have
$\|\widetilde{g}_{\beta}-g^{\circ}\|_{\infty}$ $\leq C_{1}\max_{k\in K}|a_{k}-a^{\circ}_{k}|+C_{1}\sum_{\ell=\ell_{0}}^{L}2^{-\ell\beta}\max_{\gamma\in\Gamma_{\ell}}|b_{\ell,\gamma}-b_{\ell,\gamma}^{\circ}|+2C_{0}C_{1}\sum_{\ell=L+1}^{\infty}2^{-\ell\beta}$
$\leq C_{1}\max_{k\in K}|a_{k}-a^{\circ}_{k}|+C_{1}\sum_{\ell=\ell_{0}}^{L}2^{-\ell\beta}\max_{\gamma\in\Gamma_{\ell}}|b_{\ell,\gamma}-b_{\ell,\gamma}^{\circ}|+C_{3}n^{-\beta/(2\beta+d)}.$ (45)
Define the events
$\mathcal{E}_{1}$ $\coloneqq\bigcap_{k\in K}\Bigl\{|a_{k}-a^{\circ}_{k}|\leq 2\cdot n^{-\beta/(2\beta+d)}\Bigr\},$
$\mathcal{E}_{2}$ $\coloneqq\bigcap_{\ell=\ell_{0}}^{L}\bigcap_{\gamma\in\Gamma_{\ell}}\Bigl\{|b_{\ell,\gamma}-b_{\ell,\gamma}^{\circ}|\leq 2\cdot 2^{-2(L-\ell)\beta}\Bigr\}.$
Then, on the event $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, we have by (45) that
$\|\widetilde{g}_{\beta}-g^{\circ}\|_{\infty}$ $\leq 2C_{1}n^{-\beta/(2\beta+d)}+2C_{1}2^{-L\beta}\sum_{\ell=\ell_{0}}^{L}2^{-(L-\ell)\beta}+C_{3}n^{-\beta/(2\beta+d)}\leq C_{4}n^{-\beta/(2\beta+d)}.$
Therefore,
$\mathbb{P}\bigl(\|\widetilde{g}_{\beta}-g^{\circ}\|_{\infty}\leq C_{4}n^{-\beta/(2\beta+d)}\bigr)$ $\geq\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})=\mathbb{P}(\mathcal{E}_{1})\mathbb{P}(\mathcal{E}_{2}).$ (46)
Now, by (31), we have
$\mathbb{P}(\mathcal{E}_{1})\geq\bigl(c_{0}n^{-\beta/(2\beta+d)}\bigr)^{2^{\ell_{0}d}}=\exp(-C_{5}\log n).$ (47)
Similarly,
$\mathbb{P}(\mathcal{E}_{2})>\prod_{\ell=\ell_{0}}^{L}\bigl(c_{0}2^{-2(L-\ell)\beta}\bigr)^{2^{(\ell+1)d}}$ $=\exp\biggl((\log c_{0})\sum_{\ell=\ell_{0}}^{L}2^{(\ell+1)d}-(2^{d+1}\log 2)\beta\sum_{\ell=\ell_{0}}^{L}(L-\ell)2^{\ell d}\biggr)$
$\geq\exp(-C_{6}n^{d/(2\beta+d)}),$ (48)
where the final inequality follows since if we set $A\coloneqq\sum_{\ell=\ell_{0}}^{L}(L-\ell)2^{\ell d}$, then $2^{d}A-A\leq\sum_{\ell=0}^{L}2^{\ell d}\leq 2^{(L+1)d}/(2^{d}-1)$, so $A\leq 2^{(L+1)d}/(2^{d}-1)^{2}$. The conclusion (43) for $n$ large enough that $L\geq\ell_{0}$ follows from (46), (47) and (48). By increasing $C>0$, depending only on $\beta,d,c_{0},C_{0}$ and $S$, we can then ensure that (43) holds for all $n\in\mathbb{N}$.
For (44), note that by Lemma 24(a) and (b), we have that if $g\in\mathcal{G}_{\beta}$ satisfies $\|g-g^{\circ}\|_{\infty}^{2}\leq Cn^{-2\beta/(2\beta+d)}$, then
$\mathrm{KL}(P_{g^{\circ}},P_{g})\leq C_{1}^{\prime}Cn^{-2\beta/(2\beta+d)}\quad\text{and}\quad\mathrm{V}_{2}(P_{g^{\circ}},P_{g})\leq C_{1}^{\prime}Cn^{-2\beta/(2\beta+d)}.$
Writing $\epsilon_{n,\beta}\coloneqq\sqrt{C_{1}^{\prime}C}\cdot n^{-\beta/(2\beta+d)}$, we deduce from (43) that
$\pi_{\beta}\bigl(B_{\beta}(P_{g^{\circ}},\epsilon_{n,\beta})\bigr)\geq\exp\bigl(-Cn^{d/(2\beta+d)}\bigr)\geq\exp(-C_{2}^{\prime}n\epsilon_{n,\beta}^{2}).$
Thus (44) holds with $C^{\prime}\coloneqq\sqrt{C_{1}^{\prime}C}\vee C_{2}^{\prime}$. ∎
Lemma 19 below provides an upper bound on the prior mass.
Lemma 19.
Let $\alpha\in\mathcal{A}$, $m>0$ and assume that (42) holds. Suppose that $P_{X}$ is a Borel distribution on $0,1]^{d}$ with the property that there exist $c_{0}>0$ and a hypercube $A\subseteq[0,1]^{d}$ of side length $\tau>0$ such that $P_{X}(A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0})$ for all measurable $A_{0}\subseteq A$. Then there exist $c\equiv c(\alpha,d,c_{0},C_{0},\tau)>0$ and $n_{0}\equiv n_{0}(\alpha,d,c_{0},C_{0},S,m,\tau)\in\mathbb{N}$ such that
$\widetilde{\pi}_{\alpha}\Bigl(\Bigl\{g\in\mathcal{G}_{\alpha}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq\frac{c_{0}}{4}m^{2}n^{-2\alpha/(2\alpha+d)}\Bigr\}\Bigr)\leq\exp\bigl(-cm^{-d/\alpha}n^{d/(2\alpha+d)}\bigr)$ (49)
for $n\geq n_{0}$. Moreover, there exists $c^{\prime}>0$, depending only on $\alpha,\beta,d,c_{0},C_{0},S,\tau$ and $\sigma$, such that for $n\geq n_{0}$,
$\pi_{\alpha}\bigl(B^{\prime}_{\alpha}(P_{g^{\circ}},c^{\prime}mn^{-\alpha/(2\alpha+d)})\bigr)\leq\exp\bigl(-cm^{-d/\alpha}n^{d/(2\alpha+d)}\bigr)$ (50)
where $B^{\prime}_{\alpha}(P_{g^{\circ}},c^{\prime}mn^{-\alpha/(2\alpha+d)})$ is defined as in (38).
Proof.
Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that may depend only on $\alpha,d,c_{0},C_{0}$ and $\tau$, while $C_{1}^{\prime},C_{2}^{\prime},\ldots$ may depend in addition on $\beta,S$ and $\sigma$. Let $\epsilon\coloneqq n^{-\alpha/(2\alpha+d)}$. Fix $g^{\prime}\in\mathcal{G}_{\alpha}$ such that $\|g^{\prime}-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq\frac{c_{0}}{4}m^{2}\epsilon^{2}$; we may assume without loss of generality that such a $g^{\prime}$ exists, since otherwise the left-hand side of (49) would be zero. Suppose that $g^{\prime}$ has wavelet decomposition
$g^{\prime}=\sum_{k\in K}C_{0}2^{-\ell_{0}d/2}a_{k}^{\prime}\Phi_{k}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}C_{0}2^{-\ell(\alpha+d/2)}b_{\ell,\gamma}^{\prime}\Psi_{\ell,\gamma},$
where $|a_{k}^{\prime}|\leq 1$ and $|b_{\ell,\gamma}^{\prime}|\leq 1$. Note that if $g\in\mathcal{G}_{\alpha}$ satisfies $\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq\frac{c_{0}}{4}m^{2}\epsilon^{2}$, then $\|g-g^{\prime}\|_{L^{2}(P_{X})}^{2}\leq 2\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}+2\|g^{\prime}-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq c_{0}m^{2}\epsilon^{2}$. Thus,
$\widetilde{\pi}_{\alpha}\Bigl(\Bigl\{g\in\mathcal{G}_{\alpha}:\|g$ $-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq\frac{c_{0}}{4}m^{2}\epsilon^{2}\Bigr\}\Bigr)\leq\widetilde{\pi}_{\alpha}\Bigl(\Bigl\{g\in\mathcal{G}_{\alpha}:\|g-g^{\prime}\|_{L^{2}(P_{X})}^{2}\leq c_{0}m^{2}\epsilon^{2}\Bigr\}\Bigr).$ (51)
Let $\widetilde{g}_{\alpha}\sim\widetilde{\pi}_{\alpha}$, and write $a_{k}\equiv a_{k}^{(\alpha)}$ and $b_{\ell,\gamma}\equiv b_{\ell,\gamma}^{(\alpha)}$ with representation (8), where the random variables $(a_{k}^{(\alpha)})$ and $(b_{\ell,\gamma}^{(\alpha)})$ satisfy the conditions below (8). Further, let $C_{1}\coloneqq\lceil\log_{2}(2/\tau)\rceil$. Then, by the assumption on $P_{X}$, there exists $C_{2}\in[0:2^{C_{1}}-1]$ such that $\bigl[\frac{C_{2}}{2^{C_{1}}},\frac{C_{2}+1}{2^{C_{1}}}\bigr]^{d}\subseteq A$. Choose $n_{0}\equiv n_{0}(\alpha,d,c_{0},C_{0},S,m,\tau)\in\mathbb{N}$ large enough that $L\coloneqq\bigl\lfloor\frac{1}{2\alpha}\log_{2}\bigl(\frac{c_{0}C_{0}^{2}}{6m^{2}\epsilon^{2}2^{C_{1}d}}\bigr)\bigr\rfloor$ is at least $\ell_{0}\vee C_{1}$ for $n\geq n_{0}$. Assume henceforth that $n\geq n_{0}$. Let $\Gamma_{L}^{\prime}\coloneqq\bigl\{\gamma\in\Gamma_{L}:\mathrm{supp}(\Psi_{L,\gamma})\subseteq\bigl[\frac{C_{2}}{2^{C_{1}}},\frac{C_{2}+1}{2^{C_{1}}}\bigr]^{d}\bigr\}$, $b_{L}\coloneqq(b_{L,\gamma})_{\gamma\in\Gamma_{L}^{\prime}}$ and $b_{L}^{\prime}\coloneqq(b_{L,\gamma}^{\prime})_{\gamma\in\Gamma_{L}^{\prime}}$. Since $\mathrm{supp}(\Psi_{L,\gamma})$ is contained in a dyadic cube of side length $2^{-L}$ (see Section D.2), we deduce that $|\Gamma_{L}^{\prime}|\geq 2^{(L-C_{1})d}$ and that
$C_{0}^{2}2^{-L(2\alpha+d)}\mathbb{E}\bigl(\|b_{L}-b_{L}^{\prime}\|_{2}^{2}\bigr)$ $\geq C_{0}^{2}2^{-L(2\alpha+d)}\mathrm{tr}\bigl(\mathrm{Cov}(b_{L})\bigr)$
$\geq|\Gamma_{L}^{\prime}|C_{0}^{2}2^{-L(2\alpha+d)}\cdot\frac{c_{0}}{3}\geq\frac{c_{0}C_{0}^{2}2^{-2\alpha L-C_{1}d}}{3}\geq 2m^{2}\epsilon^{2}.$
Therefore, by the assumption on $P_{X}$ and Lemma 25, we have
$\mathbb{P}\bigl(\|\widetilde{g}_{\alpha}-g^{\prime}\|_{L^{2}(P_{X})}^{2}\leq c_{0}$ $m^{2}\epsilon^{2}\bigr)\leq\mathbb{P}\bigl(c_{0}\|\widetilde{g}_{\alpha}-g^{\prime}\|_{L^{2}([C_{2}2^{-C_{1}},(C_{2}+1)2^{-C_{1}}]^{d})}^{2}\leq c_{0}m^{2}\epsilon^{2}\bigr)$
$\leq\mathbb{P}\biggl\{\|b_{L}-b_{L}^{\prime}\|_{2}^{2}-\mathbb{E}\bigl(\|b_{L}-b_{L}^{\prime}\|_{2}^{2}\bigr)\leq\frac{m^{2}\epsilon^{2}}{C_{0}^{2}2^{-L(2\alpha+d)}}-\mathbb{E}\bigl(\|b_{L}-b_{L}^{\prime}\|_{2}^{2}\bigr)\biggr\}$
$\leq\mathbb{P}\biggl\{\|b_{L}-b_{L}^{\prime}\|_{2}^{2}-\mathbb{E}\bigl(\|b_{L}-b_{L}^{\prime}\|_{2}^{2}\bigr)\leq-\frac{m^{2}\epsilon^{2}}{C_{0}^{2}2^{-L(2\alpha+d)}}\biggr\}.$ (52)
Since $(b_{L,\gamma}-b_{L,\gamma}^{\prime})^{2}\in[0,4]$ for $\gamma\in\Gamma_{L}^{\prime}$, we have by Hoeffding’s inequality and (31) that
$\mathbb{P}\biggl\{\|b_{L}$ $-b_{L}^{\prime}\|_{2}^{2}-\mathbb{E}\bigl(\|b_{L}-b_{L}^{\prime}\|_{2}^{2}\bigr)\leq-\frac{m^{2}\epsilon^{2}}{C_{0}^{2}2^{-L(2\alpha+d)}}\biggr\}\leq\exp\biggl(-\frac{m^{4}\epsilon^{4}}{8C_{0}^{4}2^{-(4\alpha+d)L-C_{1}d}}\biggr)$
$\leq\exp\biggl\{-\frac{c_{0}^{2}}{288\cdot 2^{4\alpha+(C_{1}+1)d}}\biggl(\frac{c_{0}C_{0}^{2}}{6m^{2}\epsilon^{2}2^{C_{1}d}}\biggr)^{d/(2\alpha)}\biggr\}\eqqcolon\exp\bigl(-cm^{-d/\alpha}n^{d/(2\alpha+d)}\bigr),$ (53)
where $c>0$ depends only on $\alpha,d,c_{0},C_{0}$ and $\tau$. Combining (51), (52) and (53) yields (49).
For the second claim, note that by (32) and (33),
$\|g^{\circ}\|_{\infty}\leq C_{3}\max_{k\in K}|a_{k}^{\circ}|+C_{3}\sum_{\ell=\ell_{0}}^{\infty}2^{-\ell\beta}\max_{\gamma\in\Gamma_{\ell}}|b_{\ell,\gamma}^{\circ}|\leq C_{1}^{\prime},$ (54)
and similarly, $\|\widetilde{g}_{\alpha}\|_{\infty}\leq C_{2}^{\prime}$ almost surely. Thus, if $d_{\mathrm{H}}(P_{g^{\circ}},P_{g})\leq c^{\prime}mn^{-\alpha/(2\alpha+d)}$ for some $c^{\prime}>0$, then by Lemma 24(d),
$\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq C_{3}^{\prime}d_{\mathrm{H}}^{2}(P_{g^{\circ}},P_{g})\leq C_{3}^{\prime}(c^{\prime})^{2}m^{2}n^{-2\alpha/(2\alpha+d)}.$
Therefore, taking $c^{\prime}\coloneqq\sqrt{c_{0}/(4C_{3}^{\prime})}$ and applying (49) proves (50). ∎
We are now in a position to prove Proposition 4.
Proof of Proposition 4.
Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that do not depend on $n$. By an argument similar to (54), there exists $C_{1}>0$ such that $\|g^{\circ}\|_{\infty}\leq C_{1}$ and $\|g\|_{\infty}\leq C_{1}$ for all $g\in\mathcal{G}$. We now verify the assumptions of Corollary 17 with $H=1$. For each $\alpha\in\mathcal{A}$, let $C^{\prime}>0$ (which does not depend on $n$) be the quantity in Lemma 18 such that (44) holds, and let $\epsilon_{n,\alpha}\coloneqq C^{\prime}n^{-\alpha/(2\alpha+d)}$ for $\alpha\in\mathcal{A}$. Then by Lemma 18, condition (C2) holds with $F=C^{\prime}$. By Lemma 24(d), (9) and Giné and Nickl (2021, Eq. (4.184)), for each $\alpha\in\mathcal{A}$, we have
$\sup_{\epsilon\geq\epsilon_{n,\alpha}}\log\mathcal{N}\bigl(\epsilon/3,\,B^{\prime}_{\alpha}(P_{0},2\epsilon),\,d_{\mathrm{H}}\bigr)$ $\leq\log\mathcal{N}\bigl(\epsilon_{n,\alpha}/3,\,\mathcal{P}_{\alpha},\,d_{\mathrm{H}}\bigr)$
$\leq\log\mathcal{N}\bigl(2\sigma\epsilon_{n,\alpha}/3,\,\mathcal{G}_{\alpha},\,\|\cdot\|_{\infty}\bigr)$
$\leq\log\mathcal{N}\bigl(2\sigma\epsilon_{n,\alpha}/3,\,B^{\alpha}_{\infty,\infty}([0,1]^{d},2C_{0}),\,\|\cdot\|_{\infty}\bigr)$
$\leq E_{\alpha}n\epsilon_{n,\alpha}^{2},$
for some $E_{\alpha}>0$ not depending on $n$; this verifies condition (C1). Moreover, since $\mathcal{A}$ is finite, condition (C3) is satisfied with $\mu_{n,\alpha}=1$ for all $\alpha\in\mathcal{A}$ when $n$ is sufficiently large. Similarly, condition (C5) is satisfied for some $E>0$ not depending on $n$. Finally, take $M>0$ such that $M^{2}>243(F+1)(HE+1)$. Letting $c,c^{\prime}>0$ be the quantities in Lemma 19 and applying Lemma 19 with $m\coloneqq MC^{\prime}/c^{\prime}$, we deduce that for $n$ sufficiently large,
$\sum_{\alpha\in\mathcal{A}_{<\beta}}\frac{\lambda_{\alpha}}{\lambda_{\beta}}\cdot\pi_{\alpha}\bigl(B^{\prime}_{\alpha}(P_{0},M\epsilon_{n,\alpha})\bigr)\leq\sum_{\alpha\in\mathcal{A}:\alpha<\beta}\frac{\lambda_{\alpha}}{\lambda_{\beta}}\cdot\exp\bigl(-cm^{-d/\alpha}n^{d/(2\alpha+d)}\bigr)\leq e^{-(F+3)n\epsilon_{n,\beta}^{2}},$
which verifies condition (C4). Therefore, by Corollary 17, the claim follows for sufficiently large $n$, since $e^{-c_{3}n\epsilon_{n,\beta}^{2}}\leq\epsilon_{n,\beta}^{2}$ for $n$ large enough. Moreover, since $g_{\pi}$ and $g^{\circ}$ are bounded, the conclusion now follows for all $n\in\mathbb{N}$ by increasing $C>0$ in the statement if necessary. ∎
F.2 Proof of Proposition 6
We work in the setting of Section 4.2. For $(\alpha,p)\in\mathcal{A}^{\prime}$, we define $\mathcal{G}_{\alpha}^{(p)}$ similarly to the definition of $\mathcal{G}_{\alpha}$ in Section 4.1 but replacing the dimensionality $d$ by $p$, i.e. we let $\mathcal{G}_{\alpha}^{(p)}$ be the set of all functions $h_{\alpha}^{(p)}:[0,1]^{p}\to\mathbb{R}$ such that $h$ has wavelet decomposition
$h_{\alpha}^{(p)}=\sum_{k\in K}C_{0}2^{-\ell_{0}r/2}a_{k}\Phi_{k}^{(p)}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}C_{0}2^{-\ell(\beta+r/2)}b_{\ell,\gamma}\Psi_{\ell,\gamma}^{(p)},$
where $(\Phi_{k}^{(p)})$ and $(\Psi_{\ell,\gamma}^{(p)})$ is the tensor product CDV wavelet basis for $L^{2}([0,1]^{p})$, $|a_{k}|\leq 1$ and $|b_{\ell,\gamma}|\leq 1$. Thus, similarly to (9),
$B_{\infty,\infty}^{\alpha}([0,1]^{p},C_{0})\subseteq\mathcal{G}_{\alpha}^{(p)}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0}).$
The proof of Proposition 6 relies on three preliminary lemmas that we now formulate. We may write
$g^{\circ}(x)=h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr),$ (55)
where $U^{\circ}\in V_{r}(\mathbb{R}^{d})$ and $h^{\circ}\in\mathcal{G}_{\beta}^{(r)}$ has wavelet decomposition
$h^{\circ}=\sum_{k\in K}C_{0}2^{-\ell_{0}r/2}a_{k}^{\circ}\Phi_{k}^{(r)}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}C_{0}2^{-\ell(\beta+r/2)}b_{\ell,\gamma}^{\circ}\Psi_{\ell,\gamma}^{(r)},$ (56)
where $|a^{\circ}_{k}|\leq 1$ and $|b^{\circ}_{\ell,\gamma}|\leq 1$.
Lemma 20.
Let $n\in\mathbb{N}$ and assume that (55) and (56) hold. Then there exists $C>0$, depending only on $\beta,d,r,c_{0},C_{0}$ and $S$, such that
$\widetilde{\pi}_{\beta,r}\bigl(\bigl\{g\in\mathcal{G}_{\beta,r}:\|g-g^{\circ}\|_{\infty}^{2}\leq Cn^{-2\beta/(2\beta+r)}\bigr\}\bigr)\geq\exp\bigl(-Cn^{r/(2\beta+r)}\bigr).$ (57)
Moreover, there exists $C^{\prime}>0$, depending only on $\beta,d,r,c_{0},C_{0},S$ and $\sigma$, such that if $\epsilon_{n,\beta,r}\coloneqq C^{\prime}n^{-\beta/(2\beta+r)}$, then
$\pi_{\beta,r}\bigl(B_{\beta,r}(P_{g^{\circ}},\epsilon_{n,\beta,r})\bigr)\geq\exp\bigl(-C^{\prime}n\epsilon_{n,\beta,r}^{2}\bigr),$ (58)
where $B_{\beta,r}(P_{g^{\circ}},\epsilon_{n,\beta,r})\coloneqq\bigl\{P\in\mathcal{P}_{\beta,r}:\mathrm{KL}(P_{g^{\circ}},P)\leq\epsilon_{n,\beta,r}^{2},\,\mathrm{V}_{2}(P_{g^{\circ}},P)\leq\epsilon_{n,\beta,r}^{2}\bigr\}$, similarly to (37).
Proof.
Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that may depend only on $\beta,d,r,c_{0},C_{0}$ and $S$, while $C_{1}^{\prime},C_{2}^{\prime},\ldots$ may in addition depend on $\sigma$. Let $\widetilde{g}_{\beta,r}(x)=\widetilde{g}_{\beta}^{(r)}\bigl(\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}\bigr)$ be defined as in (11) so that $\widetilde{g}_{\beta,r}\sim\widetilde{\pi}_{\beta,r}$. Observe that $B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})\subseteq B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$ (e.g. Giné and Nickl, 2021, Proposition 4.3.10(ii)). When $\beta<1$, the space $B_{\infty,\infty}^{\beta\wedge 1}$ is equal to the Hölder space $H^{\beta}$ with equivalent norms (e.g. Giné and Nickl, 2021, Proposition 4.3.23), so for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$, we have $|g(x)-g(y)|\leq C_{1}\|x-y\|_{2}^{\beta}$; when $\beta\geq 1$, $B_{\infty,\infty}^{\beta\wedge 1}$ is equal to the Zygmund space of order 11 with equivalent norms (Triebel, 1983, p. 113), so by Anderson and Pitt (1989, Proposition 2.4) (see also Zygmund, 2002, Chapter II, Theorem 3.4), for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$, we have $|g(x)-g(y)|\leq C_{2}\|x-y\|_{2}\bigl\{\log(1/\|x-y\|_{2})\vee 1\bigr\}$. Therefore, for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$,
$|g(x)-g(y)|\leq C_{3}\|x-y\|_{2}^{\beta\wedge 1}\bigl\{\log(1/\|x-y\|_{2})\vee 1\bigr\}.$ (59)
On the event $E\coloneqq\bigl\{\|U^{(r)}-U^{\circ}\|_{\mathrm{op}}\leq\bigl(n^{-\beta/(2\beta+r)}/\log(en)\bigr)^{1/(\beta\wedge 1)}\bigr\}$, we have for all $x\in\mathbb{B}^{d}$ that
$|\widetilde{g}_{\beta,r}(x)-g^{\circ}(x)|$
$\leq\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}\biggr)-\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)\biggr|+\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)\biggr|$
$\leq C_{4}n^{-\beta/\{(2\beta+r)\}}+\|\widetilde{g}_{\beta}^{(r)}-h^{\circ}\|_{\infty},$ (60)
where the second inequality follows from (59), since $\widetilde{g}_{\beta}^{(r)}\in\mathcal{G}_{\beta}^{(r)}\subseteq B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$ and $\bigl\|\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}-\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\bigr\|_{2}\leq\frac{1}{2}\bigl(n^{-\beta/(2\beta+r)}/\log(en)\bigr)^{1/(\beta\wedge 1)}$ on $E$. Now by Lemma 26, we have
$\mathbb{P}(E)\geq C_{5}\biggl(\frac{n^{-\beta/(2\beta+r)}}{\log(en)}\wedge\frac{1}{2}\biggr)^{dr/(\beta\wedge 1)}\geq\exp(-C_{6}n^{r/(2\beta+r)}),$
and by (43) (with $d$ replaced by $r$ therein), we have
$\mathbb{P}\bigl(\|\widetilde{g}_{\beta}^{(r)}-h^{\circ}\|_{\infty}\leq C_{7}n^{-\beta/(2\beta+r)}\bigr)\geq\exp(-C_{7}n^{r/(2\beta+r)}).$
Hence,
$\mathbb{P}\bigl(\|\widetilde{g}_{\beta,r}-g^{\circ}\|_{\infty}^{2}$ $\leq(C_{4}+C_{7})^{2}n^{-2\beta/(2\beta+r)}\bigr)$
$\geq\mathbb{P}\bigl(\|\widetilde{g}_{\beta,r}-g^{\circ}\|_{\infty}\leq(C_{4}+C_{7})n^{-\beta/(2\beta+r)}\,\big|\,E\bigr)\mathbb{P}(E)$
$\geq\mathbb{P}\bigl(\|\widetilde{g}_{\beta}^{(r)}-h^{\circ}\|_{\infty}\leq C_{7}n^{-\beta/(2\beta+r)}\bigr)\mathbb{P}(E)\geq\exp\bigl\{-(C_{6}+C_{7})n^{r/(2\beta+r)}\bigr\},$
where the second inequality follows from (60) and the fact that $\widetilde{g}_{\beta}^{(r)}$ and $U^{(r)}$ are independent. This proves (57) with $C\coloneqq(C_{4}+C_{7})^{2}\vee(C_{6}+C_{7})$.
The proof of (58) follows the same argument as the proof of (44). By Lemma 24(a) and (b), if $g\in\mathcal{G}_{\beta,r}$ satisfies $\|g-g^{\circ}\|_{\infty}^{2}\leq Cn^{-2\beta/(2\beta+r)}$, then
$\mathrm{KL}(P_{g^{\circ}},P_{g})\leq C_{1}^{\prime}Cn^{-2\beta/(2\beta+r)}\quad\text{and}\quad\mathrm{V}_{2}(P_{g^{\circ}},P_{g})\leq C_{2}^{\prime}Cn^{-2\beta/(2\beta+r)}.$
Writing $C^{\prime}\coloneqq C^{1/3}\vee\sqrt{(C_{1}^{\prime}\vee C_{2}^{\prime})C}$ and $\epsilon_{n,\beta,r}\coloneqq C^{\prime}n^{-\beta/(2\beta+r)}$, we deduce from (57) that
$\pi_{\beta,r}\bigl(B_{\beta,r}(P_{g^{\circ}},\epsilon_{n,\beta,r})\bigr)\geq\exp\bigl(-Cn^{r/(2\beta+r)}\bigr)\geq\exp(-C^{\prime}n\epsilon_{n,\beta,r}^{2}).$
This proves the claim. ∎
Lemma 21.
Let $(\alpha,p)\in\mathcal{A}$, $m>0$ and assume that (55) and (56) hold. Suppose that $P_{X}$ is a Borel distribution on $\mathbb{B}^{d}$ with the property that there exist $c_{0}>0$ and a closed Euclidean ball $A\subseteq\mathbb{B}^{d}$ of radius $\tau>0$ such that $P_{X}(A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0})$ for all measurable $A_{0}\subseteq A$. Then there exist $_{0}^{\prime}\equiv c_{0}^{\prime}(c_{0},\tau,d,p), cc$$\equiv c(\alpha,d,p,c_{0},C_{0},\tau)>0$ and $n_{0}\equiv n_{0}(\alpha,d,p,c_{0},C_{0},S,m,\tau)\in\mathbb{N}$, such that
$\widetilde{\pi}_{\alpha,p}\Bigl(\Bigl\{g\in\mathcal{G}_{\alpha,p}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\Bigr\}\Bigr)\leq\exp\bigl(-cm^{-p/\alpha}n^{p/(2\alpha+p)}\bigr)$ (61)
for $n\geq n_{0}$. Moreover, there exists $c^{\prime}>0$, depending only on $\alpha,\beta,d,p,c_{0},C_{0},S,\tau$ and $\sigma$, such that for $n\geq n_{0}$,
$\pi_{\alpha,p}\bigl(B^{\prime}_{\alpha,p}(P_{g^{\circ}},c^{\prime}mn^{-\alpha/(2\alpha+p)})\bigr)\leq\exp\bigl(-cm^{-p/\alpha}n^{p/(2\alpha+p)}\bigr)$ (62)
where $B^{\prime}_{\alpha,p}(P_{g^{\circ}},c^{\prime}mn^{-\alpha/(2\alpha+p)})\coloneqq\bigl\{P\in\mathcal{P}_{\alpha,p}:d_{\mathrm{H}}(P_{g^{\circ}},P)\leq c^{\prime}mn^{-\alpha/(2\alpha+p)}\bigr\}$ is defined as in (38).
Proof.
Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that may depend only on $\alpha,d,p,c_{0},C_{0}$ and $\tau$. Let $c_{0}^{\prime}\equiv c_{0}^{\prime}(c_{0},\tau,d,p)$ be the quantity in Lemma 27. Let $\widetilde{g}_{\alpha,p}$ be defined as in (11) so that $\widetilde{g}_{\alpha,p}\sim\widetilde{\pi}_{\alpha,p}$. Then
$\widetilde{\pi}_{\alpha,p}$ $\Bigl(\Bigl\{g\in\mathcal{G}_{\alpha,p}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\Bigr\}\Bigr)$
$=\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\biggr\}$
$=\mathbb{E}\biggl[\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}\biggr].$ (63)
Now conditional on $U^{(p)}$, we fix $g^{\prime}\in\mathcal{G}_{\alpha}^{(p)}$ such that $\bigl\|g^{\prime}\bigl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\bigr)-h^{\circ}\bigl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\bigr)\bigr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}$. We may assume without loss of generality that such $g^{\prime}$ exists, since otherwise the conditional probability in the expectation in (63) would be zero. Using the same argument as that in (51), we deduce that
$\mathbb{P}$ $\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
$\leq\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
$=\mathbb{P}\bigl(\|\widetilde{g}_{\alpha}^{(p)}-g^{\prime}\|_{L^{2}(Q(U^{(p)}))}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\big|\,U^{(p)}\bigr),$ (64)
where $Q(U^{(p)})$ denotes the conditional distribution of $\frac{(U^{(p)})^{\top}X+\bm{1}_{p}}{2}$ given $U^{(p)}$, with $X\sim P_{X}$. By Lemma 27, for every $U\in V_{p}(\mathbb{R}^{d})$, there exists a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ such that $Q(U)(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$. We may thus apply (49), with $d$ replaced by $p$, $m$ replaced by $4m$ and $c_{0}$ replaced by $c_{0}^{\prime}$ therein, to deduce that
$\mathbb{P}\bigl\{\bigl\|\widetilde{g}_{\alpha}^{(p)}-g^{\prime}\bigr\|_{L^{2}(Q(U^{(p)}))}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\big|\,U^{(p)}\bigr\}\leq\exp\bigl(-cm^{-p/\alpha}n^{p/(2\alpha+p)}\bigr),$
for $n\geq n_{0}$. Combining this with (63) and (64) proves (61).
Finally, the proof of (62) follows a very similar argument to that in (50). ∎
Lemma 22.
Let $(\alpha,p)\in\mathcal{A}$, $C>0$, $P_{0}\in\mathcal{P}_{\alpha,p}$ and $\epsilon_{n,\alpha,p}\coloneqq Cn^{-\alpha/(2\alpha+p)}$. Then there exists $C^{\prime}$ depending only on $\alpha,p,d,\sigma,C$ and $C_{0}$ such that
$\sup_{\epsilon\geq\epsilon_{n,\alpha,p}}\log\mathcal{N}\bigl(\epsilon/3,B_{\alpha,p}^{\prime}(P_{0},2\epsilon),d_{\mathrm{H}}\bigr)\leq C^{\prime}n\epsilon_{n,\alpha,p}^{2}.$
Proof.
Throughout this proof, $C_{1},C_{2},\ldots$ will denote positive quantities depending only on $\alpha,p,d,\sigma,C$ and $C_{0}$. By Lemma 24(d),
$\sup_{\epsilon\geq\epsilon_{n,\alpha,p}}\log\mathcal{N}\bigl(\epsilon/3,\,B^{\prime}_{\alpha,p}(P_{0},2\epsilon),\,d_{\mathrm{H}}\bigr)$ $\leq\log\mathcal{N}\bigl(\epsilon_{n,\alpha,p}/3,\,\mathcal{P}_{\alpha,p},\,d_{\mathrm{H}}\bigr)$
$\leq\log\mathcal{N}\bigl(2\sigma\epsilon_{n,\alpha,p}/3,\,\mathcal{G}_{\alpha,p},\,\|\cdot\|_{\infty}\bigr).$
Recall that $\mathcal{G}_{\alpha}^{(p)}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ and $\mathcal{G}_{\alpha,p}=\mathcal{G}_{\alpha}^{(p)}\circ\bigl\{x\mapsto\frac{U^{\top}x+\bm{1}_{p}}{2}:U\in V_{p}(\mathbb{R}^{d})\bigr\}$. Now let $\mathcal{N}_{1}$ be a $(\sigma\epsilon_{n,\alpha,p}/3)$-cover of $B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ with respect to $\|\cdot\|_{\infty}$ such that $|\mathcal{N}_{1}|\leq\exp(C_{1}n\epsilon_{n,\alpha,p}^{2})$ for some $C_{1}>0$, which exists by, e.g. Giné and Nickl (2021, Eq. (4.184)). Recall from (59) that for all $x,y\in[0,1]^{p}$,
$|g(x)-g(y)|\leq C_{2}\|x-y\|_{2}^{\alpha\wedge 1}\bigl\{\log(1/\|x-y\|_{2})\vee 1\bigr\}.$ (65)
The function $z\mapsto z^{\alpha\wedge 1}\log(1/z)$ is increasing on $(0,e^{-1/(\alpha\wedge 1)}]$, so let $\delta\coloneqq\frac{1}{\sqrt{pd}}\bigl(\frac{\sigma\epsilon_{n,\alpha,p}}{3C_{3}\log(en)}\bigr)^{1/(\alpha\wedge 1)}\wedge\frac{1}{\sqrt{pd}}e^{-1/(\alpha\wedge 1)}$ for $C_{3}>0$ large enough such that
$C_{2}(\delta\sqrt{pd})^{\alpha\wedge 1}\biggl\{\log\biggl(\frac{1}{\delta\sqrt{pd}}\biggr)\vee 1\biggr\}\leq\frac{\sigma\epsilon_{n,\alpha,p}}{3}.$ (66)
Let $\Delta\coloneqq\{i\delta:i\in\mathbb{Z},|i\delta|\leq 1\}$ and $\mathcal{N}_{2}\coloneqq\Delta^{d\times p}\subseteq\mathbb{R}^{d\times p}$. Then for any $g\in\mathcal{G}_{\alpha}^{(p)}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ and $U\in V_{p}(\mathbb{R}^{d})$, there exist $g^{\prime}\in\mathcal{N}_{1}$ and $U^{\prime}\in\mathcal{N}_{2}$ such that $\|g^{\prime}-g\|_{\infty}\leq\sigma\epsilon_{n,\alpha,p}/3$ and $\|U^{\prime}-U\|_{\mathrm{op}}\leq\sqrt{pd}\|U^{\prime}-U\|_{\infty}\leq\delta\sqrt{pd}$. Thus, for all $x\in\mathbb{B}^{d}$, we have $\|U^{\top}x-(U^{\prime})^{\top}x\|_{2}\leq\delta\sqrt{pd}$, so by (65) and (66), we deduce that
$\biggl|g\biggl($ $\frac{U^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|$
$\leq\biggl|g\biggl(\frac{U^{\top}x+\bm{1}_{p}}{2}\biggr)-g\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|+\biggl|g\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|$
$\leq\frac{2\sigma\epsilon_{n,\alpha,p}}{3}.$
Therefore, $\mathcal{N}_{1}\circ\mathcal{N}_{2}$ is a $(2\sigma\epsilon_{n,\alpha,p}/3)$-cover of $\mathcal{G}_{\alpha,p}$ with respect to $\|\cdot\|_{\infty}$. Hence, we have
$\log\mathcal{N}\bigl(2\sigma\epsilon_{n,\alpha,p}/3,\,\mathcal{G}_{\alpha,p},\,\|\cdot\|_{\infty}\bigr)\leq\log|\mathcal{N}_{1}|+\log|\mathcal{N}_{2}|\leq C^{\prime}n\epsilon_{n,\alpha,p}^{2},$ (67)
which proves the claim. ∎
Proof of Proposition 6.
The proof of Proposition 6 follows the same arguments as the proof of Proposition 4, but instead we apply Lemmas 20, 21 and 22 to verify the assumptions of Corollary 17. ∎
F.3 Proofs of Theorems 5 and 7
Proof of Theorem 5.
For any test distribution $\mu$ on $\mathcal{P}_{\beta}$ such that $\chi^{2}(\mu,\pi_{\beta})\leq\kappa$, we have
$\chi^{2}(\mu,\pi)=\mathbb{E}_{\pi}\biggl\{\biggl(\frac{\mathrm{d}\mu}{\mathrm{d}\pi}\biggr)^{2}\biggr\}-1\leq\max_{\alpha\in\mathcal{A}}\frac{1}{\lambda_{\alpha}}\cdot\mathbb{E}_{\pi_{\beta}}\biggl\{\biggl(\frac{\mathrm{d}\mu}{\mathrm{d}\pi_{\beta}}\biggr)^{2}\biggr\}-1\leq(\kappa+1)\max_{\alpha\in\mathcal{A}}\frac{1}{\lambda_{\alpha}}-1\eqqcolon\kappa^{\prime}.$
Now, by Proposition 1, then by taking $d_{\mathrm{model}}^{\circ}\geq d+2$, $d_{\mathrm{ffn}}^{\circ}\in\mathbb{N}$ and $T\in\mathbb{N}$ large enough that Proposition 3 holds with $\epsilon=1/\{16R^{2}n^{2}(\kappa^{\prime}+1)\}$ and applying Proposition 4, we have
$\mathcal{R}_{\mu}^{\mathrm{ICL}}(\widetilde{f}_{T})$ $\leq 4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]$
$\leq\frac{1}{n}+C^{\prime}n^{-2\beta/(2\beta+d)}\leq Cn^{-2\beta/(2\beta+d)},$
where $C^{\prime},C>0$ depend on neither $n$ nor $\kappa$. ∎
Proof of Theorem 7.
As in the proof of Theorem 5, for any test distribution $\mu$ on $\mathcal{P}_{\beta,r}$ such that $\chi^{2}(\mu,\pi_{\beta,r})\leq\kappa$, we have $\chi^{2}(\mu,\pi)\leq(\kappa+1)\max_{(\alpha,p)\in\mathcal{A}^{\prime}}\lambda_{\alpha,p}^{-1}-1\eqqcolon\kappa^{\prime}$. The result therefore follows by the same argument as in the proof of Theorem 5, except that we apply Proposition 6 in place of Proposition 4 to obtain the improved bound $\mathcal{R}_{\mu}^{\mathrm{ICL}}(\widetilde{f}_{T})\leq Cn^{-2\beta/(2\beta+r)}$ for sufficiently large $T$, where $C>0$ depends on neither $n$ nor $\kappa$. ∎
F.4 Proof of Theorem 8
Proof of Theorem 8.
We use the Bayes risk lower bound techniques developed by Chen, Guntuboyina and Zhang (2016). Let $(\mathcal{P}_{*},\mathcal{G}_{*})$ be either $(\mathcal{P}_{\beta},\mathcal{G}_{\beta})$ defined as in Section 4.1 or $(\mathcal{P}_{\beta,r},\mathcal{G}_{\beta,r})$ defined as in Section 4.2, and let $\widetilde{\mu}$ denote the distribution on $\mathcal{G}_{*}$ induced by $\mu$ on $\mathcal{P}_{*}$, so that $\widetilde{\mu}$ is the distribution of the random function $x\mapsto\mathbb{E}_{P}(Y\,|\,X=x)$, where the randomness comes from $P\sim\mu$. By Chen, Guntuboyina and Zhang (2016, Corollary 12(i) and Eq. (48)), we have
$\inf_{\widehat{g}_{n}\in\widehat{\mathcal{G}}_{n}}\mathcal{R}_{\mu}^{\mathrm{ICL}}(\widehat{g}_{n})$ $=\inf_{\widehat{g}_{n}\in\widehat{\mathcal{G}}_{n}}\mathbb{E}_{P\sim\mu}\mathbb{E}_{\mathcal{D}_{n}|P\sim P^{\otimes n}}\bigl\{\|\widehat{g}_{n}(\cdot,\mathcal{D}_{n})-g_{P}\|_{L^{2}(P_{X})}^{2}\bigr\}$
$=\inf_{\widehat{g}_{n}\in\widehat{\mathcal{G}}_{n}}\mathbb{E}_{\widetilde{g}\sim\widetilde{\mu}}\mathbb{E}_{\mathcal{D}_{n}|\widetilde{g}\sim P_{\widetilde{g}}^{\otimes n}}\bigl\{\|\widehat{g}_{n}(\cdot,\mathcal{D}_{n})-\widetilde{g}\|_{L^{2}(P_{X})}^{2}\bigr\}$
$\geq\frac{1}{2}\sup\biggl\{t>0:\sup_{g^{\circ}\in\mathcal{G}_{*}}\widetilde{\mu}\Bigl(\bigl\{g\in\mathcal{G}_{*}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}<t\bigr\}\Bigr)<\frac{1}{4}e^{-2I^{\mathrm{up}}_{\mathrm{KL}}}\biggr\},$ (68)
where
$I^{\mathrm{up}}_{\mathrm{KL}}\coloneqq\inf_{\eta>0}\bigl\{\log\mathcal{N}(\eta^{2}/n,\mathcal{P}_{*},\mathrm{KL})+\eta^{2}\bigr\}$
and we recall that $\mathcal{N}(\eta^{2}/n,\mathcal{P}_{*},\mathrm{KL})$ is the $(\eta^{2}/n)$-covering number of $\mathcal{P}_{*}$ with respect to the Kullback–Leibler divergence. By Lemma 24(a), we deduce that
$I^{\mathrm{up}}_{\mathrm{KL}}\leq\inf_{\eta>0}\bigl\{\log\mathcal{N}\bigl(\sigma\eta\sqrt{2/n},\mathcal{G}_{*},\|\cdot\|_{\infty}\bigr)+\eta^{2}\bigr\}.$
From now on, $C_{1},C_{2},\ldots$ will denote positive quantities that do not depend on $n$.
(a) By choosing $\eta\coloneqq n^{d/(2\beta+d)}$ and Giné and Nickl (2021, Eq. (4.184)), we deduce that
$I^{\mathrm{up}}_{\mathrm{KL}}\leq C_{1}n^{d/(2\beta+d)}.$
By the measurable bijection between $\mathcal{G}_{\beta}$ and $\mathcal{P}_{\beta}$ and our assumption on $\mu$, we have that $\sup_{g\in\mathcal{G}_{\beta}}\frac{\mathrm{d}\widetilde{\mu}}{\mathrm{d}\widetilde{\pi}_{\beta}}(g)<\infty$. Thus, by Lemma 19, there exists $C_{2}>0$ small enough that for all $g^{\circ}\in\mathcal{G}_{\beta}$,
$\widetilde{\mu}\Bigl(\bigl\{g\in\mathcal{G}_{\beta}$ $:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}<C_{2}n^{-2\beta/(2\beta+d)}\bigr\}\Bigr)$
$\leq\biggl\{\sup_{g\in\mathcal{G}_{\beta}}\frac{\mathrm{d}\widetilde{\mu}}{\mathrm{d}\widetilde{\pi}_{\beta}}(g)\biggr\}\widetilde{\pi}_{\beta}\Bigl(\bigl\{g\in\mathcal{G}_{\beta}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}<C_{2}n^{-2\beta/(2\beta+d)}\bigr\}\Bigr)<\frac{1}{4}e^{-2I^{\mathrm{up}}_{\mathrm{KL}}}.$
The claim then follows from (68).
(b) By choosing $\eta\coloneqq n^{d/(2\beta+r)}$ and applying (67) with $(\beta,r)$ replacing $(\alpha,p)$, we deduce that
$I^{\mathrm{up}}_{\mathrm{KL}}\leq C_{3}n^{r/(2\beta+r)}.$
Similarly, by Lemma 21, there exists $C_{4}>0$ small enough that for all $g^{\circ}\in\mathcal{G}_{\beta,r}$,
$\widetilde{\mu}\Bigl(\bigl\{g$ $\in\mathcal{G}_{\beta,r}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}<C_{4}n^{-2\beta/(2\beta+r)}\bigr\}\Bigr)$
$\leq\biggl\{\sup_{g\in\mathcal{G}_{\beta,r}}\frac{\mathrm{d}\widetilde{\mu}}{\mathrm{d}\widetilde{\pi}_{\beta,r}}(g)\biggr\}\widetilde{\pi}_{\beta,r}\Bigl(\bigl\{g\in\mathcal{G}_{\beta,r}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}<C_{4}n^{-2\beta/(2\beta+r)}\bigr\}\Bigr)<\frac{1}{4}e^{-2I^{\mathrm{up}}_{\mathrm{KL}}}.$
The claim then follows again from (68). ∎
Appendix G Auxiliary lemmas
The lemma below is a modification of Ghosal, Lember and van der Vaart (2008, Lemma 6.3), where we assume that the log-likelihood ratio is sub-Gamma in the left tail with uniform sub-Gamma parameters in order to obtain an exponential concentration.
Lemma 23.
Let $\mathcal{P}$ be a measurable space of distributions on a measurable space $\mathcal{Z}$, let $P_{0}$ be a distribution on $\mathcal{Z}$ and let $Z_{1},\ldots,Z_{n}\overset{\mathrm{iid}}{\sim}P_{0}$. For $\epsilon>0$, define
$B(P_{0},\epsilon)\coloneqq\Bigl\{P\in\mathcal{P}:\mathrm{KL}(P_{0},P)\leq\epsilon^{2},\,\mathrm{V}_{2}(P_{0},P)\leq\epsilon^{2}\Bigr\}.$
For $i\in[n]$ and $P\in\mathcal{P}$ with $P\ll P_{0}$, let
$W(P,Z_{i})\coloneqq\log\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i})-\mathbb{E}\log\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i}).$
Assume that there exist $a,c>0$ such that for all $P\in B(P_{0},\epsilon)$, we have that $W(P,Z_{1})$ is sub-Gamma in the left tail with variance parameter $a\epsilon^{2}$ and scale parameter $c$. Then, for any probability measure $\Pi$ on $\mathcal{P}$ and any $\epsilon,D>0$, we have with $P_{0}^{\otimes n}$-probability at most $\exp\bigl(-\frac{D^{2}n\epsilon^{2}}{2(a+cD)}\bigr)$ that
$\int_{B(P_{0},\epsilon)}\prod_{i=1}^{n}\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i})\,\mathrm{d}\Pi(P)<\Pi\bigl(B(P_{0},\epsilon)\bigr)e^{-(1+D)n\epsilon^{2}}.$ (69)
Proof.
Writing $B\coloneqq B(P_{0},\epsilon)$, we observe that the conclusion is clear if $\Pi(B)=0$ because the right-hand side of (69) is zero. Moreover, by replacing $\Pi$ with its conditional distribution on $B$ if necessary, we may assume without loss of generality that $\Pi(B)=1$. Now, by Jensen’s inequality,
$\frac{1}{n}\log\int_{B}\prod_{i=1}^{n}\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i})\,\mathrm{d}\Pi(P)$ $\geq\frac{1}{n}\sum_{i=1}^{n}\int_{B}\log\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i})\,\mathrm{d}\Pi(P)$
$=\frac{1}{n}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)-\int_{B}\mathrm{KL}(P_{0},P)\,\mathrm{d}\Pi(P)$
$\geq\frac{1}{n}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)-\epsilon^{2},$ (70)
where the final inequality follows from the definition of $B$. Moreover, for $\lambda\in(-1/c,0]$, by Jensen’s inequality and Fubini’s theorem,
$\mathbb{E}_{Z_{1}}\exp\biggl\{\lambda\int_{B}W(P,Z_{1})\,\mathrm{d}\Pi(P)\biggr\}$ $\leq\mathbb{E}_{Z_{1}}\biggl[\int_{B}\exp\bigl\{\lambda W(P,Z_{1})\bigr\}\,\mathrm{d}\Pi(P)\biggr]$
$=\int_{B}\mathbb{E}_{Z_{1}}\exp\bigl\{\lambda W(P,Z_{1})\bigr\}\,\mathrm{d}\Pi(P)\leq\exp\biggl(\frac{a\epsilon^{2}\lambda^{2}}{2(1-c\lambda)}\biggr),$
where the final inequality follows from the sub-Gamma assumption on $W(P,Z_{1})$. This shows that $\int_{B}W(P,Z_{1})\,\mathrm{d}\Pi(P)$ is sub-Gamma in the left tail with variance parameter $a\epsilon^{2}$ and scale parameter $c$, so $n^{-1}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)$ is also sub-Gamma in the left tail with variance parameter $a\epsilon^{2}/n$ and scale parameter $c/n$ (e.g. Samworth and Shah, 2025+, Exercise 10.6.12). Hence, by (70),
$\mathbb{P}\biggl(\frac{1}{n}\log\int_{B}\prod_{i=1}^{n}$ $\frac{\mathrm{d}P}{\mathrm{d}P_{0}}(Z_{i})\,\mathrm{d}\Pi(P)<-(1+D)\epsilon^{2}\biggr)$
$\leq\mathbb{P}\biggl(\frac{1}{n}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)<-D\epsilon^{2}\biggr)\leq\exp\biggl(-\frac{D^{2}n\epsilon^{2}}{2(a+cD)}\biggr),$
where the final inequality follows from Samworth and Shah (2025+, Proposition 10.6.8(a)). ∎
We next derive some properties about the distributions of nonparametric regression models with Gaussian noise.
Lemma 24.
Let $P_{X}$ be a Borel distribution on $\mathcal{X}\subseteq\mathbb{R}^{d}$, let $X\sim P_{X}$ and let $\xi\sim N(0,\sigma^{2})$ be independent of $X$. Further, let $R>0$, $g_{1},g_{2}:\mathcal{X}\to[-R,R]$ be Borel measurable, let $Y_{1}=g_{1}(X)+\xi$ and $Y_{2}=g_{2}(X)+\xi$. Finally, let $P_{1}$ and $P_{2}$ denote the distributions of $(X,Y_{1})$ and $(X,Y_{2})$ respectively. Then
  • (a)
    $\mathrm{KL}(P_{1},P_{2})=\frac{1}{2\sigma^{2}}\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}$.
  • (b)
    $\mathrm{V}_{2}(P_{1},P_{2})=\frac{1}{2\sigma^{2}}\mathrm{Var}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}+2\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}$.
  • (c)
    If $\mathrm{KL}(P_{1},P_{2})\leq\epsilon^{2}$ and $\mathrm{V}_{2}(P_{1},P_{2})\leq\epsilon^{2}$, then $W\coloneqq\log\frac{\mathrm{d}P_{2}}{\mathrm{d}P_{1}}(X,Y_{1})-\mathbb{E}\log\frac{\mathrm{d}P_{2}}{\mathrm{d}P_{1}}(X,Y_{1})$ is sub-Gamma in both tails with variance parameter $\bigl(8\vee\frac{2}{\sigma^{2}}\bigr)\epsilon^{2}$ and scale parameter $c\coloneqq\frac{2R}{\sigma}\vee\frac{4R^{2}}{3\sigma^{2}}$.
  • (d)
    $\frac{1-e^{-R^{2}/(2\sigma^{2})}}{2R^{2}}\cdot\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}\leq d_{\mathrm{H}}^{2}(P_{1},P_{2})\leq\frac{1}{4\sigma^{2}}\cdot\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}$.
Proof.
(a) Let $P_{Y_{1}|X}$ and $P_{Y_{2}|X}$ denote the conditional distributions of $Y_{1}$ given $X$ and $Y_{2}$ given $X$ respectively. Then, using the fact that $Y_{1}=g_{1}(X)+\xi$,
$\log\frac{\mathrm{d}P_{2}}{\mathrm{d}P_{1}}(X,Y_{1})=\log\frac{\mathrm{d}P_{Y_{2}|X}}{\mathrm{d}P_{Y_{1}|X}}(X,Y_{1})$ $=-\frac{\bigl(Y_{1}-g_{2}(X)\bigr)^{2}}{2\sigma^{2}}+\frac{\bigl(Y_{1}-g_{1}(X)\bigr)^{2}}{2\sigma^{2}}$
$=\frac{-\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}-2\bigl(g_{1}(X)-g_{2}(X)\bigr)\xi}{2\sigma^{2}}.$ (71)
Hence,
$\mathrm{KL}(P_{1},P_{2})=-\mathbb{E}\log\frac{\mathrm{d}P_{2}}{\mathrm{d}P_{1}}(X,Y_{1})=\frac{1}{2\sigma^{2}}\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}.$ (72)
(b) By (71) and (72), we have
$\mathrm{V}_{2}(P_{1},P_{2})$ $=\frac{1}{2\sigma^{2}}\mathbb{E}\Bigl[\Bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}+2\bigl(g_{1}(X)-g_{2}(X)\bigr)\xi-\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}\Bigr\}^{2}\Bigr]$
$=\frac{1}{2\sigma^{2}}\mathrm{Var}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}+2\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\},$
where in the final equality, we used the fact that $\xi\perp\!\!\!\perp X$, $\mathbb{E}(\xi)=0$ and $\mathbb{E}(\xi^{2})=\sigma^{2}$.
(c) From (71) and (72), we deduce that
$W$ $=\log\frac{\mathrm{d}P_{2}}{\mathrm{d}P_{1}}(X,Y_{1})-\mathbb{E}\log\frac{\mathrm{d}P_{2}}{\mathrm{d}P_{1}}(X,Y_{1})$
$=\frac{\bigl(g_{2}(X)-g_{1}(X)\bigr)\xi}{\sigma^{2}}-\frac{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}-\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\}}{2\sigma^{2}}\eqqcolon W_{1}+W_{2}.$
For $\lambda\in[0,\sigma/R)$, using the fact that $\log(1+x)\leq x$ for $x\geq 0$ and Fubini’s theorem,
$\log\mathbb{E}e^{\lambda W_{1}}$ $=\log\mathbb{E}\exp\biggl(\frac{\lambda\bigl(g_{2}(X)-g_{1}(X)\bigr)\xi}{\sigma^{2}}\biggr)=\log\mathbb{E}\exp\biggl(\frac{\lambda^{2}\bigl(g_{2}(X)-g_{1}(X)\bigr)^{2}}{2\sigma^{2}}\biggr)$
$\leq\sum_{r=1}^{\infty}\frac{\lambda^{2r}\mathbb{E}\bigl\{\bigl(g_{2}(X)-g_{1}(X)\bigr)^{2r}\bigr\}}{(2\sigma^{2})^{r}\cdot r!}\leq\mathbb{E}\bigl\{\bigl(g_{2}(X)-g_{1}(X)\bigr)^{2}\bigr\}\sum_{r=1}^{\infty}\frac{\lambda^{2r}(2R)^{2r-2}}{(2\sigma^{2})^{r}\cdot r!}$
$\leq\lambda^{2}\epsilon^{2}\sum_{r=1}^{\infty}\frac{2^{r-1}}{r!}\biggl(\frac{\lambda R}{\sigma}\biggr)^{2r-2}\leq\lambda^{2}\epsilon^{2}\sum_{r=1}^{\infty}\biggl(\frac{\lambda R}{\sigma}\biggr)^{r-1}=\frac{2\lambda^{2}\epsilon^{2}}{2\bigl(1-\frac{\lambda R}{\sigma}\bigr)}.$
The same holds true if $W_{1}$ is replaced by $W_{1}$. Thus $W_{1}$ is sub-Gamma in both tails with variance parameter $2\epsilon^{2}$ and scale parameter $R/\sigma$. By part (b) and the assumption that $\mathrm{V}_{2}(P_{1},P_{2})\leq\epsilon^{2}$, we have $\mathbb{E}(W_{2}^{2})\leq\epsilon^{2}/(2\sigma^{2})$. Moreover, since $W_{2}$ is bounded by $2R^{2}/\sigma^{2}$, we have for $r\geq 3$ that
$\mathbb{E}(|W_{2}|^{r})\leq\mathbb{E}(W_{2}^{2})\cdot\Bigl(\frac{2R^{2}}{\sigma^{2}}\Bigr)^{r-2}\leq\frac{r!\epsilon^{2}}{2\cdot 2\sigma^{2}}\Bigl(\frac{2R^{2}}{3\sigma^{2}}\Bigr)^{r-2}.$
The same bound holds for $W_{2}$, so by Samworth and Shah (2025+, Proposition 10.6.8(b)), $W_{2}$ is sub-Gamma in both tails with variance parameter $\frac{\epsilon^{2}}{2\sigma^{2}}$ and scale parameter $\frac{2R^{2}}{3\sigma^{2}}$. Finally, by Samworth and Shah (2025+, Exercise 10.6.12), we have that $W=W_{1}+W_{2}$ is sub-Gamma in both tails with variance parameter $\bigl(8\vee\frac{2}{\sigma^{2}}\bigr)\epsilon^{2}$ and scale parameter $c=\frac{2R}{\sigma}\vee\frac{4R^{2}}{3\sigma^{2}}$.
(d) We have
$d_{\mathrm{H}}^{2}(P_{1},P_{2})=\mathbb{E}\,d_{\mathrm{H}}^{2}(P_{Y_{1}|X},P_{Y_{2}|X})$ $=2\,\mathbb{E}\biggl\{1-\exp\biggl(-\frac{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}}{8\sigma^{2}}\biggr)\biggr\}$
$\geq\frac{1-e^{-R^{2}/(2\sigma^{2})}}{2R^{2}}\cdot\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\},$
where the inequality follows since $x\mapsto(1-e^{-x})/x$ is decreasing on $(0,\infty)$ and moreover $\frac{(g_{1}(X)-g_{2}(X))^{2}}{8\sigma^{2}}\leq\frac{R^{2}}{2\sigma^{2}}$. On the other hand, using the fact that $1-e^{-x}\leq x$ for $x\geq 0$, we also have
$2\mathbb{E}\biggl\{1-\exp\biggl(-\frac{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}}{8\sigma^{2}}\biggr)\biggr\}\leq\frac{1}{4\sigma^{2}}\cdot\mathbb{E}\bigl\{\bigl(g_{1}(X)-g_{2}(X)\bigr)^{2}\bigr\},$
as required. ∎
Lemma 25.
Suppose $f\in L^{2}([0,1]^{d})$ has wavelet decomposition
$f=\sum_{k\in K}a_{k}\Phi_{k}+\sum_{\ell=\ell_{0}}^{\infty}\sum_{\gamma\in\Gamma_{\ell}}b_{\ell,\gamma}\Psi_{\ell,\gamma}.$
Let $0\leq c_{1}<c_{2}\leq 1$ and let $\Lambda\coloneqq\bigl\{(\ell,\gamma):\mathrm{supp}(\Psi_{\ell,\gamma})\subseteq[c_{1},c_{2}]^{d}\bigr\}$. Then
$\|f\|_{L^{2}([c_{1},c_{2}]^{d})}^{2}\geq\sum_{(\ell,\gamma)\in\Lambda}b^{2}_{\ell,\gamma}.$
Proof.
Note that $\mathrm{supp}(\Psi_{\ell,\gamma})\subseteq[c_{1},c_{2}]^{d}$ for all $(\ell,\gamma)\in\Lambda$, so by the orthonormality of the wavelet basis, we have that $(\Psi_{\ell,\gamma})_{(\ell,\gamma)\in\Lambda}$ is also a sequence of orthonormal functions in the Hilbert space $L^{2}([c_{1},c_{2}]^{d})$ equipped with the inner product
$\langle f_{1},f_{2}\rangle_{[c_{1},c_{2}]^{d}}\coloneqq\int_{[c_{1},c_{2}]^{d}}f_{1}(x)f_{2}(x)\,\mathrm{d}x.$
Moreover, for $(\ell,\gamma)\in\Lambda$, we have
$\langle f,\Psi_{\ell,\gamma}\rangle_{[c_{1},c_{2}]^{d}}=\int_{[c_{1},c_{2}]^{d}}f(x)\Psi_{\ell,\gamma}(x)\,\mathrm{d}x=\int_{[0,1]^{d}}f(x)\Psi_{\ell,\gamma}(x)\,\mathrm{d}x=b_{\ell,\gamma}.$
Thus, an application of Bessel’s inequality (e.g. Rudin, 1987, Theorem 4.17) yields the desired result. ∎
Lemma 26.
Let $U^{(r)}$ be uniformly distributed on the Stiefel manifold $V_{r}(\mathbb{R}^{d})$ and let $U\in V_{r}(\mathbb{R}^{d})$ be deterministic. Then for any $\epsilon>0$, there exists $c>0$, depending only on $d$ and $r$, such that
$\mathbb{P}\bigl(\|U^{(r)}-U\|_{\mathrm{op}}\leq\epsilon\bigr)\geq c\biggl(\epsilon\wedge\frac{1}{2}\biggr)^{dr}.$
Proof.
We may assume that $\epsilon\leq 1/2$. By Chikuse (2003, Theorem 2.2.1(iii)), we may assume without loss of generality that $U^{(r)}=Z(Z^{\top}Z)^{-1/2}\eqqcolon F(Z)$, where $Z$ is a $d\times r$ random matrix with independent $N(0,1)$ entries. Let $\Delta\in\mathbb{R}^{d\times r}$ be such that $\delta\coloneqq\|\Delta\|_{\mathrm{op}}\leq 1/6$, and define
$A\coloneqq U^{\top}\Delta+\Delta^{\top}U+\Delta^{\top}\Delta\in\mathbb{R}^{r\times r}.$
Then $\|A\|_{\mathrm{op}}\leq 3\delta\leq 1/2$, so $(U+\Delta)^{\top}(U+\Delta)=I_{r}+A$ is positive definite. Thus
$\|F(U+\Delta)-U\|_{\mathrm{op}}$ $=\|(U+\Delta)(I_{r}+A)^{-1/2}-U\|_{\mathrm{op}}$
$\leq\|U\|_{\mathrm{op}}\|(I_{r}+A)^{-1/2}-I_{r}\|_{\mathrm{op}}+\|\Delta\|_{\mathrm{op}}\|(I_{r}+A)^{-1/2}\|_{\mathrm{op}}.$
Moreover, we can write $I_{r}+A=VDV^{\top}$ where $V\in\mathbb{R}^{r\times r}$ is orthogonal and $D\in\mathbb{R}^{r\times r}$ is diagonal with diagonal elements bounded between $1-3\delta$ and $1+3\delta$. Thus, $\|(I_{r}+A)^{-1/2}-I_{r}\|_{\mathrm{op}}\leq 3\delta$ and $\|(I_{r}+A)^{-1/2}\|_{\mathrm{op}}\leq 2$, so
$\|F(U+\Delta)-U\|_{\mathrm{op}}\leq 3\delta+2\delta=5\delta.$
Hence, if $\|Z-U\|_{\mathrm{op}}\leq\epsilon/5<1/6$, then $\|F(Z)-U\|_{\mathrm{op}}\leq\epsilon$, so
$\mathbb{P}\bigl(\|U^{(r)}$ $-U\|_{\mathrm{op}}\leq\epsilon\bigr)=\mathbb{P}\bigl(\|F(Z)-U\|_{\mathrm{op}}\leq\epsilon\bigr)\geq\mathbb{P}\biggl(\|Z-U\|_{\mathrm{op}}\leq\frac{\epsilon}{5}\biggr)$
$\geq\mathbb{P}\biggl(\|Z-U\|_{\infty}\leq\frac{\epsilon}{5\sqrt{dr}}\biggr)\geq\biggl\{\mathbb{P}\biggl(|Z_{11}-1|\leq\frac{\epsilon}{5\sqrt{dr}}\biggr)\biggr\}^{dr}\geq\biggl(\frac{2\epsilon}{25\sqrt{dr}}\biggr)^{dr},$
where the third inequality follows since $\|U\|_{\infty}\leq 1$ and $Z$ has independent $N(0,1)$ entries, and the final inequality follows since $\frac{\epsilon}{5\sqrt{dr}}\leq\frac{1}{10}$ and the standard normal density is bounded below by $1/5$ on the interval $[0.9,1.1]$. ∎
Lemma 27.
Let $P_{X}$ be a distribution on $\mathbb{B}^{d}$ that satisfies the assumptions of Lemma 21. Let $p\in[d]$, $U\in V_{p}(\mathbb{R}^{d})$, $X\sim P_{X}$ and let $Q$ denote the distribution of $\frac{U^{\top}X+\bm{1}_{p}}{2}$. Then there exist $c_{0}^{\prime}\equiv c_{0}^{\prime}(c_{0},\tau,d,p)$ and a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ with the property that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$.
Proof.
If $p=d$, then $U\in\mathbb{R}^{d\times d}$ is orthogonal, so for all measurable $A_{0}\subseteq A$,
$Q\biggl(\frac{U^{\top}A_{0}+\bm{1}_{d}}{2}\biggr)=\mathbb{P}(X\in A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0}).$
Moreover, $(U^{\top}A+\bm{1}_{d})/2$ is a Euclidean ball of radius $\tau/2$, so it contains a hypercube of side length $\frac{\tau}{2\sqrt{p}}$. Now suppose that $p\in[d-1]$. For $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$ and $1\leq j_{1}\leq j_{2}\leq d$, we write $v_{j_{1}:j_{2}}\coloneqq(v_{j_{1}},\ldots,v_{j_{2}})^{\top}\in\mathbb{R}^{j_{1}-j_{2}+1}$. Let $o\in\mathbb{R}^{d}$ be the centre of $A$ and let $K\subseteq A$ be the Euclidean ball of radius $\tau/2$ and centred at $o$. We may write $U=VP$ where $V\in\mathbb{R}^{d\times d}$ is orthogonal and $P\in\mathbb{R}^{d\times p}$ is the first $p$ columns of $I_{d}$. Let $S\subseteq K$ be Borel measurable, and define $T\coloneqq\{x\in V^{\top}A:x_{1:p}\in P^{\top}V^{\top}S\}$. Then $VT\subseteq A$; moreover, if $y=(V^{\top}x)_{1:p}$ for some $x\in S\subseteq K$ and $z\in\mathbb{R}^{d-p}$ satisfies $\|z-(V^{\top}o)_{(p+1):d}\|_{2}\leq\tau/2$, then by the choice of $K$,
$\biggl\|\begin{pmatrix}y\\ z\end{pmatrix}-V^{\top}o\biggr\|_{2}^{2}=\|y-(V^{\top}o)_{1:p}\|_{2}^{2}+\|z-(V^{\top}o)_{(p+1):d}\|_{2}^{2}\leq(\tau/2)^{2}+(\tau/2)^{2}<\tau^{2},$
so $(y^{\top},z^{\top})^{\top}\in T$. Thus
$T^{\prime}\coloneqq\bigl\{(y^{\top},z^{\top})^{\top}\in\mathbb{R}^{p+(d-p)}:y\in P^{\top}V^{\top}S,\,\|z-(V^{\top}o)_{(p+1):d}\|_{2}\leq\tau/2\bigr\}\subseteq T.$ (73)
Therefore, writing $v(\tau,d,p)$ for the Lebesgue measure of a Euclidean ball in $\mathbb{R}^{d-p}$ of radius $\tau/2$, we deduce that
$Q\biggl(\frac{U^{\top}S+\bm{1}_{p}}{2}\biggr)$ $=\mathbb{P}\biggl(\frac{P^{\top}V^{\top}X+\bm{1}_{p}}{2}\in\frac{P^{\top}V^{\top}S+\bm{1}_{p}}{2}\biggr)$
$\geq\mathbb{P}(V^{\top}X\in T)=\mathbb{P}(X\in VT)\geq c_{0}\mathrm{Vol}_{d}(VT)=c_{0}\mathrm{Vol}_{d}(T)\geq c_{0}\mathrm{Vol}_{d}(T^{\prime})$
$=c_{0}\int_{y\in P^{\top}V^{\top}S}\int_{z\in\mathbb{R}^{d-p}:\|z-(V^{\top}o)_{(p+1):d}\|_{2}\leq\tau/2}\,\mathrm{d}z\,\mathrm{d}y$
$=c_{0}v(\tau,d,p)\mathrm{Vol}_{p}(P^{\top}V^{\top}S)\eqqcolon c_{0}^{\prime}\mathrm{Vol}_{p}\biggl(\frac{U^{\top}S+\bm{1}_{p}}{2}\biggr).$
Since this holds for all measurable $S\subseteq K$, we deduce that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all $A_{0}^{\prime}\subseteq\frac{U^{\top}K+\bm{1}_{p}}{2}$. Now $\frac{U^{\top}K+\bm{1}_{p}}{2}$ is a ball of radius $\tau/4$ in $\mathbb{R}^{p}$, so it contains a hypercube $A^{\prime}$ of side length $\frac{\tau}{2\sqrt{p}}$. ∎

### v2

@@
 Our goal is to understand the performance of the pretrained Transformer when evaluated on a different test distribution $\mu$, consisting of tasks of fixed difficulty $\beta\in\mathcal{A}$, and with potential distribution shift relative to $\pi_{\beta}$, subject to the chi-squared divergence $\chi^{2}(\mu,\pi_{\beta})$ being at most $\kappa$.
-In particular, we consider nonparametric regression problems with random smoothness, and multi-index models with random smoothness as well as random effective dimension.
+In particular, we consider nonparametric regression problems with random smoothness, and multi-index models with both random smoothness and random effective dimension.
 We prove that a large Transformer pretrained on sufficient data achieves the optimal rate of convergence corresponding to the difficulty level $\beta$, uniformly over test distributions $\mu$ in the chi-squared divergence ball.
@@
 A striking ability of pretrained Transformers, first observed by Brown et al.
-(2020) in Large Language Models (LLMs), is the phenomenon of in-context learning (ICL): given a prompt containing examples and a query, Transformers can learn the underlying pattern from the examples and produce accurate output for the query, without updating its parameters.
-For instance, given the values of the function $(x,y)\mapsto xy^{2}$ or $\text{subfield}\mapsto\text{field}$’ at a few distinct points, LLMs are able to recover the value of the function at a new point:
-$\overbrace{\underbrace{f(1,1)=1,\;f(1,2)=4,\;f(2,2)=8,\;f(2,3)=18,}_{\text{examples}}\;\underbrace{f(3,3)=}_{\text{query}}}^{\text{prompt}}\underbrace{27\vphantom{f(2,2)}}_{\text{output}},$ (1)
-$\overbrace{\underbrace{\text{Genetics}\to\text{Biology},\;\text{Relativity}\to\text{Physics},}_{\text{examples}}\;\underbrace{\text{Topology}\to}_{\text{query}}}^{\text{prompt}}\underbrace{\text{Mathematics}\vphantom{y}}_{\text{output}}.$
+(2020) in Large Language Models (LLMs), is the phenomenon of in-context learning (ICL): given a prompt (context) containing examples and a query, Transformers can learn the underlying pattern from the examples and produce accurate output for the query, without updating its parameters.
 The ICL ability of Transformers was also exploited by Hollmann et al.
-(2025) to build tabular foundation models, which they show to outperform many classical machine learning methods, such as tree-based and boosting algorithms, on regression and classification problems for tabular data.
-Extensive studies on ICL have been conducted from various perspectives (e.g.
-Garg et al., 2022; Xie et al., 2022; Von Oswald et al., 2023; Bai et al., 2023; Kim, Nakamaki and Suzuki, 2024); we refer the readers to Section 1.1 for a more detailed literature review.
-However, most theoretical analyses of ICL have focused on the setting where the distribution of pretraining data (which we call the pretraining prior) and the distribution of the test data (which we call the test distribution) are the same.
-In practice, Transformers are pretrained on a variety of tasks of different difficulties, and the test distribution will typically exhibit distribution shift from the pretraining prior.
-For example, we may train a Transformer on translation tasks between different languages, where Dutch–English translation is typically easier than Chinese–English, since Dutch and English are both West Germanic languages with shared structure.
-At test time, the Transformer is required to translate between two specific languages, while some linguistic patterns may appear more frequently in the test data than in the pretraining data.
-To formulate a theoretical framework, we focus on regression problems (in line with most prior work) and consider a Transformer pretrained on a mixture distribution containing tasks of different difficulties, but where the test distribution consists of tasks of a specific difficulty with potential distribution shift from the corresponding mixture component of the pretraining prior.
-We study the following natural but important questions:
+(2025) to build tabular foundation models, which they show to outperform many machine learning methods, such as tree-based and boosting algorithms, on regression and classification problems for tabular data.
+GeneticsBiologyRelativityPhysicsTopology?MathematicsEasy taskCat10Fish25Deer17Dog13Bird?9Difficult taskPrompt(Context)Output
+Figure 1: Two ICL tasks.
+Blue indicates the (in-context) examples, red indicates the queries, and green indicates the outputs.
+The left column is an easy task, while the right column is a harder task which requires more examples to produce accurate output.
+[…]
 Do pretrained Transformers adapt to the difficulty of the test distribution, and are they robust to distribution shift?
-Our contributions
+1.1 Contributions
   • In Proposition 1, we provide a key test-time risk upper bound, which decomposes into a term that captures both the degree of distribution shift and the proximity of the pretrained Transformer to the posterior regression function, and another term that represents the rate of convergence of the posterior regression function to the true regression function.
@@
 In combination with the results mentioned above, this establishes the adaptivity and distributional robustness of pretrained Transformers.
-  • Finally, in Section 4.3, we show that the expected risk of any estimator over the test distribution $\mu$ cannot achieve a faster rate of convergence than our pretrained Transformers (even if the estimator is allowed to depend on $\mu$).
+  • In Section 4.3, we show that the expected risk of any estimator over the test distribution $\mu$ cannot achieve a faster rate of convergence than our pretrained Transformers (even if the estimator is allowed to depend on $\mu$).
 This provides a more appropriate form of optimality guarantee for pretrained Transformers than that yielded by minimax lower bounds usually considered in ICL literature.
+  • Finally, in Section 5, we verify our theoretical results with empirical experiments.
+These confirm our main conclusions that pretrained Transformers are able to adapt to different task difficulties and are robust to distribution shifts at test time.
 1.2 Related Work
-In-context learning
-There is a large body of theoretical research that aims to explain the ICL ability of Transformers.
+In-context learning   There is a large body of theoretical research that aims to explain the ICL ability of Transformers.
 One line of work shows that Transformers can perform ICL by approximating certain algorithms, such as gradient descent (Ahn et al., 2023; Akyürek et al., 2023; Bai et al., 2023; von Oswald et al., 2023), functional gradient descent (Cheng, Chen and Sra, 2024), basis function regression (Kim, Nakamaki and Suzuki, 2024), reinforcement learning (Lin, Bai and Mei, 2024) and domain adaptation (Hataya, Matsui and Imaizumi, 2024).
@@
 Their results are complementary to ours, and neither subsumes the other.
-Posterior contraction theory
-Posterior contraction theory is an active area of research in statistics that studies the concentration of the posterior distribution around a true data generating distribution (i.e.
+Posterior contraction theory   Posterior contraction theory is an active area of research in statistics that studies the concentration of the posterior distribution around a true data generating distribution (i.e.
 from a frequentist perspective).
@@
 Given $a,b\in\mathbb{R}$, we write $a\wedge b\coloneqq\min\{a,b\}$ and $a\vee b\coloneqq\max\{a,b\}$.
-For $d\in\mathbb{N}$, we define $\bm{0}_{d}\coloneqq(0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ and $\bm{1}_{d}\coloneqq(1,\ldots,1)^{\top}\in\mathbb{R}^{d}$.
+For $d\in\mathbb{N}$, we define $0_{d}\coloneqq(0,\ldots,0)^{\top}\in\mathbb{R}^{d}$, $1_{d}\coloneqq(1,\ldots,1)^{\top}\in\mathbb{R}^{d}$ and $\mathbb{B}^{d}\coloneqq\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}$.
 For $R>0$, we define the clipping operator $\mathsf{clip}_{R}:\mathbb{R}\to[-R,R]$ by $\mathsf{clip}_{R}(x)\coloneqq-R\vee x\wedge R$ for $x\in\mathbb{R}$.
 Given a Borel measurable set $A\subseteq\mathbb{R}^{d}$, we write $\mathrm{Vol}_{d}(A)$ for the $d$-dimensional Lebesgue measure of $A$.
-For two $\sigma$-finite measures $P$ and $Q$ on a measurable space $(\mathcal{X},\Sigma)$, we say that $P$ is absolutely continuous with respect to $Q$, written as $P\ll Q$, if $P(A)=0$ whenever $A\in\Sigma$ is such that $Q(A)=0$, and in that case write $\frac{\mathrm{d}P}{\mathrm{d}Q}$ for the Radon–Nikodym derivative.
-When $P$ and $Q$ are probability measures, the $\chi^{2}$-divergence from $Q$ to $P$ is defined as
-$\chi^{2}(P,Q)\coloneqq\begin{cases}\int_{\mathcal{X}}\bigl(\frac{\mathrm{d}P}{\mathrm{d}Q}\bigr)^{2}\,\mathrm{d}Q-1\quad&\text{if }P\ll Q\\ \infty&\text{otherwise}.\end{cases}$
+For two $\sigma$-finite measures $P$ and $Q$ on a measurable space $(\mathcal{X},\Sigma)$, we say that $P$ is absolutely continuous with respect to $Q$, written as $P\ll Q$, if $P(A)=0$ for all $A\in\Sigma$ such that $Q(A)=0$, and in that case write $\frac{\mathrm{d}P}{\mathrm{d}Q}$ for the Radon–Nikodym derivative.
+When $P$ and $Q$ are probability measures, the $\chi^{2}$-divergence from $Q$ to $P$ is defined as $\chi^{2}(P,Q)\coloneqq\int_{\mathcal{X}}\bigl(\frac{\mathrm{d}P}{\mathrm{d}Q}\bigr)^{2}\,\mathrm{d}Q-1$ if $P\ll Q$, and $\chi^{2}(P,Q)\coloneqq\infty$ otherwise.
 Further, we write $P^{\otimes n}$ for the $n$-fold product measure of $P$ and $\mathrm{supp}(P)$ for the support of $P$, i.e.
@@
 Transformers), define the empirical risk minimiser $\widehat{f}_{T}$ over $\mathcal{F}$ by
-$\widehat{f}_{T}\in\operatorname*{argmin}_{f\in\mathcal{F}}\widehat{\mathcal{R}}_{T}(f),\quad\text{where }\widehat{\mathcal{R}}_{T}(f)\coloneqq\frac{1}{T}\sum_{t=1}^{T}\bigl\{Y^{(t)}-f(X^{(t)},\mathcal{D}_{n}^{(t)})\bigr\}^{2}$ (2)
-and $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$ are generated according to $\pi$.
+$\widehat{f}_{T}\in\operatorname*{argmin}_{f\in\mathcal{F}}\widehat{\mathcal{R}}_{T}(f),$ (1)
+where $\widehat{\mathcal{R}}_{T}(f)\coloneqq\frac{1}{T}\sum_{t=1}^{T}\bigl\{Y^{(t)}-f(X^{(t)},\mathcal{D}_{n}^{(t)})\bigr\}^{2}$ and $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$ are generated according to $\pi$.
 We remark that, given a new set of examples $D_{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$, the prediction at a query $x\in\mathcal{X}$ is given by $\widehat{f}_{T}(x,D_{n})$, and we do not need to update $\widehat{f}_{T}$ accordingly.
-For a distribution $\nu$ on $\mathcal{P}$, we define the $\nu$-risk of a measurable $g:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ as [^1] We adopt the convention that for $P\in\mathcal{P}$, $\mathbb{E}_{P}$ means that the random variables inside the expectation are generated from $P$, i.e.
+For a distribution $\nu$ on $\mathcal{P}$, we define the $\nu$-risk of a measurable $g:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ as [^1] We adopt the convention that for $P\in\mathcal{P}$, $\mathbb{E}_{P}$ means that the random variables are generated from $P$, i.e.
 $(\mathcal{D}_{n},X,Y)\sim P^{\otimes(n+1)}$, whereas for a distribution $\nu$ on $\mathcal{P}$, $\mathbb{E}_{\nu}\mathbb{E}_{P}$ means that the distribution $P$ is also random and drawn from $\nu$.
-$\mathcal{R}_{\nu}(g)$ $\coloneqq\mathbb{E}_{P\sim\nu}\mathbb{E}_{(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]\equiv\mathbb{E}_{\nu}\mathbb{E}_{P}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr].$
+$\mathcal{R}_{\nu}(g)$ $\coloneqq\mathbb{E}_{P\sim\nu}\mathbb{E}_{(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\bigl[\bigl\{Y\!-\!g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]$
+$\hskip 2.27626pt\equiv\mathbb{E}_{\nu}\mathbb{E}_{P}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr].$
 If $\widetilde{f}_{T}$ is random and depends on the pretraining data $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$, then we interpret $\mathcal{R}_{\nu}(\widetilde{f}_{T})$ as the expectation conditional on the pretraining data, which is assumed to be independent of the test data $(\mathcal{D}_{n},X,Y)$, so $\mathcal{R}_{\nu}(\widetilde{f}_{T})$ is also random.
@@
 $\mathcal{R}^{\mathrm{ICL}}_{\mu}(\widetilde{f}_{T})$ $\coloneqq\mathbb{E}\mathcal{R}_{\mu}(\widetilde{f}_{T})-\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
-$\hskip 2.56064pt=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$ (3)
-where the term $\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ is the oracle risk assuming knowledge of the distribution $P$ that generates $(X,Y)$, and in particular of the true regression function $\mathbb{E}_{P}(Y\,|\,X)$ under $P$, and the equality is shown in the proof of Proposition 1.
+$=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$ (2)
+where the term $\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ is the oracle risk assuming knowledge of the distribution $P$ that generates $(X,Y)$, and in particular of the true regression function $\mathbb{E}_{P}(Y\,|\,X)$ under $P$; the equality (2) is shown in the proof of Proposition 1, see (17).
 Finally, we define the posterior regression function $g_{\pi}:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ (with respect to $\pi$) by
@@
 Thus $g_{\pi}(x,D_{n})$ is the posterior mean of $Y$ given $X=x$ and $\mathcal{D}_{n}=D_{n}$, where $(\mathcal{D}_{n},X,Y)\,|\,P\sim P^{\otimes(n+1)}$ and $P\sim\pi$.
-The following proposition provides a key upper bound on the ICL excess $\mu$-risk, enabling us to study the adaptivity and distributional robustness of ICL.
+The following proposition provides a key upper bound on the ICL excess $\mu$-risk, which is crucial for our study of the adaptivity and distributional robustness of ICL.
 Proposition 1.
 Let $R>0$, and suppose that for all $P\in\mathcal{P}$, we have $|\mathbb{E}_{P}(Y\,|\,X)|\leq R$ almost surely.
-Let $\widehat{f}_{T}$ be defined as in (2), let $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$ and let $\mathcal{E}(\widetilde{f}_{T})\coloneqq\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})-\mathcal{R}_{\pi}(g_{\pi})$.
-Then
+Further let $\widehat{f}_{T}$ be defined as in (1), let $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$ and let $\mathcal{E}(\widetilde{f}_{T})\coloneqq\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})-\mathcal{R}_{\pi}(g_{\pi})$.
+Then we have
 $\mathcal{R}^{\mathrm{ICL}}_{\mu}(\widetilde{f}_{T})\leq 4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr].$ (4)
@@
 Indeed, this is proved in our Proposition 3.
-The second term reflects the proximity of the posterior regression function to the true regression function.
+The second term reflects the proximity of the posterior regression function to the oracle regression function.
 In fact, in Sections 4.1 and 4.2, we will control $\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ uniformly for all $P$ in the support of $\mu$.
 This reveals the interesting feature that only the first term on the right-hand side of (4) depends on the degree of distribution shift, and this term can be made arbitrarily small by increasing the size of the Transformer and the pretraining dataset, so that the second term ultimately determines the rate of convergence.
+Bai et al.
+(2023) showed that Transformers can approximate algorithm selection based on a train-validation split.
+Their arguments proceeded via a different risk decomposition to that in Proposition 1, where Transformers are regarded as approximators of the oracle regression functions $\mathbb{E}_{P}(Y\,|\,X)$; see also Kim, Nakamaki and Suzuki (2024).
+By contrast, in Proposition 1 we view Transformers as approximators of the posterior regression function, which has three key advantages.
+First, we are able to control the ICL excess risk under distribution shift between the pretraining prior and the test distribution.
+Second, we are able to demonstrate adaptivity of Transformers to the difficulty of the test task, instead of obtaining a rate of convergence that is a weighted average over difficulty levels (typically dominated by the difficult tasks).
+Finally, by approximating the posterior regression function, Proposition 1 can be applied to settings where no simple closed form algorithms are available or optimal, thereby (in combination with Theorem 2 below) explaining the empirically-observed superiority of Transformer-based ICL compared to classical algorithms for nonparametric regression (e.g.
+Hollmann et al., 2025).
 3 Transformers, Universal Approximation and Learnability
@@
 In this section, we formally define the classes of (encoder-only) Transformers that we will consider.
+Given $d_{\mathrm{model}}\geq d+2$, called the model dimension, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a query $x\in\mathcal{X}$, the input matrix $Z_{\mathrm{in}}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ is defined as
+$Z_{\mathrm{in}}\equiv Z_{\mathrm{in}}(D_{n},x)\coloneqq\begin{pmatrix}x_{1}^{\top}&y_{1}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ x_{n}^{\top}&y_{n}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ x^{\top}&0&0_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}$ (5)
+where the last column of $Z_{\mathrm{in}}$ is our positional encoding used to identify the query, and we pad each row with zeros to make them $d_{\mathrm{model}}$-dimensional.
 For a vector $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$, let $\operatorname{\mathsf{Softmax}}(v)\in\mathbb{R}^{d}$ denote the vector whose $j$th entry is given by $\exp(v_{j})/\sum_{\ell=1}^{d}\exp(v_{\ell})$.
 Often, we will apply the $\operatorname{\mathsf{Softmax}}$ function to matrices, by which we mean that the $\operatorname{\mathsf{Softmax}}$ function is applied row-wise.
-Given $d_{\mathrm{model}}\geq d+2$, called the model dimension, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a query $x\in\mathcal{X}$, the input matrix is defined as
-$Z_{\mathrm{in}}\equiv Z_{\mathrm{in}}(D_{n},x)\coloneqq\begin{pmatrix}x_{1}^{\top}&y_{1}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ x_{n}^{\top}&y_{n}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ x^{\top}&0&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}},$ (6)
-where the last column of $Z_{\mathrm{in}}$ is our positional encoding used to identify the query, and we pad each row with zeros to make them $d_{\mathrm{model}}$-dimensional.
 A Transformer iteratively applies attention layers and feed forward network (FFN) layers to the input matrix $Z_{\mathrm{in}}$.
@@
 The FFN layer with parameters $\theta_{\mathrm{ffn}}\coloneqq(W_{1},W_{2},v)$ (and activation function $\rho$) is the function $\mathsf{FFN}_{\theta_{\mathrm{ffn}}}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ such that for $n\in[N]$ and $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$,
-$\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(Z)\coloneqq Z+\rho(ZW_{1}+\bm{1}_{n+1}v^{\top})W_{2}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}},$
+$\mathsf{FFN}_{\theta_{\mathrm{ffn}}}\!(Z)\!\coloneqq\!Z\!+\!\rho(ZW_{1}\!+\!1_{n+1}v^{\top})W_{2}\in\mathbb{R}^{(n\!+\!1)\times d_{\mathrm{model}}},$
 where $\rho$ is applied entrywise.
@@
 Finally, a Transformer is a composition of Transformer blocks.
-Definition 4 (Transformers).
-We define $\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$ to be the set of all Transformers $\mathsf{TF}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ of the form
+Definition 4 (Transformer).
+Let $\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$ be the set of all Transformers $\mathsf{TF}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ of the form
 $\mathsf{TF}=\mathsf{TFBlock}^{(L)}\circ\mathsf{TFBlock}^{(L-1)}\circ\cdots\circ\mathsf{TFBlock}^{(1)},$
@@
 For a Transformer $\mathsf{TF}\in\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}$ and a query $x$, we will slightly abuse our notation by writing $\mathsf{TF}(x,D_{n})\coloneqq\mathsf{TF}\bigl(Z_{\mathrm{in}}(D_{n},x)\bigr)$ where the input matrix $Z_{\mathrm{in}}(D_{n},x)$ is defined by (5).
-Finally, we define $\mathsf{Read}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\mathbb{R}$ by $\mathsf{Read}(Z)\coloneqq Z_{n+1,d+1}$ for $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ and $n\in[N]$.
+Finally, we define [^3] More generally, $\mathsf{Read}(Z)$ can be an affine transformation of the last row of $Z$.
+$\mathsf{Read}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\mathbb{R}$ by $\mathsf{Read}(Z)\coloneqq Z_{n+1,d+1}$ for $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ and $n\in[N]$.
 Thus, $\mathsf{Read}\circ\mathsf{TF}(x,D_{n})$ can be interpreted as the predicted value of the response at the query $x$ based on examples $D_{n}$.
@@
 Then for any $\epsilon>0$ and $N\in\mathbb{N}$, there exist $d_{\mathrm{model}}\geq d+2$, $d_{\mathrm{ffn}}\in\mathbb{N}$ and a Transformer $\mathsf{TF}\in\mathcal{F}_{\mathrm{TF}}(3,1,d_{\mathrm{model}},d_{\mathrm{ffn}})$ such that
-$\max_{n\in[N]}\mathbb{E}_{P\sim\pi,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\Bigl\{\bigl(\mathsf{Read}\circ\mathsf{TF}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}\leq\epsilon.$
+$\max_{n\in[N]}\mathbb{E}_{\pi}\mathbb{E}_{P}\Bigl\{\bigl(\mathsf{Read}\circ\mathsf{TF}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}\leq\epsilon.$
 From now on, we will assume that the activation $\rho$ belongs to the set $\{\mathsf{ReLU},\mathsf{GELU},\mathsf{SiLU}\}$, which are commonly used activation functions in modern Transformer models; see (13) for their definitions.
@@
 This proposition confirms formally the intuition provided after Proposition 1 that the first term in the decomposition (4) can be made negligible.
+In other words, Transformers can learn the posterior regression function from data, without knowing the pre-training prior.
 4 Adaptivity and Robustness of ICL
@@
 Moreover, there exists $C_{\alpha,S,d}>0$ such that $\sup_{g\in\mathcal{G}_{\alpha}}\|g\|_{\infty}\leq C_{\alpha,S,d}$; see (56).
+See Figure 2 for realisations of random Besov functions with $\alpha=2$ and $4$.
 Now let $P_{X}$ be a Borel probability distribution on $0,1]^{d}$ with the property that there exists a hypercube $A\subseteq[0,1]^{d}$ such that $P_{X}(A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0})$ for all measurable $A_{0}\subseteq A$.
@@
 Then there exists $C>0$, not depending on $n$, such that
-$\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+d)}$
+$\mathbb{E}_{P_{0}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+d)}$
 for all $n\in\mathbb{N}$.
@@
 Finally, we remark that $d_{\mathrm{model}}^{\circ},d_{\mathrm{ffn}}^{\circ}$ and $T$ should increase with $\kappa$ in order to keep $C$ independent of $\kappa$.
-Thus, for a Transformer with fixed architecture, fine tuning for downstream tasks may still be beneficial in cases where (i) the Transformer is not large enough, (ii) the amount of pretraining data is not sufficient, or (iii) there is significant distribution shift ($\kappa$ is large).
+Thus, for a Transformer with fixed architecture, fine tuning for downstream tasks may still be beneficial in cases where (i) the Transformer is not large enough, (ii) the amount of pretraining data is not sufficient, (iii) there is significant distribution shift ($\kappa$ is large), or (iv) there is shift in the support of the pretraining prior and the test distribution.
 4.2 In-context Learning for Multi-index Models
 To avoid the curse of dimensionality in the previous section, we will now consider multi-index models with [^5] Again, our results in this section extend to any compact domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ by a scaling argument.
-$\mathcal{X}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}\eqqcolon\mathbb{B}^{d}$ and $\mathcal{Y}=\mathbb{R}$.
+$\mathcal{X}=\mathbb{B}^{d}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}$ and $\mathcal{Y}=\mathbb{R}$.
 For $p\in[d]$, let $V_{p}(\mathbb{R}^{d})\coloneqq\bigl\{U\in\mathbb{R}^{d\times p}:U^{\top}U=I_{p}\bigr\}$ denote the set of all projection matrices from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ (also called the Stiefel manifold).
-We will consider regression functions on $\mathbb{B}^{d}$ of the form $x\mapsto g\bigl((U^{\top}x+\bm{1}_{p})/2\bigr)$, where $U\in V_{p}(\mathbb{R}^{d})$ for some effective dimension $p\in[d]$ and Borel measurable $g:[0,1]^{p}\to\mathbb{R}$.
+We will consider regression functions on $\mathbb{B}^{d}$ of the form $x\mapsto g\bigl((U^{\top}x+1_{p})/2\bigr)$, where $U\in V_{p}(\mathbb{R}^{d})$ for some effective dimension $p\in[d]$ and Borel measurable $g:[0,1]^{p}\to\mathbb{R}$.
 Note that we translate and scale $U^{\top}x$ to ensure that it belongs to $0,1]^{p}$.
@@
 Let $\widetilde{\pi}_{\alpha,p}$ be the distribution of the random function $\widetilde{g}_{\alpha,p}:\mathbb{B}^{d}\to\mathbb{R}$ defined by
-$\widetilde{g}_{\alpha,p}(x)\coloneqq\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr),$ (11)
+$\widetilde{g}_{\alpha,p}(x)\coloneqq\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr),$ (10)
 where $U^{(p)}$ is uniformly distributed [^6] One way to generate such a random matrix is to set $U^{(p)}\coloneqq Z(Z^{\top}Z)^{-1/2}$, where $Z$ is a $d\times p$ random matrix with independent $N(0,1)$ entries (e.g.
@@
 We then define $\pi_{\alpha,p}$ to be the distribution of the random measure $P_{\widetilde{g}_{\alpha,p}}$, where the randomness comes from the random regression function $\widetilde{g}_{\alpha,p}\sim\widetilde{\pi}_{\alpha,p}$, and let $\mathcal{P}_{\alpha,p}\coloneqq\mathrm{supp}(\pi_{\alpha,p})$.
+We remark that we index difficulty by the pair of parameters $(\alpha,p)$ in this section.
 For a finite set $\mathcal{A}^{\prime}\subseteq(0,S)\times[d]$ and a sequence of positive weights $(\lambda_{\alpha,p})_{(\alpha,p)\in\mathcal{A}^{\prime}}$ such that $\sum_{(\alpha,p)\in\mathcal{A}^{\prime}}\lambda_{\alpha,p}=1$, our pretraining prior $\pi$ for multi-index models is defined to be the mixture distribution
@@
 Finally, we define $\mathcal{P}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{P}_{\alpha,p}$ and $\mathcal{G}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{G}_{\alpha,p}$.
-Proposition 6 is the analogue for multi-index models of Proposition 4.
+Proposition 6 below is the analogue for multi-index models of Proposition 4.
 Proposition 6.
@@
 Then there exists $C>0$, not depending on $n$, such that
-$\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+r)}$
+$\mathbb{E}_{P_{0}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+r)}$
 for all $n\in\mathbb{N}$.
-Combining Propositions 13 and 6, we can now conclude that a pretrained Transformer is adaptive to the effective dimension of the multi-index models, as well as adapting to the smoothness of the regression function and being robust to distribution shift.
+Combining Propositions 13 and 6, we can now conclude that a pretrained Transformer simultaneously adapts to both the effective dimension of the multi-index models and the smoothness of the regression function, and is robust to distribution shift.
 Theorem 7.
@@
 We remark that, if $\widetilde{f}_{T}$ is defined as in Theorems 5 or 7, then conditional on the pretraining data, we have $\widetilde{f}_{T}\in\widehat{\mathcal{G}}_{n}$.
-Acknowledgement: The research of the first and third author was supported by RJS’s European Research Council Advanced Grant 101019498.
+5 Simulations
+We consider single-index models with $n=30$, $d=5$, $p=1$, $\sigma=0.01$ and four levels of Besov smoothness $\alpha\in\{2,\,2.5,\,3,\,4\}$.
+The generating mechanism of our pretraining data is as described in Section 4.2, with $C_{0}=2$, $(a_{k}^{(\alpha)}),(b_{\ell,\gamma}^{(\alpha)})\overset{\mathrm{iid}}{\sim}\mathrm{Unif}[-1,1]$ in (7) and $P_{X}$ being the uniform distribution on $\mathbb{B}^{d}$.
+For our Transformer, we use a GPT-2 medium-style model (Radford et al., 2019) with $L=24$, $H=16$, $d_{\mathrm{model}}=1024$, $d_{\mathrm{ffn}}=4096$ and GELU activation function.
+Further training details are provided in Appendix I.
+At test time, we randomly generate new single-index models with fixed smoothness levels under two settings:
+  • (i)
+    No distribution shift: the random test functions are generated from the same distribution as the pretraining data of that particular smoothness,
+  • (ii)
+    With distribution shift: similarly as (i), except that the random variables $(a_{k}^{(\alpha)})$ in (7) are sampled from $\mathrm{Unif}[0,1]$ instead of $\mathrm{Unif}[-1,1]$.
+See Figure 2 for random Besov functions with and without distribution shift (blue: no distribution shift; orange: with distribution shift).
+Figure 3(a) shows the mean squared error (MSE) of our Transformer for different levels of $\alpha$; again, blue denotes ‘no distribution shift’ and orange denotes ‘with distribution shift’ at test time.
+We see that as the smoothness level $\alpha$ increases, our Transformer is able to adapt to the smoothness and achieve a faster rate of convergence, moreover, our Transformer is also robust to distribution shift at test time.
+Figure 3(b) plots the logarithm of the MSE against $2\alpha/(2\alpha+1)$, revealing that the log MSE is approximately linear in $2\alpha/(2\alpha+1)$.
+This is in line with the conclusion of Theorem 7.
+(a) $\alpha=2$.
+(b) $\alpha=4$.
+Figure 2: Random Besov functions with smoothness $\alpha\in\{2,4\}$. Blue represents functions with no distribution shift and orange represents functions with distribution shift.
+(a) MSE against $\alpha$.
+(b) Log MSE against $\frac{2\alpha}{2\alpha+1}$.
+Figure 3: MSE and log MSE at test time based on 100,000 random test functions. Error bars represent 90% confidence intervals.
+6 Conclusion
+In this work, we show that pretrained Transformers adapt to different task difficulties and are robust to distribution shifts at test time.
+This is proved by first showing that Transformers are universal approximators of posterior regression functions, and then establishing rates of convergence of the posterior regression functions on each task.
+Our results connect ICL with posterior contraction theory, and provide a rigorous explanation for the empirical success of ICL.
+A natural extension of this work would be to combine our theory with the quantitative approximation theory derived in independent, concurrent work by Wakayama and Suzuki (2025) to provide more explicit values of $d_{\mathrm{model}}^{\circ}$, $d_{\mathrm{ffn}}^{\circ}$ and $T$ that guarantee the conclusions of Theorems 5 and 7.
+More broadly, our work opens the door to providing adaptive and distributionally robust ICL guarantees for Transformers in other settings, including heavy-tailed noise, classification and dependent data.
+Acknowledgement
+The research of the first and third author was supported by RJS’s European Research Council Advanced Grant 101019498.
 Appendix A Notation for the Appendix
@@
 $\leq\sqrt{\mathbb{E}_{\pi}(Z_{P}^{2})\cdot\mathbb{E}_{\pi}\biggl\{\biggl(\frac{\mathrm{d}\mu}{\mathrm{d}\pi}(P)\biggr)^{2}\biggr\}}\leq\sqrt{4R^{2}\mathbb{E}_{\pi}(Z_{P})\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}}=2R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}.$ (14)
-Hence,
-$\mathbb{E}\mathcal{R}_{\mu}(\widetilde{f}_{T})$ $-\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
-$=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{-2Y\widetilde{f}_{T}(X,\mathcal{D}_{n})+\widetilde{f}_{T}^{2}(X,\mathcal{D}_{n})+2Y\mathbb{E}_{P}(Y\,|\,X)-\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}$
-$\hskip-7.00032pt\overset{(i)}{=}\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
-$\leq 2\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
-$\hskip-8.72289pt\overset{(ii)}{\leq}4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr].$
-Here, $(ii)$ follows from (15), and $(i)$ follows since
+Further,
 $\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{Y\widetilde{f}_{T}(X,\mathcal{D}_{n})\bigr\}$ $=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\mathbb{E}_{P}\bigl\{Y\widetilde{f}_{T}(X,\mathcal{D}_{n})\,|\,X,\mathcal{D}_{n}\bigr\}\bigr]$
 $=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P}(Y\,|\,X,\mathcal{D}_{n})\bigr\}$
-$=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P}(Y\,|\,X)\bigr\},$
+$=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P}(Y\,|\,X)\bigr\},$ (15)
 and similarly,
-$\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\bigr\}=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\,|\,X\bigr\}\bigr]=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}.$
-This proves the claim. ∎
+$\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\bigr\}=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\,|\,X\bigr\}\bigr]=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}.$ (16)
+Hence, we deduce that
+$\mathcal{R}_{\mu}^{\mathrm{ICL}}(\widetilde{f}_{T})$ $=\mathbb{E}\mathcal{R}_{\mu}(\widetilde{f}_{T})-\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
+$=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{-2Y\widetilde{f}_{T}(X,\mathcal{D}_{n})+\widetilde{f}_{T}^{2}(X,\mathcal{D}_{n})+2Y\mathbb{E}_{P}(Y\,|\,X)-\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}$
+$\hskip-5.92851pt\overset{(i)}{=}\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ (17)
+$\leq 2\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
+$\hskip-7.34315pt\overset{(ii)}{\leq}4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$
+where $(i)$ follows from (15) and (16), and $(ii)$ follows from (14). ∎
 Appendix C Proofs for Section 3.2
@@
 Throughout this section, we assume that $\bigl((X_{1},Y_{1}),\ldots,(X_{N},Y_{N}),(X,Y)\bigr)\,|\,P\sim P^{\otimes(N+1)}$ with $P\sim\pi$. For $n\in[N]$ and $d_{\mathrm{model}}\geq d+2$, define
-$Z_{\mathrm{in}}^{(n,d_{\mathrm{model}})}\coloneqq\begin{pmatrix}X_{1}^{\top}&Y_{1}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ X_{n}^{\top}&Y_{n}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ X^{\top}&0&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}.$ (17)
+$Z_{\mathrm{in}}^{(n,d_{\mathrm{model}})}\coloneqq\begin{pmatrix}X_{1}^{\top}&Y_{1}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ X_{n}^{\top}&Y_{n}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ X^{\top}&0&0_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}.$ (19)
 We first write down the rows of attention layers and FFN layers before we start.
@@
 $\mathbb{E}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\bigg|\,\widetilde{g}_{1},\ldots,\widetilde{g}_{M}\biggr)\biggr\}$ $=\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\biggr)$
-$\hskip 113.81102pt=\mathbb{E}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\bigg|\,X_{1},Y_{1},\ldots,X_{N},Y_{N},X,Y\biggr)\biggr\}\leq\delta.$
+$=\mathbb{E}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\bigg|\,X_{1},Y_{1},\ldots,X_{N},Y_{N},X,Y\biggr)\biggr\}\leq\delta.$
 Therefore,
@@
 Hence, by definition,
-$\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\biggr)$
-$\hskip 113.81102pt=\mathbb{E}\biggl\{\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}f_{\xi}\bigl(y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\,\biggm|\,X_{1},\ldots,X_{n}\biggr)\biggr\}=0.$
+$\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\biggr)=\mathbb{E}\biggl\{\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}f_{\xi}\bigl(y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\,\biggm|\,X_{1},\ldots,X_{n}\biggr)\biggr\}=0.$
 This holds for all $n\in[N]$, which proves the claim. ∎
@@
 $\mu\Bigl(\bigl\{x\in\mathbb{R}^{d_{1}}:\bigl|f(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)$
-$\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|f(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}$
-$\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|h(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}+\frac{\delta}{3}\leq\delta.$
+$\hskip 113.81102pt\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|f(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}$
+$\hskip 113.81102pt\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|h(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}+\frac{\delta}{3}\leq\delta.$
 This proves the claim when $d_{2}=1$. For $d_{2}>1$, we may approximate the component functions in parallel and apply a union bound to deduce the final result. ∎
@@
 Proof.
-Define a Borel measurable function $f:\mathbb{R}^{d_{\mathrm{model}}}\to\mathbb{R}^{d_{\mathrm{model}}}$ by $f(x,y,\bm{0}_{d_{\mathrm{model}}-d-2},0)\coloneqq f_{1}(x,y)-(x,y,\bm{0}_{d_{\mathrm{model}}-d-2},0)$ and $f(x,0,\bm{0}_{d_{\mathrm{model}}-d-2},1)\coloneqq f_{2}(x)-(x,0,\bm{0}_{d_{\mathrm{model}}-d-2},1)$ for $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$ and set $f$ to be zero otherwise.
+Define a Borel measurable function $f:\mathbb{R}^{d_{\mathrm{model}}}\to\mathbb{R}^{d_{\mathrm{model}}}$ by $f(x,y,0_{d_{\mathrm{model}}-d-2},0)\coloneqq f_{1}(x,y)-(x,y,0_{d_{\mathrm{model}}-d-2},0)$ and $f(x,0,0_{d_{\mathrm{model}}-d-2},1)\coloneqq f_{2}(x)-(x,0,0_{d_{\mathrm{model}}-d-2},1)$ for $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$ and set $f$ to be zero otherwise.
 Let $\mu_{1}$ denote the distribution of $(X,Y)$ and let $\mu_{2}$ denote the marginal distribution of $X$.
 Define a distribution $\mu$ on $\mathbb{R}^{d_{\mathrm{model}}}$ by
-$\mu\biggl(\prod_{j=1}^{d_{\mathrm{model}}}A_{j}\biggr)\coloneqq\frac{\mu_{1}(\prod_{j=1}^{d+1}A_{j})\mathbbm{1}_{\{0\in A_{d_{\mathrm{model}}}\}}+\mu_{2}(\prod_{j=1}^{d}A_{j})\mathbbm{1}_{\{1\in A_{d_{\mathrm{model}}}\}}}{2}\cdot\prod_{j=d+2}^{{d_{\mathrm{model}}}-1}\mathbbm{1}_{\{0\in A_{j}\}},$
+$\mu\biggl(\prod_{j=1}^{d_{\mathrm{model}}}A_{j}\biggr)\coloneqq\frac{\mu_{1}(\prod_{j=1}^{d+1}A_{j})\mathbbold{1}_{\{0\in A_{d_{\mathrm{model}}}\}}+\mu_{2}(\prod_{j=1}^{d}A_{j})\mathbbold{1}_{\{1\in A_{d_{\mathrm{model}}}\}}}{2}\cdot\prod_{j=d+2}^{{d_{\mathrm{model}}}-1}\mathbbold{1}_{\{0\in A_{j}\}},$
 for all measurable sets $A_{1},\ldots,A_{d_{\mathrm{model}}}\subseteq\mathbb{R}$.
 By Lemma 11 and the definition of $\mu$ above, there exists an FFN layer $\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ such that
-$\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ $(X,Y,\bm{0}_{{d_{\mathrm{model}}}-d-2},0)-f_{1}(X,Y)\bigr\|_{\infty}>\epsilon\Bigr)$
-$+\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(X,0,\bm{0}_{{d_{\mathrm{model}}}-d-2},1)-f_{2}(X)\bigr\|_{\infty}>\epsilon\Bigr)\leq\frac{\delta}{N+1}.$
+$\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ $(X,Y,0_{{d_{\mathrm{model}}}-d-2},0)-f_{1}(X,Y)\bigr\|_{\infty}>\epsilon\Bigr)$
+$+\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(X,0,0_{{d_{\mathrm{model}}}-d-2},1)-f_{2}(X)\bigr\|_{\infty}>\epsilon\Bigr)\leq\frac{\delta}{N+1}.$
 We conclude the claim by a union bound. ∎
@@
 Since the final column of $W_{2}$ only affects the final column of $\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})$, we can choose the final column of $W_{2}$ to be zero, so that
-$\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\bigr]_{i,{d_{\mathrm{model}}}}=\bigl[Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})}\bigr]_{i,{d_{\mathrm{model}}}}=\mathbbm{1}_{\{i=n+1\}}=\bigl[U^{(n)}\bigr]_{i,{d_{\mathrm{model}}}}\quad\text{for }i\in[n+1].$ (18)
-Next, we define $Q^{(2)}=K^{(2)}\coloneqq\bm{0}_{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ and define $V^{(2)}\in\mathbb{R}^{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ by $V^{(2)}_{j,2M+j}\coloneqq 1$ for $j\in[2M]$, $V^{(2)}_{4M+1,4M+1}\coloneqq 1$ and zero otherwise.
+$\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\bigr]_{i,{d_{\mathrm{model}}}}=\bigl[Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})}\bigr]_{i,{d_{\mathrm{model}}}}=\mathbbold{1}_{\{i=n+1\}}=\bigl[U^{(n)}\bigr]_{i,{d_{\mathrm{model}}}}\quad\text{for }i\in[n+1].$ (20)
+Next, we define $Q^{(2)}=K^{(2)}\coloneqq 0_{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ and define $V^{(2)}\in\mathbb{R}^{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ by $V^{(2)}_{j,2M+j}\coloneqq 1$ for $j\in[2M]$, $V^{(2)}_{4M+1,4M+1}\coloneqq 1$ and zero otherwise.
 Let $\theta^{(2)}_{\mathrm{attn}}\coloneqq(Q^{(2)},K^{(2)},V^{(2)})$, $\widetilde{Z}^{(n)}\coloneqq\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$ and $\bar{Z}^{(n)}\coloneqq\mathsf{Attn}_{\theta^{(2)}_{\mathrm{attn}}}(\widetilde{Z}^{(n)})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$.
@@
 For $r\in\mathbb{N}$, define the $r$th order difference operator $\Delta_{h}^{r}:L^{1}([0,1]^{d})\rightarrow L^{1}([0,1]^{d})$ by
-$\Delta_{h}^{r}(f)(x)=\mathbbm{1}_{\{x+rh\in[0,1]^{d}\}}\sum_{k=0}^{r}\binom{r}{k}(-1)^{r-k}f(x+kh).$
-Note that $\Delta_{h}^{1}(f)(x)=\{f(x+h)-f(x)\}\mathbbm{1}_{\{x+h\in[0,1]^{d}\}}$ and $\Delta_{h}^{r}(f)(x)=\Delta_{h}^{1}\bigl(\Delta_{h}^{r-1}(f)(x)\bigr)$ if $x+rh\in[0,1]^{d}$.
+$\Delta_{h}^{r}(f)(x)=\mathbbold{1}_{\{x+rh\in[0,1]^{d}\}}\sum_{k=0}^{r}\binom{r}{k}(-1)^{r-k}f(x+kh).$
+Note that $\Delta_{h}^{1}(f)(x)=\{f(x+h)-f(x)\}\mathbbold{1}_{\{x+h\in[0,1]^{d}\}}$ and $\Delta_{h}^{r}(f)(x)=\Delta_{h}^{1}\bigl(\Delta_{h}^{r-1}(f)(x)\bigr)$ if $x+rh\in[0,1]^{d}$.
 Now, for $p\in[1,\infty]$, $f\in L^{p}([0,1]^{d})$ and $t\geq 0$, define the $r$th modulus of smoothness
@@
 $\Phi_{k}(x)\coloneqq\prod_{j=1}^{d}\phi_{k_{j}}(x_{j}).$ (30)
-Further, for $\tau=(\tau_{1},\ldots,\tau_{d})^{\top}\in\{0,1\}^{d}\setminus\{\bm{0}_{d}\}$, $\ell\in[{\ell_{0}}:\infty)$, $k=(k_{1},\ldots,k_{d})^{\top}\in[0:2^{\ell}-1]^{d}$ and $x=(x_{1},\ldots,x_{d})^{\top}\in[0,1]^{d}$, define
+Further, for $\tau=(\tau_{1},\ldots,\tau_{d})^{\top}\in\{0,1\}^{d}\setminus\{0_{d}\}$, $\ell\in[{\ell_{0}}:\infty)$, $k=(k_{1},\ldots,k_{d})^{\top}\in[0:2^{\ell}-1]^{d}$ and $x=(x_{1},\ldots,x_{d})^{\top}\in[0,1]^{d}$, define
 $\Psi_{\ell,k}^{(\tau)}(x)\coloneqq\prod_{j=1}^{d}\psi^{(\tau_{j})}_{\ell,k_{j}}(x_{j}),$ (31)
@@
 To simplify notation, define
-$K\coloneqq[0:2^{\ell_{0}}-1]^{d}\quad\text{and}\quad\Gamma_{\ell}\coloneqq[0:2^{\ell}-1]^{d}\times\bigl(\{0,1\}^{d}\setminus\{\bm{0}_{d}\}\bigr)\text{ for }\ell\in[\ell_{0}:\infty),$ (30)
+$K\coloneqq[0:2^{\ell_{0}}-1]^{d}\quad\text{and}\quad\Gamma_{\ell}\coloneqq[0:2^{\ell}-1]^{d}\times\bigl(\{0,1\}^{d}\setminus\{0_{d}\}\bigr)\text{ for }\ell\in[\ell_{0}:\infty),$ (32)
 and for $\gamma=(k,\tau)\in\Gamma_{\ell}$, we write $\Psi_{\ell,\gamma}\coloneqq\Psi_{\ell,k}^{(\tau)}$. Note that
@@
 for $k\in K$, $\ell\in[\ell_{0}:\infty)$ and $\gamma\in\Gamma_{\ell}$, and that
-$\sum_{k\in K}\mathbbm{1}_{\{\Phi_{k}(x)\neq 0\}}\vee\sup_{\ell\in[\ell_{0}:\infty)}\sum_{\gamma\in\Gamma_{\ell}}\mathbbm{1}_{\{\Psi_{\ell,\gamma}(x)\neq 0\}}\leq C^{\prime}\quad\text{for all }x\in[0,1]^{d}.$ (33)
+[…]
 For $f\in L^{2}([0,1]^{d})$, we have the wavelet decomposition
@@
 $\mathbb{E}_{P_{0}}\bigl[\bigl\{g_{\pi}(X$ $,\mathcal{D}_{n})-g^{0}(X)\bigr\}^{2}\bigr]=\mathbb{E}_{P_{0}}\biggl[\biggl\{\int_{\mathcal{P}}g_{P}(X)\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})-g^{0}(X)\biggr\}^{2}\biggr]$
-$\hskip-7.00032pt\overset{(i)}{\leq}\mathbb{E}_{P_{0}}\biggl[\int_{\mathcal{P}}\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
-$\hskip-8.72289pt\overset{(ii)}{=}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl(\int_{\mathcal{P}}\mathbb{E}_{(X,Y)\sim P_{0}}\bigl[\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\bigr]\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr)$
-$\hskip-10.44545pt\overset{(iii)}{\leq}C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl[\int_{\mathcal{P}}d_{\mathrm{H}}^{2}(P,P_{0})\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
-$\hskip-9.60333pt\overset{(iv)}{\leq}C_{1}^{\prime}M^{2}\epsilon_{n,\beta}^{2}+2C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\bigl[\pi\bigl(\{P\in\mathcal{P}:d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}\}\,\big|\,\mathcal{D}_{n}\bigr)\bigr]$
-$\hskip-7.88077pt\overset{(v)}{\leq}C_{1}\epsilon_{n,\beta}^{2}+C_{2}e^{-c_{3}n\epsilon_{n,\beta}^{2}}.$
+$\hskip-5.92851pt\overset{(i)}{\leq}\mathbb{E}_{P_{0}}\biggl[\int_{\mathcal{P}}\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
+$\hskip-7.34315pt\overset{(ii)}{=}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl(\int_{\mathcal{P}}\mathbb{E}_{(X,Y)\sim P_{0}}\bigl[\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\bigr]\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr)$
+$\hskip-8.7578pt\overset{(iii)}{\leq}C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl[\int_{\mathcal{P}}d_{\mathrm{H}}^{2}(P,P_{0})\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
+$\hskip-8.05411pt\overset{(iv)}{\leq}C_{1}^{\prime}M^{2}\epsilon_{n,\beta}^{2}+2C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\bigl[\pi\bigl(\{P\in\mathcal{P}:d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}\}\,\big|\,\mathcal{D}_{n}\bigr)\bigr]$
+$\hskip-6.63947pt\overset{(v)}{\leq}C_{1}\epsilon_{n,\beta}^{2}+C_{2}e^{-c_{3}n\epsilon_{n,\beta}^{2}}.$
 Here, $(i)$ follows from Jensen’s inequality, $(ii)$ follows by Fubini’s theorem, $(iii)$ follows from Lemma 24(d), $(iv)$ follows by splitting the integral into regions where $d_{\mathrm{H}}(P,P_{0})<M\epsilon_{n,\beta}$ and $d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}$ and $(v)$ follows from Proposition 16 where the sub-Gamma assumption is verified in Lemma 24(c).
@@
 The proof of Proposition 6 relies on three preliminary lemmas that we now formulate. We may write
-$g^{\circ}(x)=h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr),$ (55)
+$g^{\circ}(x)=h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr),$ (57)
 where $U^{\circ}\in V_{r}(\mathbb{R}^{d})$ and $h^{\circ}\in\mathcal{G}_{\beta}^{(r)}$ has wavelet decomposition
@@
 Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that may depend only on $\beta,d,r,c_{0},C_{0}$ and $S$, while $C_{1}^{\prime},C_{2}^{\prime},\ldots$ may in addition depend on $\sigma$.
-Let $\widetilde{g}_{\beta,r}(x)=\widetilde{g}_{\beta}^{(r)}\bigl(\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}\bigr)$ be defined as in (11) so that $\widetilde{g}_{\beta,r}\sim\widetilde{\pi}_{\beta,r}$.
+Let $\widetilde{g}_{\beta,r}(x)=\widetilde{g}_{\beta}^{(r)}\bigl(\frac{(U^{(r)})^{\top}x+1_{r}}{2}\bigr)$ be defined as in (10) so that $\widetilde{g}_{\beta,r}\sim\widetilde{\pi}_{\beta,r}$.
 Observe that $B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})\subseteq B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$ (e.g.
@@
 $|\widetilde{g}_{\beta,r}(x)-g^{\circ}(x)|$
-$\leq\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}\biggr)-\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)\biggr|+\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)\biggr|$
+$\leq\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{(r)})^{\top}x+1_{r}}{2}\biggr)-\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr)\biggr|+\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr)\biggr|$
 $\leq C_{4}n^{-\beta/\{(2\beta+r)\}}+\|\widetilde{g}_{\beta}^{(r)}-h^{\circ}\|_{\infty},$ (62)
-where the second inequality follows from (59), since $\widetilde{g}_{\beta}^{(r)}\in\mathcal{G}_{\beta}^{(r)}\subseteq B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$ and $\bigl\|\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}-\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\bigr\|_{2}\leq\frac{1}{2}\bigl(n^{-\beta/(2\beta+r)}/\log(en)\bigr)^{1/(\beta\wedge 1)}$ on $E$.
+where the second inequality follows from (61), since $\widetilde{g}_{\beta}^{(r)}\in\mathcal{G}_{\beta}^{(r)}\subseteq B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$ and $\bigl\|\frac{(U^{(r)})^{\top}x+1_{r}}{2}-\frac{(U^{\circ})^{\top}x+1_{r}}{2}\bigr\|_{2}\leq\frac{1}{2}\bigl(n^{-\beta/(2\beta+r)}/\log(en)\bigr)^{1/(\beta\wedge 1)}$ on $E$.
 Now by Lemma 26, we have
@@
 $\widetilde{\pi}_{\alpha,p}$ $\Bigl(\Bigl\{g\in\mathcal{G}_{\alpha,p}:\|g-g^{\circ}\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\Bigr\}\Bigr)$
-$=\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\biggr\}$
-$=\mathbb{E}\biggl[\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}\biggr].$ (63)
-Now conditional on $U^{(p)}$, we fix $g^{\prime}\in\mathcal{G}_{\alpha}^{(p)}$ such that $\bigl\|g^{\prime}\bigl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\bigr)-h^{\circ}\bigl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\bigr)\bigr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}$.
+$=\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\biggr\}$
+$=\mathbb{E}\biggl[\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}\biggr].$ (65)
+Now conditional on $U^{(p)}$, we fix $g^{\prime}\in\mathcal{G}_{\alpha}^{(p)}$ such that $\bigl\|g^{\prime}\bigl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\bigr)-h^{\circ}\bigl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\bigr)\bigr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}$.
 We may assume without loss of generality that such $g^{\prime}$ exists, since otherwise the conditional probability in the expectation in (65) would be zero.
 Using the same argument as that in (53), we deduce that
-$\mathbb{P}$ $\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
-$\leq\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
+$\mathbb{P}$ $\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
+$\leq\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
 $=\mathbb{P}\bigl(\|\widetilde{g}_{\alpha}^{(p)}-g^{\prime}\|_{L^{2}(Q(U^{(p)}))}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\big|\,U^{(p)}\bigr),$ (66)
-where $Q(U^{(p)})$ denotes the conditional distribution of $\frac{(U^{(p)})^{\top}X+\bm{1}_{p}}{2}$ given $U^{(p)}$, with $X\sim P_{X}$.
+where $Q(U^{(p)})$ denotes the conditional distribution of $\frac{(U^{(p)})^{\top}X+1_{p}}{2}$ given $U^{(p)}$, with $X\sim P_{X}$.
 By Lemma 27, for every $U\in V_{p}(\mathbb{R}^{d})$, there exists a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ such that $Q(U)(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$.
@@
 $\leq\log\mathcal{N}\bigl(2\sigma\epsilon_{n,\alpha,p}/3,\,\mathcal{G}_{\alpha,p},\,\|\cdot\|_{\infty}\bigr).$
-Recall that $\mathcal{G}_{\alpha}^{(p)}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ and $\mathcal{G}_{\alpha,p}=\mathcal{G}_{\alpha}^{(p)}\circ\bigl\{x\mapsto\frac{U^{\top}x+\bm{1}_{p}}{2}:U\in V_{p}(\mathbb{R}^{d})\bigr\}$.
+Recall that $\mathcal{G}_{\alpha}^{(p)}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ and $\mathcal{G}_{\alpha,p}=\mathcal{G}_{\alpha}^{(p)}\circ\bigl\{x\mapsto\frac{U^{\top}x+1_{p}}{2}:U\in V_{p}(\mathbb{R}^{d})\bigr\}$.
 Now let $\mathcal{N}_{1}$ be a $(\sigma\epsilon_{n,\alpha,p}/3)$-cover of $B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ with respect to $\|\cdot\|_{\infty}$ such that $|\mathcal{N}_{1}|\leq\exp(C_{1}n\epsilon_{n,\alpha,p}^{2})$ for some $C_{1}>0$, which exists by, e.g.
@@
 Thus, for all $x\in\mathbb{B}^{d}$, we have $\|U^{\top}x-(U^{\prime})^{\top}x\|_{2}\leq\delta\sqrt{pd}$, so by (67) and (68), we deduce that
-$\biggl|g\biggl($ $\frac{U^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|$
-$\leq\biggl|g\biggl(\frac{U^{\top}x+\bm{1}_{p}}{2}\biggr)-g\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|+\biggl|g\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|$
+$\biggl|g\biggl($ $\frac{U^{\top}x+1_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)\biggr|$
+$\leq\biggl|g\biggl(\frac{U^{\top}x+1_{p}}{2}\biggr)-g\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)\biggr|+\biggl|g\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)\biggr|$
 $\leq\frac{2\sigma\epsilon_{n,\alpha,p}}{3}.$
@@
 where the final inequality follows from the sub-Gamma assumption on $W(P,Z_{1})$.
-This shows that $\int_{B}W(P,Z_{1})\,\mathrm{d}\Pi(P)$ is sub-Gamma in the left tail with variance parameter $a\epsilon^{2}$ and scale parameter $c$, so $n^{-1}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)$ is also sub-Gamma in the left tail with variance parameter $a\epsilon^{2}/n$ and scale parameter $c/n$ (e.g.
-Samworth and Shah, 2025+, Exercise 10.6.12).
+This shows that $\int_{B}W(P,Z_{1})\,\mathrm{d}\Pi(P)$ is sub-Gamma in the left tail with variance parameter $a\epsilon^{2}$ and scale parameter $c$, so $n^{-1}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)$ is also sub-Gamma in the left tail with variance parameter $a\epsilon^{2}/n$ and scale parameter $c/n$.
 Hence, by (72),
@@
 $\leq\mathbb{P}\biggl(\frac{1}{n}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)<-D\epsilon^{2}\biggr)\leq\exp\biggl(-\frac{D^{2}n\epsilon^{2}}{2(a+cD)}\biggr),$
-where the final inequality follows from Samworth and Shah (2025+, Proposition 10.6.8(a)). ∎
+where the final inequality follows from Bernstein’s inequality (Boucheron, Lugosi and Massart, 2013, Corollary 2.11). ∎
 We next derive some properties about the distributions of nonparametric regression models with Gaussian noise.
@@
 $\mathbb{E}(|W_{2}|^{r})\leq\mathbb{E}(W_{2}^{2})\cdot\Bigl(\frac{2R^{2}}{\sigma^{2}}\Bigr)^{r-2}\leq\frac{r!\epsilon^{2}}{2\cdot 2\sigma^{2}}\Bigl(\frac{2R^{2}}{3\sigma^{2}}\Bigr)^{r-2}.$
-The same bound holds for $W_{2}$, so by Samworth and Shah (2025+, Proposition 10.6.8(b)), $W_{2}$ is sub-Gamma in both tails with variance parameter $\frac{\epsilon^{2}}{2\sigma^{2}}$ and scale parameter $\frac{2R^{2}}{3\sigma^{2}}$.
-Finally, by Samworth and Shah (2025+, Exercise 10.6.12), we have that $W=W_{1}+W_{2}$ is sub-Gamma in both tails with variance parameter $\bigl(8\vee\frac{2}{\sigma^{2}}\bigr)\epsilon^{2}$ and scale parameter $c=\frac{2R}{\sigma}\vee\frac{4R^{2}}{3\sigma^{2}}$.
+The same bound holds for $W_{2}$, so by Boucheron, Lugosi and Massart (2013, Theorem 2.10), $W_{2}$ is sub-Gamma in both tails with variance parameter $\frac{\epsilon^{2}}{2\sigma^{2}}$ and scale parameter $\frac{2R^{2}}{3\sigma^{2}}$.
+Finally, it follows that that $W=W_{1}+W_{2}$ is sub-Gamma in both tails with variance parameter $\bigl(8\vee\frac{2}{\sigma^{2}}\bigr)\epsilon^{2}$ and scale parameter $c=\frac{2R}{\sigma}\vee\frac{4R^{2}}{3\sigma^{2}}$.
 (d) We have
@@
 Let $P_{X}$ be a distribution on $\mathbb{B}^{d}$ that satisfies the assumptions of Lemma 21.
-Let $p\in[d]$, $U\in V_{p}(\mathbb{R}^{d})$, $X\sim P_{X}$ and let $Q$ denote the distribution of $\frac{U^{\top}X+\bm{1}_{p}}{2}$.
+Let $p\in[d]$, $U\in V_{p}(\mathbb{R}^{d})$, $X\sim P_{X}$ and let $Q$ denote the distribution of $\frac{U^{\top}X+1_{p}}{2}$.
 Then there exist $c_{0}^{\prime}\equiv c_{0}^{\prime}(c_{0},\tau,d,p)$ and a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ with the property that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$.
@@
 If $p=d$, then $U\in\mathbb{R}^{d\times d}$ is orthogonal, so for all measurable $A_{0}\subseteq A$,
-$Q\biggl(\frac{U^{\top}A_{0}+\bm{1}_{d}}{2}\biggr)=\mathbb{P}(X\in A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0}).$
-Moreover, $(U^{\top}A+\bm{1}_{d})/2$ is a Euclidean ball of radius $\tau/2$, so it contains a hypercube of side length $\frac{\tau}{2\sqrt{p}}$.
+$Q\biggl(\frac{U^{\top}A_{0}+1_{d}}{2}\biggr)=\mathbb{P}(X\in A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0}).$
+Moreover, $(U^{\top}A+1_{d})/2$ is a Euclidean ball of radius $\tau/2$, so it contains a hypercube of side length $\frac{\tau}{2\sqrt{p}}$.
 Now suppose that $p\in[d-1]$.
@@
 Therefore, writing $v(\tau,d,p)$ for the Lebesgue measure of a Euclidean ball in $\mathbb{R}^{d-p}$ of radius $\tau/2$, we deduce that
-$Q\biggl(\frac{U^{\top}S+\bm{1}_{p}}{2}\biggr)$ $=\mathbb{P}\biggl(\frac{P^{\top}V^{\top}X+\bm{1}_{p}}{2}\in\frac{P^{\top}V^{\top}S+\bm{1}_{p}}{2}\biggr)$
+$Q\biggl(\frac{U^{\top}S+1_{p}}{2}\biggr)$ $=\mathbb{P}\biggl(\frac{P^{\top}V^{\top}X+1_{p}}{2}\in\frac{P^{\top}V^{\top}S+1_{p}}{2}\biggr)$
 $\geq\mathbb{P}(V^{\top}X\in T)=\mathbb{P}(X\in VT)\geq c_{0}\mathrm{Vol}_{d}(VT)=c_{0}\mathrm{Vol}_{d}(T)\geq c_{0}\mathrm{Vol}_{d}(T^{\prime})$
 $=c_{0}\int_{y\in P^{\top}V^{\top}S}\int_{z\in\mathbb{R}^{d-p}:\|z-(V^{\top}o)_{(p+1):d}\|_{2}\leq\tau/2}\,\mathrm{d}z\,\mathrm{d}y$
-$=c_{0}v(\tau,d,p)\mathrm{Vol}_{p}(P^{\top}V^{\top}S)\eqqcolon c_{0}^{\prime}\mathrm{Vol}_{p}\biggl(\frac{U^{\top}S+\bm{1}_{p}}{2}\biggr).$
-Since this holds for all measurable $S\subseteq K$, we deduce that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all $A_{0}^{\prime}\subseteq\frac{U^{\top}K+\bm{1}_{p}}{2}$.
-Now $\frac{U^{\top}K+\bm{1}_{p}}{2}$ is a ball of radius $\tau/4$ in $\mathbb{R}^{p}$, so it contains a hypercube $A^{\prime}$ of side length $\frac{\tau}{2\sqrt{p}}$.
+$=c_{0}v(\tau,d,p)\mathrm{Vol}_{p}(P^{\top}V^{\top}S)\eqqcolon c_{0}^{\prime}\mathrm{Vol}_{p}\biggl(\frac{U^{\top}S+1_{p}}{2}\biggr).$
+Since this holds for all measurable $S\subseteq K$, we deduce that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all $A_{0}^{\prime}\subseteq\frac{U^{\top}K+1_{p}}{2}$.
+Now $\frac{U^{\top}K+1_{p}}{2}$ is a ball of radius $\tau/4$ in $\mathbb{R}^{p}$, so it contains a hypercube $A^{\prime}$ of side length $\frac{\tau}{2\sqrt{p}}$.
 ∎
+Appendix H ICL Example
+When we provide ChatGPT with the first two in-context examples and the query from the right column of Figure 1 (i.e.
+the prompt is ‘Cat $\to$ 10, Fish $\to$ 25, Bird $\to$ ?’), we obtain the following output (though this may vary due to uncertainty in the sampling procedure):
+There isn’t a unique answer from just Cat $\to$ 10 and Fish $\to$ 25.
+With only two example pairs, infinitely many rules can fit them, and those rules would give different values for Bird.
+So Bird $\to$ ?
+is underdetermined (cannot be determined uniquely).
+If you give one more example mapping (any other word $\to$ number), I can pin down the rule and compute Bird.
+When ChatGPT is provided with all four examples and the query from the right column of Figure 1 (i.e.
+now the prompt is ‘Cat $\to$ 10, Fish $\to$ 25, Deer $\to$ 17, Dog $\to$ 13, Bird $\to$ ?’), we obtain the following:
+The pattern is: Number = (number of letters in the word) $\times$ (alphabet position of the first letter) + 1 Check: Cat $\to$ 3 $\times$ C(3) + 1 = 10 Fish $\to$ 4 $\times$ F(6) + 1 = 25 Deer $\to$ 4 $\times$ D(4) + 1 = 17 Dog $\to$ 3 $\times$ D(4) + 1 = 13 So: Bird $\to$ 4 $\times$ B(2) + 1 = 9
+Appendix I Further Details for Simulations in Section 5
+For training, we used AdamW (Loshchilov and Hutter, 2019), with weight decay 0.001 and learning rate 0.0003 (with a cosine annealing schedule).
+We also employed pre-layer normalisation (Xiong et al., 2020) and gradient clipping to stabilise the optimisation.
+Our Transformer was trained for 200,000 optimisation steps.
+At each step, we sampled 50 independent single-index regression functions, and for each function, we randomly generated 12 prompts and queries (i.e.
+the covariate vectors and queries are random while keeping the regression function fixed), yielding a total of 600 pretraining sequences per optimisation step.
+The training was done on 12 NVIDIA A100 GPUs (80GB) over approximately 17 hours.
+To facilitate training, the difficulty of the tasks was progressively increased via a curriculum learning strategy.
+In the initial warm-up phase (the first 10,000 training steps), the ambient dimension $d$ was gradually increased from 1 to 5, while keeping $p=1$ and $\alpha=4$ fixed.
+Subsequently, for the next 50,000 steps, we introduced a smoothness curriculum, where regression functions were drawn from increasingly diverse mixtures whose smoothness parameters were gradually decreased over time.
+Finally, for the remaining 140,000 steps, training proceeded with a uniform mixture over tasks, where the smoothness parameter was sampled uniformly from $\{2,\,2.5,\,3,\,4\}$.